### v1

# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

Abstract
We prove that the Hausdorff dimension of the spectrum of a discrete Schrödinger operator with Sturmian potential of bounded type tends to one as coupling tends to zero. The proof is based on the trace map formalism.
1 Introduction
The object of study is the class of bounded self-adjoint discrete Schrödinger operators given by $H_{\lambda,\alpha,\omega}:\ell^{2}(\mathbb{Z})\rightarrow\ell^{2}(\mathbb{Z})$ via
$\left[H_{\lambda,\alpha,\omega}u\right](n)=u(n+1)+u(n-1)+\left[\lambda\chi_{[1-\alpha,1)}\circ R_{\alpha}^{n}(\omega)\right]u(n),$
[…]
This type of potential gives a self-adjoint operator that is directly associated to a mathematical model of quasicrystals. A survey of results concerning these types of operators, especially Schrödinger operators, can be found in [DEG].
We focus our attention to the case when the frequency has only finitely many different terms that appear in its continued fraction expansion (i.e. of bounded-type) and the coupling is weak (i.e. $\lambda$ is near 00), and investigate the dimension of the spectrum of $H_{\lambda,\alpha,\omega}$ in this context. Since $R_{\alpha}$ is minimal, it is well-known (see Theorem 4.9 in [DF]) that the spectrum of $H_{\lambda,\alpha,\omega}$ is independent of $\omega$, so we will denote it by $\sigma_{\lambda,\alpha}$.
It is known from [BIST] that $\sigma_{\lambda,\alpha}$ is a Cantor set of zero Lebesgue measure. It was even recently proved in [BBL] that all spectral gaps are open. It is natural to ask about the fractal dimension of $\sigma_{\lambda,\alpha}$ and how it may depend on parameters $\lambda$ or $\alpha$. In application, the fractal dimension of $\sigma_{\lambda,\alpha}$ gives information about the solution to the time-dependent Schödinger equation $i\partial_{t}\phi=H_{\lambda,\alpha,\omega}\phi$ (see [JL, L]).
We prove that
Theorem 1.
If $\alpha\in(0,1)$ is irrational of bounded-type, then
$\lim\limits_{\lambda\rightarrow 0}\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)=1.$ (1.1)
This is a generalization of a result from [M] where (1.1) is proved under the assumption that $\alpha$ has an eventually periodic continued fraction expansion. An even earlier version of this result comes from [DG3] where (1.1) is proved for the case when $\alpha=\frac{\sqrt{5}-1}{2}$. The universal technique has been to estimate the thickness of the Cantor set $\sigma_{\lambda,\alpha}$ using well-known theory from hyperbolic dynamics, but in our setting, we must introduce new methods to estimate the thickness.
In the case that $\alpha=\frac{\sqrt{5}-1}{2}$, the operator $H_{\lambda,\alpha,\omega}$ is known as the Fibonacci Hamiltonian. In this case, many questions regarding the box-counting and Hausdorff dimensions of the spectrum have been answered (see [D1, D2, DEG, DGY, Su3] for surveys).
For general coupling constant, and when $\alpha=\frac{\sqrt{5}-1}{2}$, we know from a combination of results from [C, Ca, DG1] that $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)$ is strictly between 00 and $1$. If $\alpha$ has eventually periodic continued fraction expansion, then from [P], we know that $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)$ is analytic in $\lambda>0$, and from [M], we have that $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)=\text{dim}_{B}\left(\sigma_{\lambda,\alpha}\right)$ for all $\lambda>0$.
For large coupling constant and general frequency $\alpha$, when $\lambda>20$ it is known from [LW] that $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)>0$ and there are specific necessary and sufficient conditions, regarding the continued fraction expansion of $\alpha$, for when $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)<1$. When $\lambda\geq 24$, we have a similar result from [LQW], which builds off of a series of results from [FLW, LPW], regarding the upper box counting dimension of $\sigma_{\lambda,\alpha}$. From [DEGT], we know $\text{dim}_{H}\left(\sigma_{\lambda,\frac{\sqrt{5}-1}{2}}\right)\cdot\log\lambda$ converges to $\log(1+\sqrt{5})$ as $\lambda\rightarrow\infty$. A generalization of this result for Lebesgue a.e. $\alpha\in[0,1]\setminus\mathbb{Q}$ was given in [CQ] where it was also shown that there is a set of full Lebesgue measure $\tilde{\mathbb{I}}\subset[0,1]\setminus\mathbb{Q}$ such that for each $(\alpha,\lambda)\in\tilde{\mathbb{I}}\times[24,\infty)$, the Hausdorff and box-counting dimensions of $\sigma_{\lambda,\alpha}$ coincide and are independent of $\alpha$.
For weak coupling constant, many dimension results about the spectrum have been obtained through the analysis of trace map formalism. There is a fundamental link between the spectrum $\sigma_{\lambda,\alpha}$ and the dynamics of the so-called trace maps over a family of cubic surfaces. More precisely, for each $a\in\mathbb{N}$, there is a corresponding trace map $T_{a}$ which leaves invariant a corresponding family of cubic surfaces $\{S_{\lambda}\}_{\lambda>0}$, and the dynamics of the trace maps over $S_{\lambda}$ in the prescribed order that comes exactly from the continued fraction expansion of $\alpha$, gives rich information about the spectrum $\sigma_{\lambda,\alpha}$ (see [Su1, Ra]). For example, the continued fraction expansion of $\alpha=\frac{\sqrt{5}-1}{2}$ contains only 1’s, so here, only the Fibonacci trace map $T_{1}$ plays an important role. Many authors have examined dynamical properties of this trace map (see [BGJ, BR, C, Ca, DG1, DG2, DG3, DG4, DG5, KKT]), where questions concerning hyperbolicity, Markov partitions, entropy, and more have been addressed. It turns out that the non-wandering set of the Fibonacci trace map restricted to any $S_{\lambda}$, $\lambda>0$, is a locally maximal hyperbolic set that is homeomorphic to a Cantor set. This was proved for $\lambda\geq 16$ in [C], for small $\lambda$ in [DG1], and for all $\lambda$ in [Ca]. When $\alpha$ is eventually periodic, one needs to investigate the non-wandering set of a hyperbolic map over $S_{\lambda}$, where this map is precisely a composition of finitely many trace maps, in which case several well-developed techniques from hyperbolic dynamics apply.
In this paper, to work with the weakened assumption that $\alpha$ is of bounded-type, we incorporate non-stationary versions of various dynamical notions such as stable manifolds, Markov rectangles, and a bounded distortion estimate from [BM].
1.1 Background Material
A main ingredient for achieving such results about the spectrum $\sigma_{\lambda,\alpha}$ is to investigate the dynamics of the so-called trace maps.
1.1.1 Trace Maps
Definition (Trace Maps).
We define the trace maps to be the family of maps given by
$T_{a}:\mathbb{R}^{3}\to\mathbb{R}^{3},\quad\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}\mapsto\begin{pmatrix}x_{1}U_{a}(x_{2})-x_{3}U_{a-1}(x_{2})\\ x_{1}U_{a-1}(x_{2})-x_{3}U_{a-2}(x_{2})\\ x_{2}\end{pmatrix},$
where $U_{a}(t)$ are the Chebyshev polynomials of the second kind. That is,
$U_{-1}(t)=0,\ U_{0}(t)=1,\ U_{a+1}(t)=2tU_{a}(t)-U_{a-1}(t).$
For example, the first two trace maps are given by
$T_{1}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{2}-x_{3}\\ x_{1}\\ x_{2}\end{pmatrix}$
and
$T_{2}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}4x_{1}x_{2}^{2}-2x_{2}x_{3}-x_{1}\\ 2x_{1}x_{2}-x_{3}\\ x_{2}\end{pmatrix}.$
It is a calculation to check that for the maps
$G:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ G\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{3}-x_{2}\\ x_{1}\\ x_{3}\end{pmatrix}$
and
$H:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ H\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}x_{1}\\ x_{3}\\ x_{2}\end{pmatrix},$
we have
$T_{1}=G\circ H,\ T_{a}=G^{a-1}\circ T_{1}=G\circ T_{a-1}=G^{a}\circ H.$ (1.2)
For $\lambda>0$, we consider the family of cubic surfaces given by
$S_{\lambda}:=\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-2x_{1}x_{2}x_{3}=1+\frac{\lambda^{2}}{4}\right\}.$
Each surface $S_{\lambda}$ is invariant under each trace map and from here, we will consider the restriction of each $T_{a}$ to $S_{\lambda}$. Given a sequence $\overline{a}=(a_{n})$ in $\mathbb{N}$, denote
$\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right):=\left\{\boldsymbol{x}\in S_{\lambda}:\left(T_{a_{n}}\circ\cdots T_{a_{1}}(\boldsymbol{x})\right)_{n\in\mathbb{N}}\ \text{is bounded}\right\}.$
Given irrational $\alpha\in(0,1)$, we denote its continued fraction expansion by
$\alpha=\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\cfrac{1}{a_{3}+\cdots}}}=:[a_{1},a_{2},a_{3},\ldots]$
where each $a_{k}\in\mathbb{N}$ is uniquely determined. The main link between the trace maps and the spectrum $\sigma_{\lambda,\alpha}$ is the following:
Theorem 2.
For $\lambda>0$ and $\alpha=[a_{1},a_{2},a_{3},\ldots]$ irrational, consider the sequence $\overline{a}=(a_{n})_{n\in\mathbb{N}}$. Then, a real number $E$ belongs to $\sigma_{\lambda,\alpha}$ if and only if
$\left(\frac{E-\lambda}{2},\frac{E}{2},1\right)\in\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right).$
This Theorem is due to a combination of results from [BIST] and [D3]. If we consider the line
$L_{\lambda}:=\left\{\left(\frac{E-\lambda}{2},\frac{E}{2},1\right):E\in\mathbb{R}\right\},$
which is contained in $S_{\lambda}$, then our goal is to estimate the Hausdorff dimension of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ as $\lambda\rightarrow 0$. We will estimate, from below, the Hausdorff dimension of a subset of $\Lambda_{\lambda}^{\text{bnd}}$ that consists of all points that remain in a central portion of the surface $S_{\lambda}$, under our prescribed iteration of trace maps. We will now go through some preliminaries and then outline the structure of the paper in Section 1.2.
1.1.2 Contraction Mappings
The following proposition is a non-stationary version of the contraction mapping principle.
Proposition 1.
Let $(X,d)$ be a complete metric space and $K\subset X$ a compact set. Suppose that there is a $\mu\in(0,1)$ and, for each $n\in\mathbb{N}$, a map $f_{n}:K\rightarrow K$ such that
$d\left(f_{n}(x),f_{n}(y)\right)\leq\mu d(x,y),$
for all $x,y\in K$. Then, the following hold:
  • (1)
    There is a $x^{*}\in K$ such that
    $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*},$
    for any $x\in K$.
  • (2)
    For each $\epsilon>0$, there is a $\epsilon^{\prime}>0$ such that if for each $n\in\mathbb{N}$, there is a mapping $h_{n}:K\rightarrow K$ with
    $d\left(h_{n}(x),h_{n}(y)\right)\leq\mu^{\prime}d(x,y)\ \text{and}\ \sup_{n}d_{C^{0}(K)}(g_{n},h_{n})<\epsilon^{\prime},$
    for some $\mu^{\prime}\in(0,1)$ and for all $x,y\in K$, then
    $d\left(x^{*},y^{*}\right)<\epsilon,$
    where $y^{*}$ is the element such that
    $\lim\limits_{n\rightarrow\infty}h_{1}\circ\cdots\circ h_{n}(x)=y^{*},$
    for each $x\in X$.
Proof.
  • (1)
    By compactness of $K$, there is an $M>0$ so that $d(x,y)\leq M$ for all $x,y\in M$. Then, by induction on $n$, we obtain that for any $x\in K$,
    $d\left(f_{1}\circ\cdots\circ f_{n}(x),f_{1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}d\left(x,f_{n+1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}M,$
    for every $n$ and $m$ such that $n\leq m$, so that $\left(f_{1}\circ\cdots\circ f_{n}(x)\right)$ is a Cauchy sequence. By the completeness of $X$ and compactness of $K$, there must be a $x^{*}\in K$ so that $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*}.$
    Now, let $x^{\prime}\in K$. Then, again by induction on $n$, we have
    $d\left(f_{1}\circ\cdots f_{n}(x),f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)\right)\leq\mu^{n}d\left(x,x^{\prime}\right)\leq\mu^{n}M,$
    for any $n$, and hence $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)=x^{*}.$
  • (2)
    Set $\epsilon^{\prime}=\frac{\epsilon(1-\mu)}{2}$. Then, an induction argument gives that
    $d\left(f_{1}\circ\cdots\circ f_{n}(x),h_{1}\circ\cdots\circ h_{n}(x)\right)\leq\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1},$
    for any $x\in K$. We can deduce that
    $d\left(x^{*},y^{*}\right)\leq d\left(f_{1}\circ\cdots\circ f_{n}(x),x^{*}\right)+d\left(y^{*},h_{1}\circ\cdots\circ h_{n}(x)\right)+\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1}$
    where the last quantity approaches $\frac{\epsilon^{\prime}}{1-\mu}=\frac{\epsilon}{2}$ as $n\rightarrow\infty$.
1.1.3 Hyperbolic Dynamics
Suppose that $M$ is a smooth compact Riemannian surface without boundary. For a $C^{r}$, $r\geq 1$, diffeomorphism $f:M\rightarrow M$, an $f$-invariant set $\Lambda\subset M$ is a hyperbolic set for $f$ provided there are values $0<c<1<\mu$ such that for each $x\in\Lambda$, there exist subspaces $H_{x}$ and $V_{x}$, that depend continuously on $x$, such that $T_{x}M=H_{x}\oplus V_{x}$ and for the sets
$K^{s}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{1}|\leq c|v_{2}|\},$
$K^{u}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{2}|\leq c|v_{1}|\},$
we have
  • ${D_{x}f\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f(x))$ and ${D_{x}f^{-1}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}(x)\right)$
  • $\left\|D_{x}fv\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
  • $\left\|D_{x}f^{-1}v\right\|\geq\mu\|v\|$ for $v\in K^{u}(\boldsymbol{x})$
[…]
It is well-known from [HP] that the stable and unstable sets of $\boldsymbol{x}\in M$ given by
$W^{s}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{n}(\boldsymbol{x})-f^{n}(\boldsymbol{y})\right|=0\right\}$
and
$W^{u}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{-n}(\boldsymbol{x})-f^{-n}(\boldsymbol{y})\right|=0\right\}$
are images of $C^{r}$ injective immersions from $\mathbb{R}$ to $M$. Hence, it is appropriate to refer to these sets as stable, respectively unstable, manifolds. When $\Lambda=M$, and since $\text{dim}(M)=2$, these objects can be used to construct Markov partitions of $M$ that are composed of Markov rectangles of arbitrarily small diameter (see Appendix 2 of [PT]).
If $\overline{f}=(f_{n})$ is a sequence of $C^{1}$ diffeomorphisms of $M$, then we say $\overline{f}$ satisfies a common cone condition on $M$ if for each $x\in M$, there are cones $K^{s}(x)$ and $K^{u}(x)$ as above such that for each $n\in\mathbb{N}$, we have
  • ${D_{x}f_{n}\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f_{n}(x))$ and ${D_{x}f^{-1}_{n}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}_{n}(x)\right)$
  • $\left\|D_{x}f_{n}v\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
  • $\left\|D_{x}f^{-1}_{n}\right\|\geq\mu\|v\|$ for $v\in K^{u}(x)$
[…]
$W^{s}(\boldsymbol{x},m):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{x})-f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{y})\right|=0\right\}$
is the image of an injective $C^{r}$ immersion from $\mathbb{R}$ to $M$.
This result can be taken from the theory of random hyperbolic dynamics (see Chapter 7 of [Ar] and references therein), or one can use arguments similar to Theorem 6.2.8 from [KH] in combonation with Proposition 1, which involves graph transform techniques, to obtain this result. More general versions of this statement for Anosov families have been studied in [Mu]. We use a version of this statement, and sketch the proof, for our relevant family of hyperbolic maps in Section 3.4.
1.1.4 Thickness of Cantor Sets
A fundamental tool we will use to estimate Hausdorff dimension of Cantor sets is the notion of thickness for Cantor sets. Given a Cantor set $\mathcal{C}\subset\mathbb{R}$ with convex hull $I$, we say that $U\subset I\setminus\mathcal{C}$ is a gap of $\mathcal{C}$ if it is a connected component of $I\setminus\mathcal{C}$. Any enumeration $\mathcal{U}=\{U_{n}\}$ of the gaps of $\mathcal{C}$ is called a presentation of $\mathcal{C}$. Given a boundary point $u\in\mathcal{C}$ of some gap $U_{k}$, we call the connected component $B$ contained in $I\setminus\left(U_{1}\cup\dots\cup U_{k}\right)$ that contains $u$ a bridge of $U_{k}$. We define
$\tau(\mathcal{C},\mathcal{U},u):=\frac{|B|}{|U_{k}|},$
and set the thickness of $\mathcal{C}$ to be
$\tau(\mathcal{C}):=\sup\limits_{\mathcal{U}}\inf\limits_{u}\tau(\mathcal{C},\mathcal{U},u).$
It is related to the Hausdorff dimension of $\mathcal{C}$ by the inequality
$\text{dim}_{H}(\mathcal{C})\geq\frac{\log 2}{\log\left(2+\frac{1}{\tau(\mathcal{C})}\right)}.$ (1.3)
(see Chapter 4 of [PT]). It is clear that these notions carry over to a Cantor set that belongs to a $C^{1}$ curve $\gamma\subset S^{1}$ where the gaps and bridges are measured using length relative to the curve $\gamma$.
1.1.5 Continued Fraction Expansion
The approach to our proof begins with some observations about irrational rotations on $S^{1}$ and their connections to Diophantine approximations.
Given an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$, set $p_{n}(\alpha)\in\mathbb{Z}$ and $q_{n}(\alpha)\in\mathbb{N}$ to be the coprime integers such that
$\frac{p_{n}(\alpha)}{q_{n}(\alpha)}=[a_{1},a_{2},\dots,a_{n}].$
We know that these values satisfy the following recurrence equations:
$p_{k+1}(\alpha)=a_{k+1}p_{k}(\alpha)+p_{k-1}(\alpha),\ p_{1}(\alpha)=p_{0}(\alpha)=1$ (1.4)
$q_{k+1}(\alpha)=a_{k+1}q_{k}(\alpha)+q_{k-1}(\alpha),\ q_{1}(\alpha)=a_{1},\ q_{0}(\alpha)=1$ (1.5)
and that
$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|<\frac{1}{q_{k}(\alpha)}.$ (1.6)
In regards to an irrational rotation by $\alpha=[a_{1},a_{2},\dots]$ on $S^{1}$, the Three Distance Theorem (see for example Section 2.1 of [I]) states that for each positive integer $n$, the values $\{m\alpha\}$, $1\leq m\leq n$, divide the unit circle into $n+1$ many intervals of at most three possible lengths, one of which is the sum of the other two. The precise lengths of the intervals come from how $n$ relates to the values $p_{k}(\alpha)$ and $q_{k}(\alpha)$. In particular, if we choose $n=(r+1)q_{k}(\alpha)+q_{k-1}(\alpha)-1$ for some $k\geq 1$ and $1\leq r<a_{k+1}$, then the set of points $\{m\alpha\}$, $1\leq m\leq n$, split $S^{1}$ into $n+1$ many intervals of only two possible lengths given by
$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|\ \text{or}\ \left|q_{k-1}(\alpha)\alpha-p_{k-1}(\alpha)\right|-r\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|,$
and by (1.6), both of these quantities are no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$. As a corollary, we obtain the following result which will be used later on.
Proposition 2.
Suppose that $\mathcal{A}\subset\mathbb{N}$ is finite. For each $\epsilon>0$, there is a $n=n(\epsilon,\mathcal{A})\in\mathbb{N}$ such that for any $\alpha=[a_{1},a_{2},\dots]$ with $a_{i}\in\mathcal{A}$, we must have
$\min\limits_{1\leq m\leq n}\left|m\alpha-y\right|<\epsilon$
for any $y\in S^{1}$.
Proof.
Let $\epsilon>0.$ Set $Y:=\{\alpha=[a_{1},a_{2},\dots]:a_{i}\in\mathcal{A}\}\subset[0,1]\setminus\mathbb{Q}$ and let $a^{\prime}$ be the smallest element of $\mathcal{A}$. Then using the recurrence relation (1.5) and an induction argument, we find that for each $k\geq 1$, there is an $M_{k}>0$ such that
$q_{k}\left(\left[\overline{a^{\prime}}\right]\right)\leq\sup\limits_{\alpha\in Y}|q_{k}(\alpha)|\leq M_{k}.$
Choose $k$ large enough so that
$\frac{1}{q_{k-1}\left(\left[\overline{a^{\prime}}\right]\right)}+\frac{1}{q_{k}\left(\left[\overline{a^{\prime}}\right]\right)}<\epsilon.$
Set $n:=2M_{k}+M_{k-1}-1$ and let $\alpha\in Y$ be arbitrary. By the above discussion, for $n_{k}:=2q_{k}(\alpha)+q_{k-1}(\alpha)-1\leq n$, we have that the points $\{m\alpha\}$, $1\leq m<n_{k}$, split $S^{1}$ into $n_{k}+1$ many intervals of lengths no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$ which is less than $\epsilon$. So, it follows that the set $\{m\alpha\}_{m=1}^{n}$ splits $S^{1}$ into finitely many intervals of length less than $\epsilon$.
1.2 Outline of Paper
Let us fix an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ such that the sequence $(a_{n})_{n\in\mathbb{N}}$ is bounded. Since we are working in a small range of $\lambda$, our hope is to compare the dynamics of our trace maps restricted to $S_{\lambda}$, to the dynamics of the trace maps on $S_{0}$. However, this comparison becomes difficult near the cusps of $S_{\lambda}$ due to how the curvature of the surface changes as $\lambda\rightarrow 0$.
For any points that has a bounded orbit, its orbit may enter a neighborhood of one of the cusps of $S_{\lambda}$ for some amount of time but then return to a central portion of the surface where the dynamics are comparable to those on $S_{0}$. This behavior results in the need for technical distortion estimates, and makes obtaining a common cone condition for our family of trace maps, over these cusps, rather difficult. These details have been written in Section 3 of [DG3] for the map $T_{1}$ only.
To avoid this obstacle, we will first remove neighborhoods of $S_{\lambda}$ near the cusps, so that the remaining surface is a union of five disjoint components, only one of which is bounded. We refer to this bounded component as a central portion of $S_{\lambda}$. We then turn our attention to the set of points from this central portion, whose orbits never leave this region. The points from $L_{\lambda}$ that satisfy this behavior form a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, and it is sufficient to show that this subset has Hausdorff dimension near $1$ as $\lambda$ approaches 00. Provided $\lambda$ is sufficiently small, it turns that each trace map restricted to a desired central portion of $S_{\lambda}$ is a factor of an Anosov map restricted to $\mathbb{T}^{2}$ with a small open set removed, and this map is a perturbation of a hyperbolic linear automorphism.
The structure of the paper is outlined as follows:
  • Section 2 is dedicated to analyzing basic properties of the trace maps $T_{a}$ restricted to each surface $S_{\lambda}$ for a small range of $\lambda$, and we also develop some tools and notations that will be used later in the main proof.
  • In the beginning of Section 3, using some results from 2, we find an open set $O\subset\mathbb{T}^{2}$ composed of four small neighborhoods, a sequence of maps $\left(\tilde{T}_{\lambda,a_{n}}\right)$ defined on $\mathbb{T}^{2}\setminus O$, and a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O$ that tends to the line $\left[0,\frac{1}{2}\right]\times\{0\}$ as $\lambda\rightarrow 0$, such that the set $\mathcal{C}({\lambda},O):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\right\}$ is a Cantor set in $\hat{L}_{\lambda}$ that has dimension equal to the dimension of a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ for a small range of $\lambda$. The rest of the paper is dedicated to showing that for this small range of $\lambda$, the Hausdorff dimension of $C(\lambda,O)$ is close to $1$.
  • Our approach to estimating the dimension of $C(\lambda,O)$ is to show that for small $\lambda$, there is a Cantor set of large thickness that is contained in $C(\lambda,O)$. This is the goal of Section 3, and to do so, we extend each map $\tilde{T}_{\lambda,a_{n}}$ to an Anosov diffeomorphisms on $\mathbb{T}^{2}$, and analyze the non-stationary dynamics of these extended maps. This Anosov family will be constructed to satisfy a common cone condition, and we derive a non-stationary version of the Stable Manifold Theorem for these maps in Section 3.4.
  • Section 3.5 is dedicated to constructing a particular set of dynamical rectangles, which will help us estimate the dimension of $\mathcal{C}(\lambda,O)$. Roughly speaking, we will construct a family of curvilinear rectangles that have two opposite boundaries composed of pieces of the stable manifolds developed in Section 3.4, and these rectangles will have controlled size and contain the open set $O$. We use the common cone conditions to control the size of these rectangles.
  • In Section 3.6, we use the properties of the dynamical rectangles constructed in Section 3.5 to construct a Cantor set of large thickness that is contained in $C(\lambda,O)$. Roughly speaking, this Cantor set is derived from considering the points of $\hat{L}_{\lambda}$ that do not enter a family of associated rectangles containing $O$, under our non-stationary iterations of extended Anosov maps. The dynamical nature of these rectangles allows us to estimate the thickness of the newly derived Cantor set. We use this result to prove Theorem 1.
As a choice of notation, we will always use dist to denote the Hausdorff distance between sets and $d_{C^{0}}$ to denote the $C^{0}$ distance between mappings. With a slight abuse of notation, we will denote $(x,y)$ to be the standard coordinates in both $\mathbb{T}^{2}$ and $\mathbb{R}^{2}$.
2 Trace Map Dynamics as Anosov Diffeomorphisms of $\mathbb{T}^{2}$
[…]
2.1 Torus Model for $\lambda=0$
The surface $S_{0}$ is smooth everywhere except at $P_{1}=(1,1,1),\ P_{2}=(-1,-1,1),\ P_{3}=(1,-1,-1),$ and $P_{4}=(-1,1,-1)$. Let us denote $\mathbb{S}:=S_{0}\cap\{|x_{1}|\leq 1,|x_{2}|\leq 1,|x_{3}|\leq 1\}$. We know
Lemma 1.
$T_{a}\restriction_{\mathbb{S}}$ is a factor of the linear toral hyperbolic map $\hat{T}_{0,a}:\mathbb{T}^{2}\rightarrow\mathbb{T}^{2}$ which is induced by the matrix $A_{a}=\begin{pmatrix}a&1\\ 1&0\end{pmatrix}$, via the semiconjugacy $F(x,y)=(\cos(2\pi(x+y),\cos(2\pi x),\cos(2\pi y))$.
Proof.
This is a direct calculation. ∎
For each $a\in\mathbb{N}$, $A_{a}$ has two perpendicular eigenspaces with corresponding real eigenvalues for which one is less than $1$ in absolute value, and the other is greater than $1$. We call these respective eigenspaces the stable and unstable eigenspace of $A_{a}$ and denote them by $E^{s}_{a}$ and $E^{u}_{a}=\left(E^{s}_{a}\right)^{\perp}$. The corresponding stable and unstable eigenvalues are given by
$\mu^{s}(a):=\frac{a-\sqrt{1+a^{2}}}{2}\ \text{and}\ \mu^{u}(a):=\frac{a+\sqrt{1+a^{2}}}{2}.$
For each $a>1$, the unstable eigenspace $E^{u}_{a}$ lies between the lines $E_{1}^{u}$ and the line $\{y=0\}$.
Let $\beta\in(0.0.1]$, and consider the vector spaces $V_{1}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. We let $K^{u}_{\beta}$ be the collection of vectors that are between $V_{1}(\beta)$ and $V_{2}(\beta)$, which indeed contains each $E^{u}_{a}$. Formally,
$K^{u}_{\beta}:=\left\{c_{1}\begin{pmatrix}1\\ -\beta\end{pmatrix}+c_{2}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}:c_{1}c_{2}\geq 0\right\}$
and we also set $V_{0}(\beta)$ to be the line mid-way between $V_{1}(\beta)$ and $V_{2}(\beta)$. Then, we have the following:
Lemma 2.
For each $\beta\in(0,0.1]$, for the cones given by $K^{u}_{\beta}$ and $K^{s}_{\beta}:=\left(K_{\beta}^{u}\right)^{\perp}$, there exists a $\overline{\mu}_{\beta}>1$ such that for each $a\in\mathbb{N}$, we have
  • $A_{a}^{-1}\left(K^{s}_{\beta}\right)\subset\text{int}K^{s}_{\beta}$ and $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}K^{u}_{\beta}$
  • $\overline{\mu}_{\beta}|v|\leq|A_{a}v|\leq\mu^{u}(a)|v|$ for all $v\in K_{\beta}^{u}$
  • $\overline{\mu}_{\beta}|v|\leq\left|A_{a}^{-1}v\right|\leq\left(\mu^{s}(a)\right)^{-1}|v|$ for all $v\in K_{\beta}^{s}$
Proof.
Let $a\in\mathbb{N}$. To see that $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}K^{u}_{\beta}$, it suffices to check that $A_{a}v\in\text{int}K^{u}_{\beta}$ for the vectors $v\in\partial K^{u}_{\beta}$. The lines that make up $\partial K^{u}_{\beta}$ are $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. A simple calculation gives that
$A_{a}\begin{pmatrix}1\\ -\beta\end{pmatrix}=\begin{pmatrix}a-\beta\\ 1\end{pmatrix}=\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}.$
Clearly $\frac{1}{a-\beta}>-\beta$ and also, $\beta\in(0,0.1]$ implies
$(1+2\beta)\left(a-\beta\right)>1\Rightarrow 1+2\beta>\frac{1}{a-\beta}$
and hence $\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}$ belongs to a line strictly between $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$. A similar calculation shows that the same is true for $A_{a}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}$. The weakest expansion of any vector from $K^{u}_{\beta}$ under any $A_{a}$ is given by iterating vectors furthest away from $E_{1}^{u}$ under the map $A_{1}$. In other words, for any $a$ and $v\in K^{u}_{\beta}$,
$\left|A_{a}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{v}{|v|}\right|$
$\geq\left|A_{1}\frac{\left(1,-\beta\right)^{T}}{\left|\left(1,-\beta\right)^{T}\right|}\right|=\frac{\sqrt{\left(1-\beta\right)^{2}+1}}{\sqrt{\beta^{2}+1}}=:\overline{\mu}_{\beta}.$
The vectors from $K^{u}_{\beta}$ that are expanded the most under $A_{a}$ are the ones from $E^{u}_{a}$, and these vectors are expanded by a factor of precisely $\mu^{u}(a)$. It follows that
$\overline{\mu}_{\beta}|v|\leq\left|A_{a}v\right|\leq\mu^{u}(a)|v|.$
The invariance of $K^{s}_{\beta}$ under $A_{a}^{-1}$ and expansion estimates are checked similarly. ∎
Let $\theta_{0}(\beta)$ denote the angle between boundary lines of $K^{u}_{\beta}$. Notice that $\theta_{0}(\beta)\rightarrow\frac{\pi}{4}$, and $\overline{\mu}_{\beta}\rightarrow\sqrt{2}$, as $\beta\rightarrow 0$. For the rest of the paper, we will fix a $\beta_{0}\in(0,0.1)$, so that
$\theta_{0}(\beta_{0})<\frac{\pi}{3}\ \text{and}\ \overline{\mu}_{\beta_{0}}\geq\sqrt{2}-0.01,$ (2.1)
and denote
$K^{u}:=K^{u}_{\beta_{0}},\ K^{s}:=K^{s}_{\beta_{0}},\ \text{and}\ V:=V_{0}(\beta_{0}).$ (2.2)
Lemma 3.
Let $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ be irrational. The set
$E^{s}_{\alpha}:=\bigcap_{n=1}^{\infty}A_{a_{1}}^{-1}\cdots A_{a_{n}}^{-1}\left(K^{s}\right)$
is a line through 00 of slope $-\frac{1}{\alpha}$.
Proof.
For each $n\in\mathbb{N}$, we consider $\hat{A}_{a_{n}}$ to be the action on $\mathbb{R}\mathbb{P}^{1}$ induced by $A_{a_{n}}$. Denote $\hat{K}^{s}\subset\mathbb{R}\mathbb{P}^{1}$ to be the set obtained from projecting $K^{s}$ onto $\mathbb{R}\mathbb{P}^{1}$, which is compact. Then, from Lemma 2 we have that each map $\hat{A}_{a_{n}}^{-1}$ restricted to $\hat{K}^{s}$ maps $\hat{K}^{s}$ into $\hat{K}^{s}$ and is a contraction with contraction rate no greater than the contraction rate of $\hat{A}_{1}$, which is $\mu^{s}(1)^{2}$. By Proposition 1, there is a $L^{*}\in\hat{K}^{s}$ such that
$\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}(L)=L^{*},$
for any $L\in\hat{K}^{s}$. That is to say, $E^{s}_{\alpha}$ projects uniquely onto $L^{*}$ and is hence a line. In particular, we know that the projection of $\text{span}\left\{(0,1)^{T}\right\}$ onto $\mathbb{R}\mathbb{P}^{1}$, denoted by $\hat{(0,1)}$, lives in $\hat{K}^{s}$, and hence $\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(0,1)}\right)=L^{*}.$ We however see that the line $\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}(1,0)^{T}\right\}$ projects onto $\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(1,0)}\right)$. That $E^{s}_{\alpha}$ has slope $-\frac{1}{\alpha}$ follows from
Claim 1.
The line $\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}\begin{pmatrix}0\\ 1\end{pmatrix}\right\}$ has slope $-\frac{1}{[a_{1},a_{2},\dots,a_{n}]}$.
Proof of Claim 1..
First, notice that for any number $b$ and any $a\in\mathbb{N}$, one has
$A^{-1}_{a}\begin{pmatrix}-b\\ 1\end{pmatrix}=\begin{pmatrix}0&1\\ 1&-a\end{pmatrix}\begin{pmatrix}-b\\ 1\end{pmatrix}=(a+b)\begin{pmatrix}-\frac{1}{a+b}\\ 1\end{pmatrix}.$
An induction argument then gives that
$A_{a_{1}}^{-1}\cdots A_{a_{n}}^{-1}\begin{pmatrix}-b\\ 1\end{pmatrix}=(\text{Constant})\begin{pmatrix}-\alpha_{n}(b)\\ 1\end{pmatrix}$
where
[…]
It follows that
$\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}\begin{pmatrix}0\\ 1\end{pmatrix}\right\}=\text{span}\left\{\begin{pmatrix}-\frac{1}{[a_{1},\dots,a_{n}]}\\ 1\end{pmatrix}\right\},$
and hence has slope $-\frac{1}{[a_{1},\dots,a_{n}]}$.
Projecting these lines onto $\mathbb{T}^{2}$, we have that each line $\pi\left(E^{s}_{\alpha}\right)$ is dense in $\mathbb{T}^{2}$. If $Q$ is a rational point in $\mathbb{T}^{2}$, and $L\subset\pi\left(E^{s}_{\alpha}\right)$ is a line segment with midpoint 00, the length of $L$ and continued fraction expansion of $\alpha$ determine the distance from $L$ and $Q$. From Proposition 2, we have the following:
Lemma 4.
Let $Q$ and $Q^{\prime}$ be distinct rational points on $\mathbb{T}^{2}$, and suppose that $\mathcal{A}\subset\mathbb{N}$ is a finite set.
  • 1.)
    For each $C^{*}>0$, there is a $l^{*}=l^{*}\left(C^{*},\mathcal{A}\right)>0$ such that if $L$ is a line with midpoint $Q$, irrational slope $[a_{1},a_{2},\dots]$, $a_{i}\in\mathcal{A}$, and $\text{length}(L)>l^{*}$, then $\text{dist}(L,Q^{\prime})<C^{*}$
  • 2.)
    For each $l>0$, there is a constant $C^{**}=C^{**}\left(l,Q,Q^{\prime},\mathcal{A}\right)>0$ such that if $L$ is a line with midpoint $Q$, irrational slope $[a_{1},a_{2},\dots]$, $a_{i}\in\mathcal{A}$, and $\text{length}(L)<l$, then $\text{dist}(L,Q^{\prime})>C^{**}$
Proof.
Without loss of generality, let us assume $Q=0$.
Let $C^{*}>0$. By Proposition 2, there is a $n=n\left(C^{*},\mathcal{A}\right)$ so that for any $\alpha=[a_{1},a_{2},\dots]$ where $a_{i}\in\mathcal{A}$, we must have
$\min\limits_{1\leq k\leq n}|k\alpha-y|<C^{*}$
for any $y\in S^{1}$.
Consider the vertical line through $Q^{\prime}=(x_{0},y_{0})$ given by $\text{vert}(Q^{\prime})=\left\{(x_{0},y):y\in S^{1}\right\}$. Then, there is a $l^{*}$ dependent on $n$ such that any line $L$ with midpoint 00 and irrational slope $\alpha=[a_{1},a_{2},\dots]$, where $a_{i}\in\mathcal{A}$, satisfying $\text{length}(L)>l^{*}$ must intersect $\text{vert}(Q^{\prime})$ at least $2n$ many times.
At least $n$-many of the intersection points take the form $(x_{0},z_{0}+k\alpha)$, $1\leq k\leq n$, for some $z_{0}\in S^{1}$. It follows that
$\text{dist}(L,Q^{\prime})\leq\min\limits_{1\leq k\leq n}|(x_{0},y_{0})-(x_{0},z_{0}+k\alpha)|$
$=\min\limits_{1\leq k\leq n}|(y_{0}-z_{0})-k\alpha|<C^{*},$
so that (1) holds.
Let $l>0$ and consider the set of lines $\mathbb{L}(\mathcal{A})$ in $\mathbb{R}^{2}$ that have midpoint 00, length exactly $l$, and irrational slope of the form $[a_{1},a_{2},\dots]$, $a_{i}\in\mathcal{A}$. If we consider the natural projection of $p:\mathbb{R}^{2}\setminus\{0\}\rightarrow\mathbb{R}\mathbb{P}^{1}$, then $p(\mathbb{L}(\mathcal{A})\setminus\{0\})$ is a Cantor set on the circle $\mathbb{R}\mathbb{P}^{1}$, and is hence closed. By continuity of $p$, we must have that $p^{-1}(p(\mathbb{L}(\mathcal{A})\setminus\{0\}))\cup\{0\}$ is a closed set in $\mathbb{R}^{2}$, and since $\mathbb{L}(\mathcal{A})=\left[p^{-1}(p(\mathbb{L}(\mathcal{A})\setminus\{0\}))\cup\{0\}\right]\cap B_{l}(0)$, it must be a compact set. Since the lines in $\mathbb{L}(\mathcal{A})$ have irrational slope, this set is disjoint from the finite set $\pi^{-1}\left(Q^{\prime}\right)\cap B_{2l}(0)$. Hence, there is a $C^{**}_{0}>0$ such that
$\text{dist}\left(\mathbb{L}(\mathcal{A}),\pi^{-1}\left(Q^{\prime}\right)\cap B_{2l}(0)\right)\geq C^{**}_{0},$
and the result follows from here.
Remark 1.
We note that by a symmetric argument, Lemma 4 holds if the lines $L$ are replaced with lines normal to $L$ of equal length with midpoint 00. In other words, the statement holds if we instead require the lines $L$ to have slope $-\frac{1}{[a_{1},a_{2},\dots]}$ such that $a_{i}\in\mathcal{A}$.
2.2 Torus Model for Small $\lambda>0$
Let us set $Q_{i}:=F^{-1}(\{P_{i}\})$, for $i=1,\dots,4$, so that $Q_{1}=(0,0)$, $Q_{2}=\left(\frac{1}{2},0\right)$, $Q_{3}=\left(\frac{1}{2},\frac{1}{2}\right)$, and $Q_{4}=\left(0,\frac{1}{2}\right)$. The next Lemma relates the dynamics of each $T_{a}$ on $S_{\lambda}$ to the dynamics of a toral map.
Lemma 5.
For each $\rho>0$ and $a\in\mathbb{N}$, there is a $\tilde{\lambda}(\rho,a)>0$ and $r(\rho)>0$, such that if $\lambda\in\left[0,\tilde{\lambda}(\rho,a)\right]$, then
  • (1)
    For $O_{\rho}:=\bigcup_{i=1}^{4}B_{\rho}(P_{i})$, the set $S_{\lambda}\setminus O_{\rho}$ consists of five disjoint connected components, only one of which is bounded, which we will denote by $\mathbb{S}_{\lambda,O_{\rho}}$. Furthermore, there is a smooth injective map $\pi_{\lambda}:\mathbb{S}_{\lambda,O_{\rho}}\rightarrow\mathbb{S}$, that only depends on $\lambda$, that approaches the identity map in the $C^{1}$-sense as $\lambda\rightarrow 0$
  • (2)
    There is an open set $\tilde{O}_{r(\rho)}\subset\mathbb{T}^{2}$ which consists of four disjoint neighborhoods, each centered at a distinct $Q_{i}$ and of diameter less than $\frac{r(\rho)}{2}$, and a map
    $\tilde{T}_{\lambda,a}:\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}\rightarrow\mathbb{T}^{2}$
    satisfying
    $F_{\lambda}\circ\tilde{T}_{\lambda,a}=T_{a}\restriction_{\mathbb{S}_{\lambda,O_{\rho}}}\circ F_{\lambda},$
    where $F_{\lambda}:=\pi^{-1}_{\lambda}\circ F$.
Moreover, we have $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$, as $\lambda\rightarrow 0$, and $r(\rho)\rightarrow 0$ as $\rho\rightarrow 0$.
Proof.
One can choose $\rho^{\prime}\in(0,\rho)$ such that for $O_{\rho}=\bigcup_{i=1}^{4}B_{\rho}$ and $O_{\rho^{\prime}}=\bigcup_{i=1}^{4}B_{\rho^{\prime}},$ we must have
$T_{a}^{-1}\left(O_{\rho^{\prime}}\right)\subset O_{\rho}.$
There is a $\tilde{\lambda}(\rho,a)>0$ such that for each $\lambda\in\left[0,\tilde{\lambda}(\rho,a)\right]$, we have that $S_{\lambda}\setminus O_{\rho}\ \text{and}\ S_{\lambda}\setminus O_{\rho^{\prime}}$ are both unions of five disjoint connected components, only one of which is bounded.
Denoting the bounded component of $S_{\lambda}\setminus O_{\rho^{\prime}}$ by $\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, there is a smooth injective map $\pi_{\lambda}:\mathbb{S}_{\lambda,O_{\rho^{\prime}}}\rightarrow\mathbb{S}$, that is defined as follows:
For each $\boldsymbol{y}\in\mathbb{S}$, consider the normal line $N_{\boldsymbol{y}}$ to $\mathbb{S}$ at $\boldsymbol{y}$ that is pointing outward. Either $N_{\boldsymbol{y}}$ intersects $\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$ uniquely, or not at all. In the former case, we set $\pi_{\lambda}^{-1}\left(\boldsymbol{y}\right)=N_{\boldsymbol{y}}\cap\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, and this gives us (1).
Denote the bounded component of $S_{\lambda}\setminus O_{\rho}$ by $\mathbb{S}_{\lambda,O_{\rho}}$. Then, there is a $r(\rho)>0$ and an open set $\tilde{O}_{r(\rho)}\subset\mathbb{T}^{2}$ which consists of neighborhoods of $Q_{i}$, $i=1,\dots,4$, each of diameter less than $\frac{r(\rho)}{2}$, such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho}}\right)=\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$.
Also, there is an open set $O^{\prime}\subset\tilde{O}_{r(\rho)}$ such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho^{\prime}}}\right)=\mathbb{T}^{2}\setminus O^{\prime}$. Let us denote $F_{\lambda}:=\pi_{\lambda}^{-1}\circ F$.
Let $P\in\mathbb{T}^{2}$. Then, $F_{\lambda}(P)$ is contained in $\mathbb{S}_{\lambda,O_{\rho}}$ and $\left(T_{a}\circ F_{\lambda}\right)(P)\in\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, so that
$\left(F^{-1}_{\lambda}\circ T_{a}\circ F_{\lambda}\right)(P)$
is a set consisting of two points $Q,Q^{\prime}\in\mathbb{T}^{2}\setminus O^{\prime}$. If $\tilde{\lambda}(\rho,a)$ is small enough, then one of the two points in $\{Q,Q^{\prime}\}$ is much closer to $\hat{T}_{0,a}$ than the other.
Let us choose $\tilde{T}_{\lambda,a}(P)$ to be the element from $\{Q,Q^{\prime}\}$ that is closest to $\hat{T}_{0,a}(P)$. Then, $\tilde{T}_{\lambda,a}:\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}\rightarrow\mathbb{T}^{2}\setminus O^{\prime}$ is well-defined, and satisfies
$F_{\lambda}\circ\tilde{T}_{\lambda,a}=T_{a}\restriction_{\mathbb{S}_{\lambda,O_{\rho}}}\circ F_{\lambda}.$
Since $\left\|\pi_{\lambda}\circ T_{a}\circ\pi^{-1}_{\lambda}\restriction_{\mathbb{S}_{0,O_{\rho}}}-T_{a}\restriction_{\mathbb{S}_{0,O_{\rho}}}\right\|_{C^{2}}\rightarrow 0$ as $\lambda\rightarrow 0$, it follows that $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus O_{r(\rho)}$, as $\lambda\rightarrow 0$.
Lastly, since $\text{diam}\left(B_{\rho}(P_{i})\right)\rightarrow 0$ as $\rho\rightarrow 0$ for each $i=1,\dots,4$, we can choose $r(\rho)$ so that $r(\rho)\rightarrow 0$ as $\rho\rightarrow 0$.
2.3 Main Problem Setup
We will now setup various tools and notations that will be helpful in the main proof.
2.3.1 Boxes and Local Cones
Recalling our definition of $V$ from (2.2), consider the projections $\pi_{V}:V\oplus V^{\perp}\rightarrow V$ and $\pi_{V^{\perp}}:V\oplus V^{\perp}\rightarrow V^{\perp}$, and set .$\|.\|_{V}:=\left|\pi_{V}(.)\right|$ and .$\|.\|_{V^{\perp}}:=\left|\pi_{V^{\perp}}(.)\right|$. For $\boldsymbol{Q}\in\mathbb{R}^{2}$, consider the rectangular neighborhood given by
$\text{Box}_{r}(\boldsymbol{Q}):=\left\{\boldsymbol{P}:\|\boldsymbol{P}-\boldsymbol{Q}\|_{V}\leq r\right\}\times\left\{\boldsymbol{P}:\|\boldsymbol{P}-\boldsymbol{Q}\|_{V^{\perp}}\leq r\right\}.$
Remark 2.
Suppose that $a\in\mathbb{N}$ and consider the square $\text{Box}_{1}(0)$. Then, from our choice of $\beta_{0}$ (see 2.2), we know that the length of any segment contained in $K^{u}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length no greater than $\frac{2}{\sqrt{3}}$ and length no less than $1$. If we iterate this segment under $A_{a}$, the its length will be greater than $\overline{\mu}_{\beta_{0}}$. However, from (2.1), we have
$\overline{\mu}_{\beta_{0}}\geq\sqrt{2}-0.01>\frac{2}{\sqrt{3}}.$
This calculation gives that for any line segment $L$ through 00 that is contained in $K^{u}$ and connects the boundary lines of $\text{Box}_{1}(0)$ that are parallel to $V^{\perp}$, we must have that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$ (see Figure 1).
If $\beta>\beta_{0}$ is close enough to $\beta_{0}$, then same property will hold for $K^{u}_{\beta}$ restricted to $\text{Box}_{1}(0)$. Notice that in this case, $V(\beta_{0})$ is not the center line of the cone $K^{u}_{\beta}$, but the angle between its center line and $V(\beta_{0})$ is small.
The property also holds for a sufficiently small perturbation of each map $A_{a}$.
[…]
Figure 1: Illustration of Remark 2
[…]
$\phi_{\boldsymbol{Q}}:=\pi\restriction_{\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})}:\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\rightarrow\pi\left(\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right),$ (2.3)
where $\pi\left(\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ is the rectangular neighborhood centered at $\pi(\boldsymbol{Q})$, with sides parallel to the lines $\pi(V)$ and $\pi(V^{\perp})$, respectively, of lengths given by $\frac{1}{5}$.
For each $r\in\left(0,\frac{1}{10}\right)$, let us denote $\mathbb{B}_{r}(Q):=\pi\left(\text{Box}_{r}\left(\pi^{-1}(Q)\right)\right)$, which is the rectangular neighborhood centered at $Q$ of side lengths $2r$. We set $\partial^{s}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V^{\perp}\right)$ and $\partial^{u}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V\right)$ (see Figure 2).
In particular, let us set $\mathbb{B}_{i}:=\mathbb{B}_{\frac{1}{10}}(Q_{i})$, $\partial^{s}\mathbb{B}_{i}:=\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q_{i})$ and $\partial^{u}\mathbb{B}_{i}:=\partial^{u}\mathbb{B}_{\frac{1}{10}}(Q_{i})$.
[…]
Figure 2: Illustration of local boxes.
For any cone $K$ containing $K^{u}$ and transversal to $K^{s}$, we define
$K_{-}:=K^{u}\cap\{x\leq 0\}\ \text{and}\ K_{+}:=K\cap\{x\geq 0\}.$ (2.4)
For a cone $K\subset\mathbb{R}^{2}$, define the local projected cone $\tilde{K}(Q):=\pi\left(\left(\boldsymbol{Q}+K\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$. If $K$ is a cone containing $K^{u}$ and transversal to $K^{s}$, then set $\tilde{K}_{+}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{+}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ and $\tilde{K}_{-}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{-}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ (see Figure 3).
[…]
Figure 3: Illustration of a local cone.
With this in mind, we set +$\partial^{s}_{+}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{+}^{u}(Q)$ and $\partial^{s}_{-}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{-}^{u}(Q)$.
We will let $K^{s,u}(x,y)$ be the cone $K^{s,u}$ contained in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, and set
$\mathcal{K}^{s}:=\left\{K^{s}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}\ \text{and}\ \mathcal{K}^{u}:=\left\{K^{u}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}},$ (2.5)
and refer to these collections as stable, respectively unstable, cone fields.
Remark 3.
There is a $C_{1}>0$ such that if $\gamma\subset\mathbb{T}^{2}$ is a curve tangent to $\mathcal{K}^{u}$, and is contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, for some $Q\in\mathbb{T}^{2}$, with endpoints $P_{1}$ and $P_{2}$, then
$\text{length}(\gamma)\leq C_{1}\left|P_{1}-P_{2}\right|.$
2.3.2 Dynamical Rectangles
Definition.
A dynamical rectangle (or rectangle) is a curvilinear rectangle $R$, such that two opposite boundary curves of $R$ are $C^{1}$ curves tangent to $\mathcal{K}^{s}$, referred to as the stable boundaries, denoted by $\partial^{s}R$, and the other two boundary curves are transversal to $\mathcal{K}^{s}$, referred to as the non-stable boundaries, and denoted by $\partial^{ns}R$.
We say that a $C^{1}$ curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ that is tangent to $\mathcal{K}^{s}$ connects opposite boundaries of $\tilde{K}^{u}(Q)$ if its endpoints belong to $\partial\tilde{K}_{+}^{u}(Q)$. For any $Q\in\mathbb{T}^{2}$ and points $P,P^{\prime}\in\mathbb{B}_{\frac{1}{10}}(Q)$, we say that a curve $\gamma$ connects $P$ and $P^{\prime}$ if these are the endpoints of $\gamma$, and $\gamma$ is strictly contained in $\overline{B_{\frac{1}{10}}(Q)}$. If $\gamma_{1}$ and $\gamma_{2}$ are disjoint curves contained in $B_{\frac{1}{10}}(Q)$, then $\gamma$ connects these curves if it connects points from these curves.
For each $Q\in\mathbb{T}^{2}$ and cone $K\supset K^{u}$ transversal to $K^{s}$, we say that two $C^{1}$ curves $\gamma_{1}$ and $\gamma_{2}$ contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, tangent to $\mathcal{K}^{s}$, are centered at $Q$ with respect to $\tilde{K}(Q)$, if there is a sub-curve of $\gamma_{1}$ that connects opposite boundaries of $\tilde{K}_{+}(Q)$ and does not intersect $\tilde{K}_{-}(Q)$, and if there is a sub-curve of $\gamma_{2}$ that connects opposite boundaries of $\tilde{K}_{-}(Q)$ and does not intersect $\tilde{K}_{+}(Q)$, or vice versa (see Figure 4).
Notice that any pair of curves centered at $Q$ with respect to $\tilde{K}(Q)$, must be centered at $Q$ with respect to $\tilde{K}^{u}(Q)$. If not specified otherwise, we say that two curves are centered at $Q$, if they are centered at $Q$ with respect to $\tilde{K}^{u}(Q)$.
A rectangle $R\subset B_{\frac{1}{10}}(Q)$ that contains $Q$ in its interior is centered at $Q$ with respect to $\tilde{K}(Q)$ if its stable boundary curves are centered at $Q$ with respect to $\tilde{K}(Q)$ (see Figure 4). We denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$. It should also be noted that ±$\partial^{s}_{\pm}R\cap\tilde{K}^{u}_{\mp}(Q_{i})=\emptyset$.
[…]
$\partial^{s}_{-}R$$\partial^{ns}R$$R$+$\partial^{s}_{+}R$$Q$$\tilde{K}(Q)$$\boldsymbol{-}$$\boldsymbol{+}$
Figure 4: The right figure depicts two curves $\gamma_{1}$ and $\gamma_{2}$ that are centered at $Q$ with respect to a cone $\tilde{K}(Q)$, and the left figure depicts a dynamical rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$.
Remark 4.
Notice there is a $C^{*}_{0}>0$ such that for each $i=1,\dots,4$, if $\gamma$ is a $C^{1}$ curve tangent to $\mathcal{K}^{s}$ that is contained in $\overline{\mathbb{B}}_{i}$ satisfying $\text{dist}(\gamma,Q_{i})<C^{*}_{0}$, then it is disjoint from $\partial^{s}\mathbb{B}_{i}$ (see Figure 5). Moreover, if $\kappa$ is a curve tangent to $\mathcal{K}^{u}$ with endpoints $P_{1}$ and $P_{2}$ satisfying $P_{1}\in\tilde{K}_{+}(Q_{i})$, $P_{2}\in\tilde{K}_{-}(Q_{i})$, and $\max\left\{\left|P_{1}-Q_{i}\right|,\left|P_{2}-Q_{i}\right|\right\}<C^{*}_{0}$, we must have $\kappa\subset\mathbb{B}_{i}$. That is $\kappa$ is disjoint from $\partial^{u}\mathbb{B}_{i}$.
$Q$$\tilde{K}^{s}(Q)$$\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q)$$\gamma$
$\tilde{K}(Q)$$\tilde{K}^{u}(Q)$$Q$$\kappa$
Figure 5: Illustration of Remark 4. On the top figure, any curve $\gamma$ tangent to $\mathcal{K}^{s}$ that intersects the center “diamond” must stay between the dotted lines. On the bottom figure, any curve $\kappa$ tangent to $\mathcal{K}^{u}$ with endpoints in the center region contained in $\tilde{K}(Q)$, must be contained within the dotted lines. So, one can choose a neighborhood $O$ of $Q$ accordingly.
Remark 5.
If $K\supsetneq K^{u}$ is a cone disjoint from $K^{s}$, then for any rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$, there is a neighborhood $O$ of $Q$ such that for any curve $\gamma$ tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects $O$ must not intersect $\partial^{ns}R$ (see Figure 6). The size of the neighborhood $O$ depends on the cone $K$ and $\text{dist}(\partial R,Q)$.
In particular, if $\{R_{n}\}$ is a family of rectangles that are centered at $Q$ with respect to $\tilde{K}(Q)$ and there is a $l^{*}>0$ such that $\inf_{n}\text{dist}\left(\partial R_{n},Q\right)\geq l^{*}$, then there is a $r=r(K,l^{*})>0$ such that for each $n$ and any curve $\gamma$ tangent to $\mathcal{K}^{u}$, we must have that any component of $\gamma\cap R_{n}$ that intersects $B_{r}(Q)$, must not intersect $\partial^{ns}R_{n}$. That is, for any curve $\gamma$ tangent to $\mathcal{K}^{u}$ connecting $\partial_{+}^{s}R_{n}$ and $\partial_{-}^{s}R_{n}$ with $\gamma\cap B_{r}(Q)\neq\emptyset$, we must have $\gamma\subset R_{n}$ and $\gamma\cap\partial^{ns}R=\emptyset$.
$R$$\tilde{K}(Q)$$\tilde{K}^{u}(Q)$$\gamma$
Figure 6: Illustration of Remark 5. The rectangle $R$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and has the property that for any curve $\gamma$ that is tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects the center quadrilateral must stay within the dotted lines and hence not intersect $\partial^{ns}R$.
3 Main Estimates and Non-Stationary Trace Map Dynamics
For the rest of the paper, let us consider an arbitrary finite set $\mathcal{A}\subset\mathbb{N}$ and an arbitrary sequence $\overline{a}:=(a_{n})$ such that $a_{n}\in\mathcal{A}$ for all $n\in\mathbb{N}$.
Let $M>0$. We will show that there is a $\lambda_{0}>0$ such that for all $\lambda\in[0,\lambda_{0}]$, we have
$\text{dim}_{H}\left(L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}.$ (3.1)
Our approach is to invoke Lemma 5 and choose a neighborhood $O$, that is composed of small neighborhoods containing the points $Q_{i},$ so that the set
$C(\lambda,O):=\{P\in\hat{L}_{\lambda}\setminus O:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\}$
is diffeomorphic to a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$. We will choose the neighborhood $O$ so that the dimension of this set is bigger than the quantity $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ for small $\lambda$. More specifically, the neighborhood $O$ is chosen in a way so that for a sufficiently small range of $\lambda$, the set $C(\lambda,O)$ contains a Cantor set of thickness larger than $M$.
In the next section, we introduce a series of constants that will be used throughout the proof, and these constants will determine the size of the neighborhoods that make up $O$.
3.1 Development of Constants
We define a series of constants needed in the proof. Their significance will become clear as the proof progresses.
Recalling (2.2), we fix a cone $K:=K_{\beta_{1}}^{u}$, for some $\beta_{1}\in(\beta_{0},0.1)$, so that it strictly contains $K^{u}$ and is transversal to $K^{s}$. We also choose $\beta_{1}$ close enough to $\beta_{0}$ so that the property described in Remark 2 holds for the cone $K$. If $K(x,y)$ denotes the cone $K$ in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, then we denote $\mathcal{K}:=\left\{K(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}$. We set
$C^{*}:=\frac{C^{*}_{0}}{C_{1}},$ (3.2)
where the constants $C_{1}$ and $C^{*}_{0}$ are taken from Remarks 3 and 4, respectively. Referring to Lemma 2, we set
$\mu_{1}:=\overline{\mu}_{\beta_{0}},\ \text{and}\ \mu_{2}:=\max_{n}\left\{\mu^{u}(a_{n})\right\}.$
Furthermore, choose a $\delta>0$ such that
$\mu_{1}-\delta>1,$
and choose $C_{2}>0$ such that
$\|\hat{T}_{0,a}\|_{C^{2}}\leq C_{2}$
for all $a\in\mathcal{A}$.
We say that a $C^{1}$ diffeomorphism $T:\mathbb{T}^{2}\rightarrow\mathbb{T}^{2}$, satisfies Property (C) if for each $(x,y)\in\mathbb{T}^{2}$, we have
  • $D_{(x,y)}T^{-1}\left(K^{s}(x,y)\right)\subset\text{int}K^{s}\left(T(x,y)\right)$ and $D_{(x,y)}T\left(K^{u}(x,y)\right)\subset\text{int}K^{u}\left(T(x,y)\right)$
  • $\left(\mu_{1}-\delta\right)|v|\leq\left|D_{(x,y)}T^{-1}v\right|\leq\left(\mu_{2}+\delta\right)|v|$ for $v\in K^{s}(x,y)$
  • $\left(\mu_{1}-\delta\right)|v|\leq\left|D_{(x,y)}Tv\right|\leq\left(\mu_{2}+\delta\right)|v|$ for $v\in K^{u}(x,y)$
Certainly, all of our maps $\left\{\hat{T}_{0,a_{n}}\right\}$ satisfy Property (C) by Lemma 2. We set
$C_{3}:=\sum_{n=-1}^{\infty}\frac{1}{(\mu_{1}-\delta)^{-(n+1)}},$ (3.3)
Remark 6.
There is a $\Theta>0$ such that if $(T_{n})$ is a sequence of maps satisfying Property (C) and $\sup_{n}\|T_{n}\|_{C^{2}}<C_{2}+1$, and if $\gamma$ is a smooth curve tangent to $\mathcal{K}^{u}$ so that $\sum_{i=0}^{n}\text{length}\left(T_{n}\circ\cdots\circ T_{1}(\gamma)\right)<C_{1}C_{3}\frac{\sqrt{2}}{10}$ for each $n$, then $\sum_{i=0}^{n}\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)<\Theta$ for each $n$, where $\text{Ang}(\kappa)$ denotes the supremum of the angles between tangent vectors of $\kappa$.
Let us define
$C_{4}:=\exp{\left[(C_{2}+1)^{2}\left(\Theta+C_{1}C_{3}\frac{\sqrt{2}}{10}\right)\right]}.$ (3.4)
By Lemma 4 we have
Lemma 6.
There is a $l^{*}>0$ such that if $L$ is a line through $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then for each $i=1,\dots,4$, there are segments $L^{+}_{i},\ L^{-}_{i}\subset L$ contained in $\mathbb{B}_{i}\setminus\partial^{s}\mathbb{B}_{i}$ such that
  • 1.)
    $L^{\pm}_{i}$ connects the opposite lines that make up $\partial^{u}\mathbb{B}_{i}$ and are centered at $Q_{i}$ with respect to $\tilde{K}(Q_{i})$, where $L^{\pm}_{i}\cap\tilde{K}_{\pm}(Q_{i})\neq\emptyset$
  • 2.)
    For any $P\in\left(L^{+}_{i}\cup L^{-}_{i}\right)\cap\tilde{K}(Q_{i})$, we have $|P-Q_{i}|<\cos(\theta)\frac{C^{*}}{4}$
where $\theta$ is the angle between boundary lines of $\tilde{K}(Q_{i})$ (or $K$).
Let us choose a $\tilde{n}$ so that $\frac{\left(\mu_{1}-\delta\right)^{\tilde{n}}}{10}>l^{*}$ and set $l^{**}:=\left(\mu_{2}+\delta\right)^{\tilde{n}+1}$. Then, again by Lemma 4, we have
Lemma 7.
There is a $C^{**}\in\left(0,\frac{1}{4}-\frac{1}{10}\right)$ such that if $L$ is a line through $Q_{0}$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\leq l^{**}$, then
$\text{dist}(L,Q_{i})\geq C^{**},$ (3.5)
for each $i=1,\dots,4$.
Recalling our constant $C_{1}$ from Remark 3, we will choose $N^{*}\in\mathbb{N}$ such that
$C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-N^{*}}}-C_{4}^{-1}\\ \geq M$ (3.6)
and
$C_{1}(\mu_{1}-\delta)^{-N^{*}}C^{*}\leq\frac{C^{**}}{4}.$ (3.7)
3.2 Translation to a Toral Map Problem
We are now ready to choose the neighborhood $O$ that was described at the beginning of this section.
From Remark 5, there is a $r_{0}=r_{0}\left(N^{*},C^{**},K\right)$ such that
$r_{0}<(\mu_{2}+\delta)^{-N^{*}-1}\frac{C^{**}}{4},$ (3.8)
and for each $Q\in\mathbb{T}^{2}$ and any rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$ satisfying
[…]
  • $\text{dist}\left(\partial^{s}R,Q\right)\geq(\mu_{2}+\delta)^{-N^{*}-1}\frac{C^{**}}{2}$
  • $\partial^{ns}R$ is a union of two disjoint lines with endpoints belonging to $\partial\tilde{K}(Q)$
we must have that $B_{r_{0}}(Q)\subset R$ and if $\gamma$ is a curve tangent to $\mathcal{K}^{u}$ that connects +$\partial^{s}_{+}R$ and $\partial^{s}_{-}R$ and satisfies $\gamma\cap B_{r_{0}}(Q)\neq\emptyset$, then $\gamma$ is completely contained in $R$ and does not intersect $\partial^{ns}R$.
Referring to Lemma 5, let us choose $\rho>0$ so small that $r(\rho)<\frac{r_{0}}{2}$ and set $\lambda_{1}:=\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$. We see that for $\lambda\in[0,\lambda_{1}]$, the set
$\left\{\boldsymbol{x}\in L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}:T_{a_{n}}\circ\cdots\circ T_{a_{1}}(\boldsymbol{x})\not\in O_{\rho}\ \text{for all}\ n>0\right\}$ (3.9)
is a one-dimensional compact set that is contained in $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$.
It is a straightforward calculation to check that the line $\hat{L}_{0}:=\left[0,\frac{1}{2}\right]\times\{0\}\subset\mathbb{T}^{2}$ is mapped into $L_{0}\cap\mathbb{S}$ under the map $F$. Also, the projection map $\pi_{\lambda}$ sends $L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}$ onto a curve in $\mathbb{S}$ that is $C^{2}$ close to a segment of $L_{0}$, provided $\lambda$ is small. That is, for small $\lambda$, there is a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O_{r(\rho)}$, that is $C^{2}$ close to a segment of $\hat{L}_{0}$, that is mapped into $L_{\lambda}$ under the map $F_{\lambda}$, and this curve tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ as $\lambda\rightarrow 0$.
As the map $F_{\lambda}$ preserves Hausdorff dimension, it suffices to estimate the Hausdorff dimension of the set
$\mathcal{C}\left(\lambda,O_{r(\rho)}\right):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O_{r(\rho)}\ \text{for all}\ n>0\right\}.$ (3.10)
That is, it suffices to show that there is a $\lambda_{0}\in(0,\lambda_{1})$ such that
$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,O_{r(\rho)}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (3.11)
for all $\lambda\in[0,\lambda_{0}]$.
3.3 Extended Torus Maps
Now, for each $a\in\mathcal{A}$ and $\lambda\in[0,\lambda_{1}]$, we will construct an Anosov diffeomorphism that coincides with $\tilde{T}_{\lambda,a}$ outside of $O_{r_{0}}$.
Lemma 8.
For each $a\in\mathcal{A}$ and $\lambda\in\left[0,\lambda_{1}\right]$, there is a smooth diffeomorphism $\hat{T}_{\lambda,a}:\mathbb{T}^{2}\rightarrow\mathbb{T}^{2}$, such that $\hat{T}_{\lambda,a}=\tilde{T}_{\lambda,a}\restriction_{\mathbb{T}^{2}\setminus O_{r_{0}}}$ and $\lim\limits_{\lambda\rightarrow 0}\left\|\hat{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}=0$ on $\mathbb{T}^{2}$.
Proof.
Let $\psi:\mathbb{T}^{2}\rightarrow[0,1]$ be a smooth function such that $\psi=0$ on $\mathbb{T}^{2}\setminus O_{r_{0}}$ and $\psi=1$ on $O_{\frac{r_{0}}{2}}$. Then, as $\tilde{O}_{r(\rho)}\subset O_{\frac{r_{0}}{2}}$, we must have that $\tilde{T}_{\lambda,a}$ is defined on $\mathbb{T}^{2}\setminus O_{\frac{r_{0}}{2}}$.
Writing in components
$\tilde{T}_{0,a}(x,y)=\left(\tilde{T}_{0,a}^{1}(x,y),\tilde{T}_{0,a}^{2}(x,y)\right)\ \text{and}\ \tilde{T}_{\lambda,a}(x,y)=\left(\tilde{T}_{\lambda,a}^{1}(x,y),\tilde{T}_{\lambda,a}^{2}(x,y)\right)$
where $\tilde{T}_{0,a}^{1}(x,y)=ax+y$, $\tilde{T}_{0,a}^{2}(x,y)=x$, and $\tilde{T}_{\lambda,a}^{1}(x,y),\ \tilde{T}_{\lambda,a}^{2}(x,y):\mathbb{T}^{2}\setminus O_{r(\rho)}\rightarrow S^{1}$, we set
$\hat{T}_{\lambda,a}(x,y):=\left(\hat{T}_{\lambda,a}^{1}(x,y),\hat{T}_{\lambda,a}^{2}(x,y)\right),$
where
$\hat{T}_{\lambda,a}^{1}(x,y):=\psi(x,y)\tilde{T}_{0,a}^{1}(x,y)+(1-\psi(x,y))\tilde{T}_{\lambda,a}^{1}(x,y)$
and
$\hat{T}_{\lambda,a}^{2}(x,y):=\psi(x,y)\tilde{T}_{0,a}^{2}(x,y)+(1-\psi(x,y))\tilde{T}_{\lambda,a}^{2}(x,y).$
The result follows from computing all partial derivatives up to order 2 of $\hat{T}_{\lambda,a}$, and comparing them to those of $\hat{T}_{0,a}$. ∎
To ease notation, we will denote $f_{\lambda,n,m}:=\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}$, for $n,m\geq 0$, where $\hat{T}_{\lambda,0}:=\text{id}$.
Remark 7.
Based on the compositions from (1.2) and construction of $\pi_{\lambda}$ in Lemma 5, one can check that for any $\lambda\in[0,\lambda_{1}]$, the set
$\left\{\left(T_{a_{n}}\circ\cdots\circ T_{a_{1}}\right)\left(-\sqrt{1+\frac{\lambda^{2}}{4}},0,0\right):n\in\mathbb{N}\right\}$
is finite and
$\left(\pi_{\lambda}\circ T_{a_{n}}\circ\cdots\circ T_{a_{1}}\right)\left(-\sqrt{1+\frac{\lambda^{2}}{4}},0,0\right)=T_{a_{n}}\circ\cdots\circ T_{a_{1}}(-1,0,0).$
Since $F\left(\frac{1}{4},\frac{1}{4}\right)=(-1,0,0)$, and $\left(\frac{1}{4},\frac{1}{4}\right)\in\mathbb{T}^{2}\setminus O_{r_{0}}$ (see 3.8), this gives
$\hat{T}_{\lambda,a_{n}}\circ\cdots\circ\hat{T}_{\lambda,a_{1}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n}}\circ\cdots\circ\hat{T}_{0,a_{1}}\left(\frac{1}{4},\frac{1}{4}\right)$
for any $n\in\mathbb{N}$, and this non-stationary orbit is a finite set. It is similar to check that
$\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)$
for any $m\geq 0$ and $n\in\mathbb{N}$.
Now, one may choose a $\lambda_{2}\in(0,\lambda_{1})$ such that for any $\lambda\in[0,\lambda_{2}]$ and $n\in\mathbb{N}$, we have
  • The map $\hat{T}_{\lambda,a_{n}}$ satisfies Property (C)
  • We have the estimate
    $\left\|\hat{T}_{\lambda,a_{n}}\right\|_{C^{2}}<C_{2}+1$ (3.12)
  • For any $P\in\mathbb{T}^{2}$, we have
    $D_{P}\hat{T}_{\lambda,a_{n}}\left(K_{\bullet}(P)\right)\subset K_{\bullet}\left(\hat{T}_{\lambda,a_{n}}(P)\right)$ (3.13)
[…]
  • The smooth curve $\hat{L}_{\lambda}$ is tangent to $\mathcal{K}^{u}$
The first and second part can be taken as a consequence of Lemma 8. The third part comes from the fact that $A_{a}$ has a positive unstable eigenvalue for each $a$, so if $\hat{T}_{\lambda,a}$ is sufficiently $C^{1}$-close to $\hat{T}_{0,a}$, then the differential of $\hat{T}_{\lambda,a}$ will preserve orientation of vectors from the cone $K$. The last part comes from the fact that the curve $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ in the $C^{2}$-sense.
From these cone conditions, we have
Lemma 9.
For all $\lambda\in[0,\lambda_{2}]$, $m\in\mathbb{N}$ and $n\geq 0$, we have
  • 1.)
    If $\gamma$ is a curve tangent to $\mathcal{K}^{s}$, then
    $(\mu_{1}-\delta)^{n+1}\text{length}(\gamma)\leq\text{length}\left(f_{\lambda,n,m}^{-1}(\gamma)\right)\leq(\mu_{2}+\delta)^{n+1}\text{length}(\gamma)$
  • 2.)
    If $\gamma$ is a curve tangent to $\mathcal{K}^{u}$, then
    $(\mu_{1}-\delta)^{n+1}\text{length}(\gamma)\leq\text{length}\left(f_{\lambda,n,m}(\gamma)\right)\leq(\mu_{2}+\delta)^{n+1}\text{length}(\gamma)$
Remark 8.
Suppose that $\kappa$ is a curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, for some $Q\in\mathbb{T}^{2}$, that is tangent to $\mathcal{K}$, with one of its endpoints being $Q$. Then, we must have that $\kappa\subset\tilde{K}_{\bullet}(Q)$ for =$\bullet$=+∙ = + or --. If we parameterize $\kappa$ via $\kappa(t):[0,1]\rightarrow\mathbb{T}^{2}$ so that it has positive orientation, then for each $t\in[0,1]$, we would have $\dot{\kappa}(t)\in K_{+}\left(\kappa(t)\right)$, and hence (3.13) implies that $f_{\lambda,n,m}(\kappa)\subset\tilde{K}_{\bullet}\left(f_{\lambda,n,m}(Q)\right)$, where one of its endpoints is $f_{\lambda,n,m}(Q)$. In other words, the orientation of curves tangent to $\mathcal{K}$ is preserved under the map $f_{\lambda,n,m}$.
3.3.1 Bounded Distortion Estimate
If $P$ and $Q$ are points on a curve $\gamma$, we denote the length of the segment of $\gamma$ from $P$ to $Q$ by $\text{dist}_{\gamma}(P,Q)$. Recalling (3.4), from Theorem 2 of [BM], we obtain the following bounded distortion estimate:
Lemma 10.
For each $\lambda\in[0,\lambda_{2}]$, if $\gamma$ is a sub-curve of $\hat{L}_{\lambda}$ that has endpoints $P$ and $Q$ and satisfies $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$ for some $i=1,\dots,4$, then for any point $Q^{\prime}\in\gamma\setminus\{P,Q\}$, we must have
$C_{4}^{-1}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})}$ $\leq\frac{\text{dist}_{f_{\lambda,m,0}(\gamma)}\left(f_{\lambda,m,0}(P),f_{\lambda,m,0}(Q^{\prime})\right)}{\text{dist}_{f_{\lambda,m,0}(\gamma)}\left(f_{\lambda,m,0}(Q),f_{\lambda,m,0}(Q^{\prime})\right)}$
$\leq C_{4}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})},$
for any $m\geq 0$.
Proof.
Let $\lambda\in[0,\lambda_{2}]$ and $m\geq 0$. Since $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$, it must have length less than $C_{1}\frac{\sqrt{2}}{10}$. Thus, from Lemma 9 and recalling (3.3), it must be the case that for each $k$ with $0\leq k\leq m$, we have
$\text{length}\left(f_{\lambda,k,0}(\gamma)\right)\leq(\mu_{1}-\delta)^{-(m-k)}\text{length}(f_{\lambda,m,0}(\gamma))\leq C_{1}\frac{\sqrt{2}}{10}(\mu_{1}-\delta)^{-(m-k)},$
so that
$\sum_{k=0}^{m}\text{length}(f_{\lambda,k,0}(\gamma))\leq C_{1}\frac{\sqrt{2}}{10}\sum_{k=0}^{m}(\mu_{1}-\delta)^{-(m-k)}\leq C_{1}C_{3}\frac{\sqrt{2}}{10}.$
In addition, since we have that $\hat{L}_{\lambda}$ is a smooth curve tangent to $\mathcal{K}^{u}$ and $\left\|\hat{T}_{\lambda,a_{n}}\right\|\leq C_{2}+1$ for all $n\in\mathbb{N}$, it follows from Remark 6 that
$\sum_{k=0}^{m}\text{Ang}\left(f_{\lambda,k,0}\right)<\Theta.$
Thus, using Theorem 2 of [BM], the result holds for $C_{4}$ as defined in (3.4). ∎
3.4 Non-Stationary Stable Manifolds
In this section, we establish a non-stationary version of the Stable Manifold Theorem for our collection of maps.
Consider the set
$W^{s}_{\lambda}\left(Q,m\right):=\left\{P\in\mathbb{T}^{2}:\lim\limits_{n\rightarrow\infty}\text{dist}\left(f_{\lambda,n,m}(P),f_{\lambda,n,m}(Q)\right)=0\right\},$
which we call the non-stationary stable manifold of $\left(T_{a_{n+m}}\right)_{n\geq 1}$ at $Q$. For $\eta>0$, we define
$W^{s}_{\lambda,\text{loc}}(Q,m):=\{P\in B_{\eta}(Q):f_{\lambda,n,m}(P)\in B_{\eta}(f_{\lambda,n,m}(Q))\ \text{for all}\ n\geq 0\}.$
We will denote
$\text{Orb}_{\lambda,m}(Q):=\{f_{\lambda,n,m}(Q)\}_{n\geq 0}\cup\{Q\}\ \text{and}\ \text{Orb}_{\lambda}(Q):=\bigcup_{m=1}^{\infty}\text{Orb}_{\lambda,m}(Q).$ (3.14)
Theorem 3.
There is a $\lambda_{3}\in(0,\lambda_{2})$ such that for all $\lambda\in[0,\lambda_{3}]$ and $\eta=\frac{1}{10}$, we have
  • 1.)
    For all $Q\in\mathbb{T}^{2}$ and $n\in\mathbb{N}$, we have that $W^{s}_{\lambda,\text{loc}}(Q,n)$ is a $C^{1}$ curve that is tangent to $\mathcal{K}^{s}$ and satisfies
    $\hat{T}_{\lambda,a_{n}}\left(W^{s}_{\lambda,\text{loc}}(Q,n)\right)\supset W^{s}_{\lambda,\text{loc}}\left(\hat{T}_{\lambda,a_{n}}(Q),n+1\right).$
  • 2.)
    For each $\epsilon>0$, there is a $\lambda^{\prime}(\epsilon)>0$ such that if $\lambda<\lambda^{\prime}(\epsilon)$, then
    $\text{dist}\left(W^{s}_{0,\text{loc}}(Q,k),W^{s}_{\lambda,\text{loc}}(Q,k)\right)<\epsilon$
    for all $k\in\mathbb{N}$ and any point $Q$ satisfying $f_{\lambda,n,k}(Q)=f_{0,n,k}(Q)$ for any $n,k\geq 0$ and each $\lambda$.
  • 3.)
    We have that $W^{s}(Q,k)=\bigcup_{n=0}^{\infty}f_{\lambda,n,k}^{-1}\left(W^{s}_{\lambda,\text{loc}}(f_{\lambda,n,k}(Q),k+n+1)\right)$ and there is a $C^{1}$ injective immersion $\iota_{\lambda,Q,k}:\mathbb{R}\rightarrow\mathbb{T}^{2}$ such that $\iota_{\lambda,Q,k}(\mathbb{R})=W^{s}_{\lambda}(Q,k)$ and $TW^{s}_{\lambda}(Q,k)\subset\mathcal{K}^{s}$.
Parts (1) and (3) of this Theorem are very close to a Corollary of Theorem 6.2.8 from [KH].
We will sketch the proof of these items using similar techniques from [KH], in our context.
Sketch of Proof of Theorem 3..
We prove the result first for $Q=0$. Recalling our notations from Section 2.3.1, we work locally in the neighborhood $\mathbb{B}_{1}=\pi\left(\text{Box}_{\frac{1}{10}}(0)\right)$ and in the coordinate system $\tilde{V}^{\perp}\oplus\tilde{V}$ where $\tilde{V}=\pi\left(V\cap\text{Box}_{\frac{1}{10}}(0)\right)$
Notice that there is a $l<1$ such that $\varphi:\tilde{V}^{\perp}\rightarrow\tilde{V}$ is tangent to $\mathcal{K}^{s}$ if and only if $\text{Lip}(\varphi)<l$.
We set
$X:=\left\{\text{graph}(\varphi):\varphi\in\text{Lip}\left(\tilde{V}^{\perp},\tilde{V}\right),\ \text{Lip}(\varphi)<l,\ \varphi(0)=0\right\}$
which is complete under the metric $d_{X}$ via
$d_{X}\left(\text{graph}(\varphi),\text{graph}(\varphi^{\prime})\right):=d_{C^{0}}(\varphi,\varphi^{\prime}).$
From Remark 2 and the cone conditions, we may choose a $\lambda_{3}>0$ such that for each $\lambda\in[0,\lambda_{3}]$ and each $a\in\mathcal{A}$, if $\text{graph}(\varphi)\in X$, then its image under $\hat{T}_{\lambda,a}^{-1}$ must overflow $\mathbb{B}_{1}$ and be tangent to $\mathcal{K}^{s}$.
For each $\lambda\in[0,\lambda_{3}]$ and $n\in\mathbb{N}$, let us define $\Gamma_{\lambda,a_{n}}:X\rightarrow X$ so that $\Gamma_{\lambda,a_{n}}\left(\text{graph}(\varphi)\right)$ is the segment of $\hat{T}_{\lambda,a_{n}}^{-1}\left(\text{graph}(\varphi)\right)\cap\mathbb{B}_{1}$ that contains 00. By the previous discussion, this map is well-defined.
Through the common cone conditions (Property (C)), for each $\lambda\in[0,\lambda_{3}]$, we find that $\left(\Gamma_{\lambda,a_{n}}\right)$ is a sequence of contractions with uniform contraction rate over $n$. Thus, by Proposition 1, for each $k\in\mathbb{N}$, there is a $\varphi_{\lambda,k}^{*}:V^{\perp}\rightarrow V$ such that $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)\in X$ and
$\lim\limits_{n\rightarrow\infty}\Gamma_{\lambda,a_{k}}\circ\cdots\circ\Gamma_{\lambda,a_{n+k}}\text{graph}\left(\varphi\right)=\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$
for any $\text{graph}\left(\varphi\right)\in X$. By construction, it follows that
$W^{s}_{\lambda,\text{loc}}\left(0,k\right)=\text{graph}\left(\varphi^{*}_{\lambda,k}\right).$
Again, through the cone conditions and $C^{1}$ smoothness of our maps, one finds that $\varphi_{\lambda,k}^{*}$ is differentiable and that
$T_{P}\text{graph}\left(\varphi_{\lambda,k}^{*}\right)=\bigcap_{k=1}^{\infty}D_{f_{\lambda,n,k}(P)}\left(f_{\lambda,n,k}\right)^{-1}\left[K^{s}\left(f_{\lambda,n,k}(P)\right)\right],$
for each $P\in\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ where the resulting set is a one-dimensional line in $K^{s}(P)$, and these lines vary continuously in $P$ (see Proposition 6.2.12 and Lemma 6.2.15 from [KH]). Hence, $\varphi_{\lambda,k}^{*}$ must be $C^{1}$ with its graph being tangent to the cone field $\mathcal{K}^{s}$. This establishes (1) and also (3).
To establish (2), we notice that the distance between $\Gamma_{0,n}$ and $\Gamma_{\lambda,n}$ in $(X,d_{X})$, can be made arbitrarily small for a sufficiently small range of $\lambda$ due to the fact that $\hat{T}_{\lambda,a_{n}}$ is a $C^{2}$ perturbation of $\hat{T}_{0,a_{n}}$, and this smallness can be made uniform in $n$.
Hence, by Proposition 1, we can make the corresponding fixed points $\text{graph}\left(\varphi_{0,k}^{*}\right)$ and $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ arbitrary close in the metric $d_{X}$ for sufficiently small $\lambda$, and this closeness can be made uniform in $k$.
For the case $Q\neq 0$, we repeat the process but for the family of maps given by
$P\mapsto\hat{T}_{\lambda,a_{n}}\left(P+f_{\lambda,n-1,0}(Q)\right)-f_{\lambda,n-1,0}(Q).$
Notice that in the case $\lambda=0$, we have $T_{Q}W^{s}_{0}(Q,m)=E^{s}_{[a_{m},a_{m+1},\dots]}$, but since $W^{s}_{0}(Q,m)$ is a line, from Lemma 3, we deduce that it is precisely the line through $Q$ of irrational slope given by $\frac{-1}{[a_{m},a_{m+1},\dots]}.$
Remark 9.
One could claim higher regularity of $W^{s}_{\lambda,\text{loc}}(Q,k)$ and in fact show that these curves are smooth or even real analytic, but we only provide the statement we actually need for the proof.
The next Corollary expands on (2) from the previous theorem.
Corollary 1.
Suppose that $Q_{0}\in\mathbb{T}^{2}$ is a point such that $f_{\lambda,n,k}(Q_{0})=f_{0,n,k}(Q_{0})$ for any $n,k\geq 0$ and each $\lambda>0$, and that $\text{Orb}_{0}(Q_{0})$ is finite. For each $\epsilon>0$ and $n_{0}\in\mathbb{N}$, there is a $\lambda^{\prime}\in\left(0,\lambda_{3}\right)$ such that if $\lambda\in\left[0,\lambda^{\prime}\right]$, we have that
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq n_{0}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)<\epsilon,$
for all $Q\in\text{Orb}_{\lambda}(Q_{0})$.
Proof.
Since for each $\lambda$, the collection of maps $\bigcup_{n=1}^{n_{0}}\bigcup_{m\in\mathbb{N}}\left\{f^{-1}_{\lambda,n,m}\right\}$ is finite, it is equicontinuous. In addition, since the map $g\mapsto g^{-1}$ is continuous over the space $\text{Diff}^{2}\left(\mathbb{T}^{2}\right)$, since each map $\hat{T}_{\lambda,a_{n}}$ is a $C^{2}$ perturbation of $\hat{T}_{0,a_{n}}$, and again since our collection of maps is finite, we can make $f_{\lambda,n,m}^{-1}$ and $f_{0,n,m}^{-1}$ arbitrarily close in the $C^{0}$ norm, where the closeness is uniform in $m\in\mathbb{N}$ and $n\leq n_{0}$, for a sufficiently small range of $\lambda$. Together with (2) of Theorem 3, the result follows. ∎
We note that while the result above is true for arbitrary $k$, the sets only have dynamical meaning for certain values of $k$. What we notice is that starting with any $m$, we have the nested sets
$W_{\lambda,\text{loc}}^{s}(Q,m)$
$\subset\hat{T}_{\lambda,a_{m}}^{-1}W_{\lambda,\text{loc}}^{s}\left(\hat{T}_{\lambda,a_{m}}(Q),m+1\right)=f_{\lambda,0,m}^{-1}W^{s}_{\lambda,\text{loc}}(f_{\lambda,0,m}(Q),m+1)$
$\subset\hat{T}_{\lambda,a_{m}}^{-1}\circ\hat{T}_{\lambda,a_{m+1}}^{-1}W_{\lambda,\text{loc}}^{s}\left(\hat{T}_{\lambda,a_{m+1}}\circ\hat{T}_{a_{m}}^{-1}(Q),m+2\right)=f^{-1}_{\lambda,1,m}W^{s}(f_{\lambda,1,m}(Q),m+2)$
[…]
$\subset\hat{T}_{\lambda,a_{m}}^{-1}\circ\cdots\circ\hat{T}_{\lambda,a_{m+n_{0}}}^{-1}W_{\lambda,\text{loc}}^{s}\left(\hat{T}_{\lambda,a_{m+n_{0}}}\circ\cdots\circ\hat{T}_{\lambda,a_{m+1}}(Q),m+n_{0}+1\right)$
$=f^{-1}_{\lambda,n_{0},m}W^{s}(f_{\lambda,\tilde{n},m}(Q),m+n_{0}+1)$
[…]
$\subset W^{s}_{\lambda}(Q,m)$
and Corollary 1 gives that for a small range of $\lambda$, depending on $n_{0}$, these sets are $\epsilon$-close to the corresponding sets for $\lambda=0$, and this is true for any $Q\in\text{Orb}(Q_{0})$. That is, arbitrarily long pieces of the stable manifolds $W^{s}_{\lambda}(Q,m)$ containing $Q$ can be made arbitrarily close to pieces of $W^{s}_{0}(Q,m)$ containing $Q$, uniformly in $m$, provided the pieces are of uniform size, for a sufficiently small range of $\lambda$.
3.5 Rectangle Construction
Recall that by Remark 7, for $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$, we have that
$f_{\lambda,n,m}(Q_{0})=f_{0,n,m}(Q_{0})$
for each $n,m\geq 0$ and $\lambda\in[0,\lambda_{2}]$, and it is also a calculation to see that $\text{Orb}_{0}(Q_{0})$ is a finite set and bounded away from each $Q_{i}$ by at least $\frac{1}{4}$.
Now, for each $i=1,\dots,4$, $m\in\mathbb{N}$, and $\lambda\in[0,\lambda_{3}]$, we will construct a family of rectangles containing $Q_{i}$ that have stable sides belonging to $W^{s}_{\lambda}(Q_{0},m)$ and also have controlled size.
The idea is that for each $i=1,\dots,4$ and $m\in\mathbb{N}$, we will find a rectangle $R$ that contains $B_{r_{0}}(Q_{i})$ and has stable sides belonging to $W^{s}_{\lambda}(Q_{0},m+1)$, and such that any segment from $\hat{L}_{\lambda}$ that intersects $B_{r_{0}}(Q_{i})$, under the map $f_{\lambda,m,0}$, must also intersect the stable sides of the rectangle $R$. This means that for future iterates, the points that intersected the stable boundaries of $R$ will move away from the points $Q_{i}$.
This dynamical behavior will allow us to construct a Cantor set contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has large thickness.
3.5.1 Base Set of Rectangles
The construction of our rectangles will be done in an inductive process, so we start by constructing a base set of rectangles.
Recalling the constant $C^{*}$ from (3.2) and $C^{**}$ from Lemma 7, we have
Lemma 11.
There is a $\lambda_{4}\in(0,\lambda_{3})$ such that for each $\lambda\in[0,\lambda_{4}]$, $m\in\mathbb{N}$, and $i=1,\dots,4$, there is a dynamical rectangle $R_{m}(\lambda,Q_{i})$ centered at $Q_{i}$ with respect to $\tilde{K}(Q_{i})$ such that
  • 1.)
    $\partial^{s}R_{m}(\lambda,Q_{i})\subset W^{s}_{\lambda}(Q_{0},m)$ and in particular, +$\partial^{s}_{+}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{+}(Q_{i})$ and $\partial^{s}_{-}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{-}(Q_{i})$
  • 2.)
    For any $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$
  • 3.)
    If $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, then
    $\min\left\{\text{dist}\left(P,\bigcup_{i=1}^{4}\{Q_{i}\}\right),\text{dist}\left(f_{\lambda,n,m}(P),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\right\}\geq\frac{C^{**}}{2}$
    for any $n\geq 0$
Proof.
We break the proof into steps:
Step 1: We first construct the stable sides of the rectangles for $\lambda=0.$
We know that $W^{s}_{0}(Q_{0},m)$ is a line through $Q_{0}$ of slope $-\frac{1}{[a_{m},a_{m+1},\dots]}$ and we know from Lemma 9 that
$\text{length}\left(f^{-1}_{0,\tilde{n},m}\left(W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),m+\tilde{n}+1)\right)\right)\geq l^{*},$
where $\tilde{n}$ was chosen right after Lemma 6.
From Lemma 6, we know that $f^{-1}_{0,\tilde{n},m}\left(W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),m+\tilde{n}+1)\right)$ has two segments, denoted by $\gamma^{+}_{i,m}$ and $\gamma^{-}_{i.m}$, that connect the opposite lines that make up $\partial^{u}\mathbb{B}_{i}$ and are centered at $Q_{i}$ with respect to $\tilde{K}(Q_{i})$, where $\gamma^{\pm}_{m,i}\cap\tilde{K}_{\pm}(Q_{i})\neq\emptyset$. In addition, these segments have distance from $Q_{i}$ less than $\cos(\theta)\frac{C^{*}}{4}$, where $\theta$ is the angle between boundary lines of the cone $\tilde{K}(Q_{i})$. We verify these segments satisfy certain estimates:
[…]
$|P-Q_{i}|\leq\cos(\theta)^{-1}|P^{\prime}-Q_{i}|<\frac{C^{*}}{4}.$ (3.15)
Suppose that $P\in\gamma^{+}_{i,m}\cup\gamma^{-}_{i,m}$ and let $n\geq 0$. From Lemma 7, we know $|P-Q_{i}|\geq C^{**}$ since
$\text{length}\left(f^{-1}_{0,\tilde{n},m}\left(W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),m+\tilde{n}+1)\right)\right)\leq\frac{(\mu_{2}+\delta)^{\tilde{n}+1}}{10}<l^{**},$
from Lemma 9.
Notice that $f_{0,\tilde{n},m}(P)\in W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)$, and so if $n\geq\tilde{n}$, then $f_{\lambda,n,m}(P)\in B_{\frac{1}{10}}(Q_{0})$ so that its distance from $\bigcup_{i=1}^{4}\{Q_{i}\}$ is greater than $\frac{3}{20}\frac{1}{4}-\frac{1}{10}$.
If $n<\tilde{n}$, then $f_{0,n,m}(P)\in f^{-1}_{0,\tilde{n}-n,m+n+1}W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),\tilde{n}+m+1)$, where $f^{-1}_{0,\tilde{n}-n,m+n+1}W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),\tilde{n}+m+1)$ is a line segment contained in $W^{s}_{0}(f_{0,n,m}(Q_{0}),m+n+1)$ that has slope $-\frac{1}{[a_{m+n+1},a_{m+n+2},\dots]}$, contains $f_{0,n,m}(Q_{0})\in\text{Orb}_{0,m}(Q_{0})$, and has length no greater than $\frac{(\mu_{2}+\delta)^{(\tilde{n}-n)+1}}{10}<l^{**}$. It follows from Lemma 7 that
$\min\left\{\text{dist}\left(P,\bigcup_{i=1}^{4}\{Q_{i}\}\right),\text{dist}\left(f_{0,n,m}(P),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\right\}\geq C^{**}$ (3.16)
for all $n\geq 0$.
Step 2: We now construct the stable sides of the rectangles for a small range of $\lambda$.
[…]
From Corollary 1, the quantity
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq\tilde{n}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)$ (3.17)
can be made arbitrarily small for a sufficiently small range of $\lambda$ for all $Q\in\text{Orb}_{\lambda}(Q_{0})$.
Then in particular, if $f^{-1}_{\lambda,\tilde{n},m}\left(W^{s}_{\lambda,\text{loc}}(Q_{0},m+\tilde{n}+1)\right)$ and $f^{-1}_{0,\tilde{n},m}\left(W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)\right)$ are close enough, there will be a curve $\gamma^{\bullet}_{\lambda,i,m}\subset f^{-1}_{\lambda,\tilde{n},m}\left(W^{s}_{\lambda,\text{loc}}(Q_{0},m+\tilde{n}+1)\right)$ that exhibits a sub-curve $\psi^{\bullet}_{\lambda,i,m}\subset\gamma^{\bullet}_{\lambda,i,m}$ such that $\psi^{\bullet}_{\lambda,i,m}$ connects opposite boundaries of $\tilde{K}_{\bullet}(Q_{i})$. Using estimates (3.15) and (3.16), we may find a $\lambda_{4}\in(0,\lambda_{3})$ such that for all $\lambda\in[0,\lambda_{4}]$, we have
  • For any $P\in\psi_{\lambda,i,m}^{+}\cup\psi_{\lambda,i,m}^{-}$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$
  • For any $P\in\psi_{\lambda,i,m}^{+}\cup\psi_{\lambda,i,m}^{-}$, we have
    $\min\left\{\text{dist}\left(P,\bigcup_{i=1}^{4}\{Q_{i}\}\right),\text{dist}\left(f_{\lambda,n,m}(P),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\right\}\geq\frac{C^{**}}{2}$
    for all $n\geq 0$
Step 3: We now construct the rectangles.
[…]
Figure 7: Rectangle from Lemma 11.
3.5.2 Rectangles of Controlled Size
We will now use this base set of rectangles to construct dynamical rectangles of small controlled size. We will need rectangles of smaller scale to ensure that the Cantor set of $\mathcal{C}(\lambda,O_{r(\rho)})$, that we construct later on, will have small gaps in comparison to their bridges.
The construction will come from taking the stable boundaries of rectangles from the collection $\left\{R_{\lambda}(m,Q_{i})\right\}$, and iterating them under the maps $f^{-1}_{\lambda,n,m}$, which will give two new curves tangent to $\mathcal{K}^{s}$ that have segments which are close together and centered at some $Q_{i}$, and these curves will serve as the stable boundaries for a new rectangle.
Lemma 12.
For each $\lambda\in[0,\lambda_{4}]$, $m\in\mathbb{N}$, $i=1,\dots,4$, and $N\in\mathbb{N}\cup\{0\}$, there is a dynamical rectangle $R_{m}^{\prime}(\lambda,Q_{i},N)$ centered at $Q_{i}$ with respect to $\tilde{K}(Q_{i})$ satisfying:
  • 1)
    The stable boundaries $\partial^{s}R_{m}^{\prime}(\lambda,Q_{i},N)$ satisfy the following:
[…]
    $\partial^{s}_{\bullet}R_{m}^{\prime}(\lambda,Q_{i},N)\subset f^{-1}_{\lambda,N,m}\left[\partial^{s}_{\bullet}R_{m+N+1}\left(\lambda,f_{\lambda,N,m}(Q_{i})\right)\right]$
    and in particular
    $\partial^{s}_{\bullet}R_{m}^{\prime}(\lambda,Q_{i},N)\subset\hat{T}_{\lambda,a_{m}}^{-1}\left(\partial^{s}_{\bullet}R^{\prime}_{m+1}\left(\lambda,\hat{T}_{\lambda,a_{m}}(Q_{i}),N-1\right)\right).$
  • 2)
    If $\kappa$ is a curve contained in $\mathbb{B}_{i}$ and tangent to $\mathcal{K}^{u}$ that connects $\partial_{+}^{s}R^{\prime}_{m}(\lambda,Q_{i},N)$ and $\partial_{-}^{s}R^{\prime}_{m}(\lambda,Q_{i},N)$, then
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}C^{*}$
[…]
    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}\frac{C^{*}}{2}$
  • 3)
    We have
    $\text{dist}\left(\partial_{\bullet}^{s}R^{\prime}_{m}(\lambda,Q_{i},N),Q_{i}\right)\geq(\mu_{2}+\delta)^{-N-1}\frac{C^{**}}{2},$
[…]
  • 4)
    The non-stable boundaries $\partial^{ns}R^{\prime}_{m}(\lambda,Q_{i},N)$ of $R^{\prime}_{m}(\lambda,Q_{i},N)$ are line segments with endpoints in $\partial\tilde{K}(Q_{i})$. That is, $R^{\prime}_{m}(\lambda,Q_{i},N)$ has vertices belonging to $\partial\tilde{K}(Q_{i})$, and does not intersect $\tilde{K}(Q_{i})$ otherwise.
Proof.
Let $\lambda\in[0,\lambda_{4}]$. For $N=0$, from Lemma 11, we set $R^{\prime}_{m}(\lambda,0,Q_{i})=R_{m}(\lambda,Q_{i})$, and these rectangles satisfy the desired properties.
We will proceed inductively. Let $N\geq 1$, and suppose that we have constructed the family of rectangles $\left\{R^{\prime}_{m}(\lambda,Q_{i},k):m\in\mathbb{N},i=1,\dots,4\right\}$ for each $k<N$ and that they satisfy the desired properties.
Let $m\in\mathbb{N}$ and $i\in\{1,\dots,4\}$. Consider the rectangle $R^{\prime}_{m+1}\left(\lambda,Q_{j},N-1\right)$, where $Q_{j}=f_{\lambda,0,m}(Q_{i})$, which is centered at $Q_{j}$ with respect to $\tilde{K}(Q_{j})$.
Notice from (3.13) and Remark 2, provided $\lambda$ is small, the image of the cone $\tilde{K}(Q_{i})$, under the map $f_{\lambda,0,m}$, overflows the box $\mathbb{B}_{j}$ (see Figure 8). Without loss of generality, let us assume that this holds for all $\lambda\in[0,\lambda_{4}].$
[…]
By construction and the induction hypothesis, we have
$\partial^{s}_{\bullet}R^{\prime}_{m}(\lambda,Q_{i},N)\subset\hat{T}_{\lambda,a_{m}}^{-1}\left(\partial^{s}_{\bullet}R^{\prime}_{m+1}(\lambda,Q_{j},N-1)\right)$
$\subset\hat{T}_{\lambda,a_{m}}^{-1}\left(f^{-1}_{\lambda,N-1,m+1}\left[\partial^{s}_{\bullet}R_{(m+1)+(N-1)+1}(\lambda,Q_{k})\right]\right)$
$=f^{-1}_{\lambda,N,m}\partial^{s}_{\bullet}R_{m+N+1}\left(\lambda,f_{\lambda,N,m}(Q_{i})\right).$
This establishes (1) and (4), and we now check that this rectangle satisfies (2) and (3).
[…]
$\text{length}(\kappa)\leq(\mu_{1}-\delta)^{-1}\text{length}\left(f_{\lambda,0,m}(\kappa)\right)\leq(\mu_{1}-\delta)^{-1}C_{1}(\mu_{1}-\delta)^{-(N-1)}\frac{C^{*}}{2}$
so that second part of (2) holds, and the first part is verified in an almost identical way.
In a similar manner, again from (2) of Lemma 9 and the induction hypothesis (in particular property (3)), we also have
$\text{length}(\kappa)$ $\geq(\mu_{2}+\delta)^{-1}\text{length}\left(f_{\lambda,0.m}(\kappa)\right)\geq(\mu_{2}+\delta)^{-1}\text{dist}\left(\partial^{s}_{\bullet}R^{\prime}_{m+1}(\lambda,Q_{j},N-1),Q_{j}\right)$
$\geq(\mu_{2}+\delta)^{-1}\frac{C^{**}}{2}(\mu_{2}+\delta)^{-N-2}=(\mu_{2}+\delta)^{-N-1}\frac{C^{**}}{2},$
and since $\partial^{s}_{\bullet}R^{\prime}_{m}(\lambda,Q_{i},N-1)$ is tangent to $\mathcal{K}^{s}$ and $\kappa$ was arbitrary and tangent to $\mathcal{K}^{u}$, we conclude that (3) holds. ∎
[…]
Figure 8: Illustration of the image of $\tilde{K}(Q_{i})$ under the map $f_{\lambda,0,m}$.
We note that combining (2) and (4) of this Lemma implies that
$\lim\limits_{N\rightarrow\infty}\text{diam}\left(R_{m}(\lambda,Q_{i},N)\right)=0,$ (3.18)
where this limit is uniform in $\lambda\in[0,\lambda_{4}]$, $m\in\mathbb{N}$, and $i=1,\dots,4$.
3.6 Gap Estimates and Thickness
In order to establish (3.11), we will show that for each $\lambda\in[0,\lambda_{4}]$, there is a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has thickness larger than $M$.
For each curve $\hat{L}_{\lambda}$, as it is close to a segment of $\hat{L}_{0}=\left[0,\frac{1}{2}\right]\times\{0\}$ and tangent to $\mathcal{K}^{u}$, for our range of $\lambda$, we assign a natural ordering to its elements as follows: Let $P^{*}$ be the endpoint of $\hat{L}_{\lambda}$ that is closest to $(0,0)$ and $P^{**}$ its other endpoint. We see that the endpoints will be close $(0,0)$ and $\left(\frac{1}{2},0\right)$, respectively, for small $\lambda$. We say that $P<P^{\prime}$ if
$\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{*}\right)<\text{dist}_{\hat{L}_{\lambda}}\left(P^{\prime},P^{*}\right).$
With this in mind, we have the following lemma.
Lemma 13.
Let $I\subset\hat{L}_{\lambda}$ be a segment with endpoints $P<P^{\prime}$ such that $f_{\lambda,n,m}(P)\in\tilde{K}_{+}(Q_{i})$ and $f_{\lambda,n,m}\left(P^{\prime}\right)\in\tilde{K}_{-}(Q_{j})$ for some $i,j\in\{1,\dots,4\}$. Then, there must be a $P_{0}\in\hat{L}_{\lambda}$ between $P$ and $P^{\prime}$ such that $f_{\lambda,n,m}(P_{0})\not\in\bigcup_{i=1}^{4}\mathbb{B}_{i}$.
Proof.
If $i\neq j$, then this is clear. For $i=j$, if we parametrize $I$ as a curve with positive orientation, then this follows from Remark 8. ∎
3.6.1 Gap Constructions
Recall from (3.7) we know $R^{\prime}_{k}(\lambda,Q_{i},N^{*})\subset B_{\frac{C^{**}}{4}}(Q_{i})$ for all $i=1,\dots,4$ and $k\in\mathbb{N}$.
We will construct a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ that is contained in $\mathcal{C}\left({\lambda},O_{r(\rho)}\right)$ and show that it has thickness bigger than $M$:
Consider the line $\hat{L}_{\lambda}$ and recall that it is tangent to $\mathcal{K}^{u}$ for any $\lambda\in[0,\lambda_{4}]$. As $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus O_{r_{0}}$ as $\lambda\rightarrow 0$, and $B_{r_{0}}(0,0)\subset R_{1}(\lambda,(0,0))$ and $B_{r_{0}}\left(\frac{1}{2},0\right)\subset R_{1}\left(\lambda,\left(\frac{1}{2},0\right)\right)$, we can work in a small enough range of $\lambda$ so that $\hat{L}_{\lambda}$ has its least and greatest endpoints belonging to $R_{1}(\lambda,(0,0))$ and $R_{1}\left(\lambda,\left(\frac{1}{2},0\right)\right)$ respectively. If $\lambda$ is sufficiently small, then we can assume that $\hat{L}_{\lambda}$ intersects the stable sides of these rectangles since this is the case for $\hat{L}_{0}$. Without loss of generality, let us assume this holds for all $\lambda\in[0,\lambda_{4}]$.
Then, there exist $P_{0}^{+},P_{0}^{-}\in\hat{L}_{\lambda}$ such that $P_{0}^{+}<P_{0}^{-}$ and $P_{0}^{+}\in\partial^{s}_{+}R_{1}(\lambda,(0,0))$ and $P_{0}^{-}\in\partial^{s}_{-}R_{1}\left(\lambda,\left(\frac{1}{2},0\right)\right)$.
We notice that
$\text{dist}\left(f_{\lambda,k,0}\left(P_{0}^{\bullet}\right),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\geq\frac{C^{**}}{2}$ (3.19)
[…]
Suppose that $I_{0}$ is the segment from $P_{0}^{+}$ to $P_{0}^{-}$ contained in $\hat{L}_{\lambda}$, and let $k_{0}$ be the smallest value such that $f_{\lambda,k_{0},0}\left(I_{0}\right)$ intersect $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$. Now, we consider all of the components of
$f_{\lambda,k_{0},0}(I_{0})\cap\left(\bigcup_{i=1}^{4}R^{\prime}_{k_{0}+1}\left(\lambda,Q_{i},N^{*}\right)\right).$
The gaps of order $k_{0}$ will come from taking the image of such components that intersects $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$ non-trivially, under the map $f^{-1}_{\lambda,k_{0},0}$.
Let $J$ be such a component, and suppose it is contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ for some $j=1,\dots,4$, and $J\cap O_{r_{0}}(Q_{j})\neq\emptyset$. Then, $J$ is the image of some segment $I_{0}^{\prime}\subset I_{0}$ under the map $f_{\lambda,k_{0},0}$. We will say that $\text{int}I^{\prime}_{0}$ is a gap of order $k_{0}$, where the interior is taken with respect to the subspace topology on $\hat{L}_{\lambda}$.
Since $R^{\prime}_{k_{0}+1}\left(\lambda,Q_{j},N^{*}\right)$ is centered at $Q_{j}$ with respect to $\tilde{K}(Q_{j})$ and
$f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is a curve tangent to $\mathcal{K}^{u}$, by Remark 5 and our discussion after Lemma 12, as $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)\cap O_{r_{0}}(Q_{j})\neq\emptyset$, we must have that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is fully contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and only intersects $\partial R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ non-trivially in the event that either of its endpoints belong to $\partial^{s}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. However, from (3.19), and since we chose $N^{*}$ so large that $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})\subset B_{\frac{C^{**}}{4}}$ (see (3.7)), it must be the case that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ indeed connects the stable boundaries of $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$. In particular, by Lemma 13, if $I_{0}^{\prime}$ has endpoints $P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$, then we have $f_{\lambda,k_{0},0}(P_{1})\in\partial^{s}_{-}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and $f_{\lambda,k_{0},0}(P_{2})\in\partial^{s}_{+}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
Suppose that we have constructed all of the gaps up to order $k\geq k_{0}$, and that they are all disjoint segments of $\hat{L}_{\lambda}$ of the form $\text{int}I_{m}$ where $I_{m}\subset I_{0}$ for some $m\leq k$ such that:
  • $f_{\lambda,k,0}\left(I_{m}\right)$ is a curve contained in $R^{\prime}_{m+1}(\lambda,Q_{i},N^{*})$, for some $i=1,\dots,4$
  • $I_{m}$ has endpoints $P_{m}^{-}<P_{m}^{+}$ so that $f_{\lambda,m,0}\left(P_{m}^{-}\right)\in\partial^{s}_{-}R^{\prime}_{m+1}(\lambda,Q_{i},N^{*})$ and $f_{\lambda,m,0}\left(P_{m}^{+}\right)\in\partial^{s}_{+}R^{\prime}_{m+1}(\lambda,Q_{i},N^{*})$
  • $f_{\lambda,m,0}\left(I_{m}\right)\cap O_{r_{0}}(Q_{i})\neq\emptyset$
Denote this collection of gaps by $\mathcal{U}_{k}$. Let $I$ be a component of $I_{0}\setminus\mathcal{U}_{k}$. Now, consider the components of
$f_{\lambda,k+1,0}(I)\cap\left(\bigcup_{i=1}^{4}R^{\prime}_{k+2}(\lambda,Q_{i},N^{*})\right).$
The gaps of order $k+1$ will come from taking the image of such components that intersects $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$ non-trivially, under the map $f^{-1}_{\lambda,k+1,0}$. A similar application of Remark 5 and Lemma 13 as in the base case shows that any gap of order $k+1$ will be of the form above with index $m$ replaced with $k+1$.
We have now constructed all orders of gaps of $\mathcal{C}\left(\lambda,N^{*}\right)$ and shown they have satisfied a certain form. From (3.8) and definition of $r_{0}$, we know
$B_{r(\rho)}(Q_{i})\subset B_{r_{0}}(Q_{i})\subset R^{\prime}_{k}(\lambda,Q_{i},N^{*})$
for all $k\geq 0$, $i=1,\dots,4$ and $\lambda\in[0,\lambda_{4}]$. Thus, it follows that
$C\left(\lambda,N^{*}\right)\subset C\left(\lambda,O_{r(\rho)}\right).$ (3.20)
3.6.2 Thickness Estimate
Let $\lambda_{0}=\min\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ and let $\lambda\in[0,\lambda_{0}]$. We will show that $C\left(\lambda,N^{*}\right)$ has thickness larger than $M$.
With respect to the Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ and the ordering of gaps defined above, let $U\subset I_{0}$ be a gap of order $k$ with endpoints $P<P^{\prime}$, and $B$ a bridge of it. Without loss of generality, suppose that $P$ is an endpoint of $B$, and denote its other endpoint by $P^{\prime\prime}$. We know that $f_{\lambda,k,0}\left(P\right)\in\partial^{s}_{-}R^{\prime}_{k+1}(\lambda,Q_{i},N^{*})$ and $f_{\lambda,k,0}\left(P^{\prime}\right)\in\partial^{s}_{+}R^{\prime}_{k+1}(\lambda,Q_{i},N^{*})$, and there must be some $Q^{\prime}$ between $P$ and $P^{\prime}$ such that $f_{\lambda,k,0}\left(Q^{\prime}\right)\in O_{r_{0}}(Q_{i})$.
Lemma 14.
There is a point $Q\in B$ between $P$ and $P^{\prime\prime}$ such that
$\text{dist}\left(f_{\lambda,k,0}(Q),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\geq\frac{C^{**}}{2},$
and the segment $J\subset f_{\lambda,k,0}(B\cup U)$ connecting $f_{\lambda,0,k}(Q)$ and $f_{\lambda,k,0}(P)$ is contained in $\mathbb{B}_{i}$.
Proof.
By the construction of our gaps, the endpoint $P^{\prime\prime}$ of $B$ satisfies $f_{\lambda,k^{\prime},0}\left(P^{\prime\prime}\right)\in\partial^{s}R^{\prime}_{k^{\prime}+1}(\lambda,Q_{j},N^{*})$ for some $j=1,\dots,4$ and $k^{\prime}\leq k$.
From (1) of Lemma 12, either $f_{\lambda,k,0}\left(P^{\prime\prime}\right)\in\partial^{s}_{+}R_{k+1}(\lambda,Q_{j},N^{\prime})$ for some $N^{\prime}\leq N^{*}$ and $j=1,\dots,4$, or there was some $k^{\prime\prime}\in\left[k^{\prime},k\right)$ so that $f_{\lambda,k^{\prime\prime},0}\left(P^{\prime\prime}\right)\in\partial^{s}_{+}R_{k^{\prime\prime}+1}(\lambda,Q_{l},0)$ for some $l=1,\dots,4$.
In the former case, Lemma 13 implies that there is a point $Q^{*}\in B$ such that $f_{\lambda,k,0}(Q^{*})\not\in\mathbb{B}_{i}$. In this case, we notice that the segment of $f_{\lambda,k,0}(B\cup U)$ connecting $f_{\lambda,k,0}(Q^{*})$ and $f_{\lambda,k,0}(P^{\prime})$ intersects $O_{r_{0}}$, and from our choice of neighborhood $O_{r_{0}}$ (see Section 3.2), this segment must intersect $\partial^{s}_{-}R_{k+1}(\lambda,Q_{i})$, in which case, we choose $Q$ so that $f_{\lambda,k,0}(Q)$ is this intersection point, and we recall (3) of Lemma 11.
In the latter case, by (3) of Lemma 11, $f_{\lambda,k,0}(P^{\prime\prime})$ has distance from $Q_{i}$ greater than $\frac{C^{**}}{2}$. If the segment of $f_{\lambda,k,0}(B\cup U)$ connecting $f_{\lambda,k,0}(P^{\prime\prime})$ and $f_{\lambda,k,0}(P^{\prime})$ is not contained in $\mathbb{B}_{i}$, then by the same reasoning as in the former case, there is some $Q\in B$ such that $f_{\lambda,k,0}(Q)\in\partial^{s}_{-}R_{k+1}(\lambda,Q_{i})$. ∎
This Lemma gives the following estimate:
$\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}(Q)\right)\geq\left|f_{\lambda,k,0}(P)-f_{\lambda,k,0}(Q)\right|$
$\geq\left|f_{\lambda,k,0}(Q)-Q_{i}\right|-\left|Q_{i}-f_{\lambda,k,0}\left(Q^{\prime}\right)\right|-\left|f_{\lambda,k,0}(P)-f_{\lambda,k,0}\left(Q^{\prime}\right)\right|$
$\geq\frac{C^{**}}{2}-r_{0}-\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)$
$\geq\frac{C^{**}}{4}-\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right).$
where the last inequality follows from (3.8). Therefore, since the segment $J$ is contained in $\mathbb{B}_{i}$, by Lemma 10 and (2) of Lemma 12, we have
$\frac{\text{length}(B)}{\text{length}(U)}=\frac{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime\prime}\right)}{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime}\right)}\geq\frac{\text{dist}_{\hat{L}_{\lambda}}\left(P,Q\right)}{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime}\right)}$
$\geq C_{4}^{-1}\frac{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(Q\right)\right)}{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)}\geq C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-N}}-C_{4}^{-1}$
$\geq M,$
where the last inequality follows from (3.6). We then see that for each $\lambda\in[0,\lambda_{0}]$, we have
$\tau\left(C(\lambda,N^{*})\right)\geq M$
and from (1.3), we have
$\text{dim}_{H}\left(C(\lambda,N^{*})\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}.$
From (3.20), this then implies (3.11) holds, and hence (3.1) holds, for all $\lambda\in[0,\lambda_{0}]$.
Since $L_{\lambda}\cap\Lambda^{\text{bnd}}_{\lambda}$ is affine equivalent to $\sigma_{\lambda,\alpha}$ and $M$ was arbitrary, this proves Theorem 1.
Acknowledgements
I would like to thank Anton Gorodetski for his guidance and support. This project was supported by NSF grant DMS-2247966 (PI: A. Gorodetski).

### v2

@@
 […]
-This type of potential gives a self-adjoint operator that is directly associated to a mathematical model of quasicrystals.
-A survey of results concerning these types of operators, especially Schrödinger operators, can be found in [DEG].
+This operator falls under the category of a discrete Schrödinger operator with dynamically defined potential.
+This so-called Sturmian potential has been studied since the early 1980s, and has been a popular model in the study of one-dimensional quasicrystals.
+A Sturmian sequence is an aperiodic sequence of least possible complexity, and hence relates to the structure of quasicrystals which are highly ordered and aperiodic.
+A survey of results concerning these types of operators, especially Schrödinger operators, can be found in [DEG, DF2].
 We focus our attention to the case when the frequency has only finitely many different terms that appear in its continued fraction expansion (i.e.
@@
 $\lambda$ is near 00), and investigate the dimension of the spectrum of $H_{\lambda,\alpha,\omega}$ in this context.
-Since $R_{\alpha}$ is minimal, it is well-known (see Theorem 4.9 in [DF]) that the spectrum of $H_{\lambda,\alpha,\omega}$ is independent of $\omega$, so we will denote it by $\sigma_{\lambda,\alpha}$.
+Since $R_{\alpha}$ is minimal, it is well-known (see Theorem 4.9 in [DF1]) that the spectrum of $H_{\lambda,\alpha,\omega}$ is independent of $\omega$, so we will denote it by $\sigma_{\lambda,\alpha}$.
 It is known from [BIST] that $\sigma_{\lambda,\alpha}$ is a Cantor set of zero Lebesgue measure.
@@
 In the case that $\alpha=\frac{\sqrt{5}-1}{2}$, the operator $H_{\lambda,\alpha,\omega}$ is known as the Fibonacci Hamiltonian.
-In this case, many questions regarding the box-counting and Hausdorff dimensions of the spectrum have been answered (see [D1, D2, DEG, DGY, Su3] for surveys).
+It was first discovered that the spectrum of this operator is a Cantor set of zero Lebesgue measure in [Su2] and many questions regarding the box-counting and Hausdorff dimensions of the spectrum have been answered (see [D1, D2, DEG, DGY, Su3] for surveys).
 For general coupling constant, and when $\alpha=\frac{\sqrt{5}-1}{2}$, we know from a combination of results from [C, Ca, DG1] that $\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)$ is strictly between 00 and $1$.
@@
 Many authors have examined dynamical properties of this trace map (see [BGJ, BR, C, Ca, DG1, DG2, DG3, DG4, DG5, KKT]), where questions concerning hyperbolicity, Markov partitions, entropy, and more have been addressed.
-It turns out that the non-wandering set of the Fibonacci trace map restricted to any $S_{\lambda}$, $\lambda>0$, is a locally maximal hyperbolic set that is homeomorphic to a Cantor set.
+It turns out that the set of points with bounded orbit under the Fibonacci trace map restricted to any $S_{\lambda}$, $\lambda>0$, is a locally maximal hyperbolic set that is homeomorphic to a Cantor set.
 This was proved for $\lambda\geq 16$ in [C], for small $\lambda$ in [DG1], and for all $\lambda$ in [Ca].
 When $\alpha$ is eventually periodic, one needs to investigate the non-wandering set of a hyperbolic map over $S_{\lambda}$, where this map is precisely a composition of finitely many trace maps, in which case several well-developed techniques from hyperbolic dynamics apply.
-In this paper, to work with the weakened assumption that $\alpha$ is of bounded-type, we incorporate non-stationary versions of various dynamical notions such as stable manifolds, Markov rectangles, and a bounded distortion estimate from [BM].
+In this paper, to work with the weakened assumption that $\alpha$ is of bounded-type, we incorporate non-stationary versions of various dynamical notions such as stable manifolds, Markov rectangles, and a bounded distortion estimate from [BM], since many of these classical results cannot be directly applied.
 1.1 Background Material
 A main ingredient for achieving such results about the spectrum $\sigma_{\lambda,\alpha}$ is to investigate the dynamics of the so-called trace maps.
-1.1.1 Trace Maps
-Definition (Trace Maps).
-We define the trace maps to be the family of maps given by
-$T_{a}:\mathbb{R}^{3}\to\mathbb{R}^{3},\quad\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}\mapsto\begin{pmatrix}x_{1}U_{a}(x_{2})-x_{3}U_{a-1}(x_{2})\\ x_{1}U_{a-1}(x_{2})-x_{3}U_{a-2}(x_{2})\\ x_{2}\end{pmatrix},$
-where $U_{a}(t)$ are the Chebyshev polynomials of the second kind. That is,
-$U_{-1}(t)=0,\ U_{0}(t)=1,\ U_{a+1}(t)=2tU_{a}(t)-U_{a-1}(t).$
-For example, the first two trace maps are given by
-$T_{1}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{2}-x_{3}\\ x_{1}\\ x_{2}\end{pmatrix}$
-and
-$T_{2}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}4x_{1}x_{2}^{2}-2x_{2}x_{3}-x_{1}\\ 2x_{1}x_{2}-x_{3}\\ x_{2}\end{pmatrix}.$
-It is a calculation to check that for the maps
-$G:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ G\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}2x_{1}x_{3}-x_{2}\\ x_{1}\\ x_{3}\end{pmatrix}$
-and
-$H:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3},\ \text{via}\ H\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}x_{1}\\ x_{3}\\ x_{2}\end{pmatrix},$
-we have
-$T_{1}=G\circ H,\ T_{a}=G^{a-1}\circ T_{1}=G\circ T_{a-1}=G^{a}\circ H.$ (1.2)
+Consider the maps $G,H:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ via
+$G(x_{1},x_{2},x_{3}):=(2x_{1}x_{3}-x_{2},x_{1},x_{3})\ \text{and}\ H(x_{1},x_{2},x_{3}):=(x_{1},x_{3},x_{2}).$
+For $a\in\mathbb{N}$, we define the $a^{\text{{th}}}$-trace map to be $T_{a}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ via
+$T_{a}=G^{a}\circ H.$ (1.2)
 For $\lambda>0$, we consider the family of cubic surfaces given by
 $S_{\lambda}:=\left\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-2x_{1}x_{2}x_{3}=1+\frac{\lambda^{2}}{4}\right\}.$
-Each surface $S_{\lambda}$ is invariant under each trace map and from here, we will consider the restriction of each $T_{a}$ to $S_{\lambda}$.
-Given a sequence $\overline{a}=(a_{n})$ in $\mathbb{N}$, denote
-$\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right):=\left\{\boldsymbol{x}\in S_{\lambda}:\left(T_{a_{n}}\circ\cdots T_{a_{1}}(\boldsymbol{x})\right)_{n\in\mathbb{N}}\ \text{is bounded}\right\}.$
+Each surface $S_{\lambda}$ is invariant under each trace map. Given a sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N}$, denote
+$\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right):=\left\{\boldsymbol{x}\in S_{\lambda}:\left(T_{a_{n}}\circ\cdots{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ}T_{a_{1}}(\boldsymbol{x})\right)_{n\in\mathbb{N}}\ \text{is bounded}\right\}.$
 Given irrational $\alpha\in(0,1)$, we denote its continued fraction expansion by
@@
 $L_{\lambda}:=\left\{\left(\frac{E-\lambda}{2},\frac{E}{2},1\right):E\in\mathbb{R}\right\},$
-which is contained in $S_{\lambda}$, then our goal is to estimate the Hausdorff dimension of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ as $\lambda\rightarrow 0$.
-We will estimate, from below, the Hausdorff dimension of a subset of $\Lambda_{\lambda}^{\text{bnd}}$ that consists of all points that remain in a central portion of the surface $S_{\lambda}$, under our prescribed iteration of trace maps.
-We will now go through some preliminaries and then outline the structure of the paper in Section 1.2.
-1.1.2 Contraction Mappings
-The following proposition is a non-stationary version of the contraction mapping principle.
-Proposition 1.
-Let $(X,d)$ be a complete metric space and $K\subset X$ a compact set. Suppose that there is a $\mu\in(0,1)$ and, for each $n\in\mathbb{N}$, a map $f_{n}:K\rightarrow K$ such that
-$d\left(f_{n}(x),f_{n}(y)\right)\leq\mu d(x,y),$
-for all $x,y\in K$. Then, the following hold:
-  • (1)
-    There is a $x^{*}\in K$ such that
-    $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*},$
-    for any $x\in K$.
-  • (2)
-    For each $\epsilon>0$, there is a $\epsilon^{\prime}>0$ such that if for each $n\in\mathbb{N}$, there is a mapping $h_{n}:K\rightarrow K$ with
-    $d\left(h_{n}(x),h_{n}(y)\right)\leq\mu^{\prime}d(x,y)\ \text{and}\ \sup_{n}d_{C^{0}(K)}(g_{n},h_{n})<\epsilon^{\prime},$
-    for some $\mu^{\prime}\in(0,1)$ and for all $x,y\in K$, then
-    $d\left(x^{*},y^{*}\right)<\epsilon,$
-    where $y^{*}$ is the element such that
-    $\lim\limits_{n\rightarrow\infty}h_{1}\circ\cdots\circ h_{n}(x)=y^{*},$
-    for each $x\in X$.
-Proof.
-  • (1)
-    By compactness of $K$, there is an $M>0$ so that $d(x,y)\leq M$ for all $x,y\in M$. Then, by induction on $n$, we obtain that for any $x\in K$,
-    $d\left(f_{1}\circ\cdots\circ f_{n}(x),f_{1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}d\left(x,f_{n+1}\circ\cdots\circ f_{m}(x)\right)\leq\mu^{n}M,$
-    for every $n$ and $m$ such that $n\leq m$, so that $\left(f_{1}\circ\cdots\circ f_{n}(x)\right)$ is a Cauchy sequence.
-By the completeness of $X$ and compactness of $K$, there must be a $x^{*}\in K$ so that $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*}.$
-    Now, let $x^{\prime}\in K$. Then, again by induction on $n$, we have
-    $d\left(f_{1}\circ\cdots f_{n}(x),f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)\right)\leq\mu^{n}d\left(x,x^{\prime}\right)\leq\mu^{n}M,$
-    for any $n$, and hence $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}\left(x^{\prime}\right)=x^{*}.$
-  • (2)
-    Set $\epsilon^{\prime}=\frac{\epsilon(1-\mu)}{2}$. Then, an induction argument gives that
-    $d\left(f_{1}\circ\cdots\circ f_{n}(x),h_{1}\circ\cdots\circ h_{n}(x)\right)\leq\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1},$
-    for any $x\in K$. We can deduce that
-    $d\left(x^{*},y^{*}\right)\leq d\left(f_{1}\circ\cdots\circ f_{n}(x),x^{*}\right)+d\left(y^{*},h_{1}\circ\cdots\circ h_{n}(x)\right)+\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1}$
-    where the last quantity approaches $\frac{\epsilon^{\prime}}{1-\mu}=\frac{\epsilon}{2}$ as $n\rightarrow\infty$.
-1.1.3 Hyperbolic Dynamics
-Suppose that $M$ is a smooth compact Riemannian surface without boundary.
-For a $C^{r}$, $r\geq 1$, diffeomorphism $f:M\rightarrow M$, an $f$-invariant set $\Lambda\subset M$ is a hyperbolic set for $f$ provided there are values $0<c<1<\mu$ such that for each $x\in\Lambda$, there exist subspaces $H_{x}$ and $V_{x}$, that depend continuously on $x$, such that $T_{x}M=H_{x}\oplus V_{x}$ and for the sets
-$K^{s}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{1}|\leq c|v_{2}|\},$
-$K^{u}(x):=\{(v_{1},v_{2})\in H_{x}\oplus V_{x}:|v_{2}|\leq c|v_{1}|\},$
-we have
-  • ${D_{x}f\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f(x))$ and ${D_{x}f^{-1}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}(x)\right)$
-  • $\left\|D_{x}fv\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
-  • $\left\|D_{x}f^{-1}v\right\|\geq\mu\|v\|$ for $v\in K^{u}(\boldsymbol{x})$
-[…]
-It is well-known from [HP] that the stable and unstable sets of $\boldsymbol{x}\in M$ given by
-$W^{s}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{n}(\boldsymbol{x})-f^{n}(\boldsymbol{y})\right|=0\right\}$
-and
-$W^{u}(\boldsymbol{x}):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f^{-n}(\boldsymbol{x})-f^{-n}(\boldsymbol{y})\right|=0\right\}$
-are images of $C^{r}$ injective immersions from $\mathbb{R}$ to $M$.
-Hence, it is appropriate to refer to these sets as stable, respectively unstable, manifolds.
-When $\Lambda=M$, and since $\text{dim}(M)=2$, these objects can be used to construct Markov partitions of $M$ that are composed of Markov rectangles of arbitrarily small diameter (see Appendix 2 of [PT]).
-If $\overline{f}=(f_{n})$ is a sequence of $C^{1}$ diffeomorphisms of $M$, then we say $\overline{f}$ satisfies a common cone condition on $M$ if for each $x\in M$, there are cones $K^{s}(x)$ and $K^{u}(x)$ as above such that for each $n\in\mathbb{N}$, we have
-  • ${D_{x}f_{n}\left(K^{u}(x)\right)}\subseteq\text{int}K^{u}(f_{n}(x))$ and ${D_{x}f^{-1}_{n}\left(K^{s}(x)\right)}\subseteq\text{int}K^{s}\left(f^{-1}_{n}(x)\right)$
-  • $\left\|D_{x}f_{n}v\right\|\geq\mu\|v\|$ for $v\in K^{s}(x)$
-  • $\left\|D_{x}f^{-1}_{n}\right\|\geq\mu\|v\|$ for $v\in K^{u}(x)$
-[…]
-$W^{s}(\boldsymbol{x},m):=\left\{\boldsymbol{y}\in M:\lim\limits_{n\rightarrow\infty}\left|f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{x})-f_{n+m}\circ\cdots\circ f_{m}(\boldsymbol{y})\right|=0\right\}$
-is the image of an injective $C^{r}$ immersion from $\mathbb{R}$ to $M$.
-This result can be taken from the theory of random hyperbolic dynamics (see Chapter 7 of [Ar] and references therein), or one can use arguments similar to Theorem 6.2.8 from [KH] in combonation with Proposition 1, which involves graph transform techniques, to obtain this result.
-More general versions of this statement for Anosov families have been studied in [Mu].
-We use a version of this statement, and sketch the proof, for our relevant family of hyperbolic maps in Section 3.4.
-1.1.4 Thickness of Cantor Sets
-A fundamental tool we will use to estimate Hausdorff dimension of Cantor sets is the notion of thickness for Cantor sets.
-Given a Cantor set $\mathcal{C}\subset\mathbb{R}$ with convex hull $I$, we say that $U\subset I\setminus\mathcal{C}$ is a gap of $\mathcal{C}$ if it is a connected component of $I\setminus\mathcal{C}$.
-Any enumeration $\mathcal{U}=\{U_{n}\}$ of the gaps of $\mathcal{C}$ is called a presentation of $\mathcal{C}$.
-Given a boundary point $u\in\mathcal{C}$ of some gap $U_{k}$, we call the connected component $B$ contained in $I\setminus\left(U_{1}\cup\dots\cup U_{k}\right)$ that contains $u$ a bridge of $U_{k}$.
-We define
-$\tau(\mathcal{C},\mathcal{U},u):=\frac{|B|}{|U_{k}|},$
-and set the thickness of $\mathcal{C}$ to be
-$\tau(\mathcal{C}):=\sup\limits_{\mathcal{U}}\inf\limits_{u}\tau(\mathcal{C},\mathcal{U},u).$
-It is related to the Hausdorff dimension of $\mathcal{C}$ by the inequality
-$\text{dim}_{H}(\mathcal{C})\geq\frac{\log 2}{\log\left(2+\frac{1}{\tau(\mathcal{C})}\right)}.$ (1.3)
-(see Chapter 4 of [PT]).
-It is clear that these notions carry over to a Cantor set that belongs to a $C^{1}$ curve $\gamma\subset S^{1}$ where the gaps and bridges are measured using length relative to the curve $\gamma$.
-1.1.5 Continued Fraction Expansion
-The approach to our proof begins with some observations about irrational rotations on $S^{1}$ and their connections to Diophantine approximations.
-Given an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$, set $p_{n}(\alpha)\in\mathbb{Z}$ and $q_{n}(\alpha)\in\mathbb{N}$ to be the coprime integers such that
-$\frac{p_{n}(\alpha)}{q_{n}(\alpha)}=[a_{1},a_{2},\dots,a_{n}].$
-We know that these values satisfy the following recurrence equations:
-$p_{k+1}(\alpha)=a_{k+1}p_{k}(\alpha)+p_{k-1}(\alpha),\ p_{1}(\alpha)=p_{0}(\alpha)=1$ (1.4)
-$q_{k+1}(\alpha)=a_{k+1}q_{k}(\alpha)+q_{k-1}(\alpha),\ q_{1}(\alpha)=a_{1},\ q_{0}(\alpha)=1$ (1.5)
-and that
-$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|<\frac{1}{q_{k}(\alpha)}.$ (1.6)
-In regards to an irrational rotation by $\alpha=[a_{1},a_{2},\dots]$ on $S^{1}$, the Three Distance Theorem (see for example Section 2.1 of [I]) states that for each positive integer $n$, the values $\{m\alpha\}$, $1\leq m\leq n$, divide the unit circle into $n+1$ many intervals of at most three possible lengths, one of which is the sum of the other two.
-The precise lengths of the intervals come from how $n$ relates to the values $p_{k}(\alpha)$ and $q_{k}(\alpha)$.
-In particular, if we choose $n=(r+1)q_{k}(\alpha)+q_{k-1}(\alpha)-1$ for some $k\geq 1$ and $1\leq r<a_{k+1}$, then the set of points $\{m\alpha\}$, $1\leq m\leq n$, split $S^{1}$ into $n+1$ many intervals of only two possible lengths given by
-$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|\ \text{or}\ \left|q_{k-1}(\alpha)\alpha-p_{k-1}(\alpha)\right|-r\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|,$
-and by (1.6), both of these quantities are no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$. As a corollary, we obtain the following result which will be used later on.
-Proposition 2.
-Suppose that $\mathcal{A}\subset\mathbb{N}$ is finite.
-For each $\epsilon>0$, there is a $n=n(\epsilon,\mathcal{A})\in\mathbb{N}$ such that for any $\alpha=[a_{1},a_{2},\dots]$ with $a_{i}\in\mathcal{A}$, we must have
-$\min\limits_{1\leq m\leq n}\left|m\alpha-y\right|<\epsilon$
-for any $y\in S^{1}$.
-Proof.
-Let $\epsilon>0.$ Set $Y:=\{\alpha=[a_{1},a_{2},\dots]:a_{i}\in\mathcal{A}\}\subset[0,1]\setminus\mathbb{Q}$ and let $a^{\prime}$ be the smallest element of $\mathcal{A}$.
-Then using the recurrence relation (1.5) and an induction argument, we find that for each $k\geq 1$, there is an $M_{k}>0$ such that
-$q_{k}\left(\left[\overline{a^{\prime}}\right]\right)\leq\sup\limits_{\alpha\in Y}|q_{k}(\alpha)|\leq M_{k}.$
-Choose $k$ large enough so that
-$\frac{1}{q_{k-1}\left(\left[\overline{a^{\prime}}\right]\right)}+\frac{1}{q_{k}\left(\left[\overline{a^{\prime}}\right]\right)}<\epsilon.$
-Set $n:=2M_{k}+M_{k-1}-1$ and let $\alpha\in Y$ be arbitrary.
-By the above discussion, for $n_{k}:=2q_{k}(\alpha)+q_{k-1}(\alpha)-1\leq n$, we have that the points $\{m\alpha\}$, $1\leq m<n_{k}$, split $S^{1}$ into $n_{k}+1$ many intervals of lengths no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$ which is less than $\epsilon$.
-So, it follows that the set $\{m\alpha\}_{m=1}^{n}$ splits $S^{1}$ into finitely many intervals of length less than $\epsilon$.
-1.2 Outline of Paper
-Let us fix an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ such that the sequence $(a_{n})_{n\in\mathbb{N}}$ is bounded.
-Since we are working in a small range of $\lambda$, our hope is to compare the dynamics of our trace maps restricted to $S_{\lambda}$, to the dynamics of the trace maps on $S_{0}$.
-However, this comparison becomes difficult near the cusps of $S_{\lambda}$ due to how the curvature of the surface changes as $\lambda\rightarrow 0$.
-For any points that has a bounded orbit, its orbit may enter a neighborhood of one of the cusps of $S_{\lambda}$ for some amount of time but then return to a central portion of the surface where the dynamics are comparable to those on $S_{0}$.
+which is contained in $S_{\lambda}$, then our goal is to estimate the Hausdorff dimension of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ as $\lambda\rightarrow 0$ (see Figure 1).
+Figure 1: Surface $S_{\frac{1}{4}}$ and line $L_{\frac{1}{4}}$.
+1.2 Approach and Outline of Paper
+Consider the points $P_{1}=(1,1,1),\ P_{2}=(-1,-1,1),\ P_{3}=(1,-1,-1),$ and $P_{4}=(-1,1,-1)$.
+For $\rho>0$, let us set $O_{\rho}:=\bigcup_{i=1}^{4}B_{\rho}\left(P_{i}\right),$ so that for small enough $\lambda$, we will have that $S_{\lambda}\setminus O_{\rho}$ is composed of five disjoint components, only one of which is bounded (see Figure 2).
+Denote this bounded component by $\mathbb{S}_{\lambda,O_{\rho}}$.
+For a fixed sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N},$ and for small enough $\lambda$, we define
+$\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right):=\left\{\boldsymbol{x}\in\mathbb{S}_{\lambda,O_{\rho}}:T_{\lambda,a_{n}}\circ\cdots\circ T_{\lambda,a_{1}}(\boldsymbol{x})\in\mathbb{S}_{\lambda,O_{\rho}},\ \text{for all}\ n\in\mathbb{N}\right\}.$
+This set can be thought of as a non-stationary survival set. We will then prove
+Theorem 3.
+Let $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ be an irrational of bounded type and set $\overline{a}=(a_{n})$. For each $\epsilon>0$, there is a $\rho>0$ and $\lambda_{0}>0$ such that
+$\text{dim}_{H}\left(L_{\lambda}\cap\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\right)\geq 1-\epsilon$
+for all $\lambda\in[0,\lambda_{0}]$.
+Since $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\subset\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, together with Theorem 2, this will immediately imply Theorem 1.
+Figure 2: Depiction of $S_{\lambda}\setminus O_{\rho}$.
+The motivation for studying the set $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)$ comes from the fact that near each $P_{i}$, the curvature of the surface $S_{\lambda}$ becomes large as $\lambda\rightarrow 0$.
+This makes analyzing the trace map dynamics near these cusps complicated.
+For any point that has a bounded orbit, its orbit may enter a neighborhood of one of the cusps of $S_{\lambda}$ for some amount of time but then return to a central portion of the surface where the dynamics are comparable to those on $S_{0}$.
 This behavior results in the need for technical distortion estimates, and makes obtaining a common cone condition for our family of trace maps, over these cusps, rather difficult.
 These details have been written in Section 3 of [DG3] for the map $T_{1}$ only.
-To avoid this obstacle, we will first remove neighborhoods of $S_{\lambda}$ near the cusps, so that the remaining surface is a union of five disjoint components, only one of which is bounded.
-We refer to this bounded component as a central portion of $S_{\lambda}$.
-We then turn our attention to the set of points from this central portion, whose orbits never leave this region.
+So, to avoid this obstacle, we will first remove neighborhoods of $S_{\lambda}$ near the cusps, and instead study the set of points whose orbits remain (or survive) in the left over bounded component.
 The points from $L_{\lambda}$ that satisfy this behavior form a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, and it is sufficient to show that this subset has Hausdorff dimension near $1$ as $\lambda$ approaches 00.
-Provided $\lambda$ is sufficiently small, it turns that each trace map restricted to a desired central portion of $S_{\lambda}$ is a factor of an Anosov map restricted to $\mathbb{T}^{2}$ with a small open set removed, and this map is a perturbation of a hyperbolic linear automorphism.
+Provided $\lambda$ is sufficiently small, it turns out that each trace map restricted to a desired central portion of $S_{\lambda}$ is a factor of an Anosov map restricted to the two torus $\mathbb{T}^{2}:=\mathbb{R}^{2}/\mathbb{Z}^{2}$ with a small open set removed, and this map is a perturbation of a hyperbolic linear automorphism.
+Thus, as we will see later in the paper (see Section 4), studying the set $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)$ can be translated to a problem involving a sequence of toral Anosov maps.
 The structure of the paper is outlined as follows:
   • Section 2 is dedicated to analyzing basic properties of the trace maps $T_{a}$ restricted to each surface $S_{\lambda}$ for a small range of $\lambda$, and we also develop some tools and notations that will be used later in the main proof.
-  • In the beginning of Section 3, using some results from 2, we find an open set $O\subset\mathbb{T}^{2}$ composed of four small neighborhoods, a sequence of maps $\left(\tilde{T}_{\lambda,a_{n}}\right)$ defined on $\mathbb{T}^{2}\setminus O$, and a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O$ that tends to the line $\left[0,\frac{1}{2}\right]\times\{0\}$ as $\lambda\rightarrow 0$, such that the set $\mathcal{C}({\lambda},O):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\right\}$ is a Cantor set in $\hat{L}_{\lambda}$ that has dimension equal to the dimension of a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$ for a small range of $\lambda$.
-The rest of the paper is dedicated to showing that for this small range of $\lambda$, the Hausdorff dimension of $C(\lambda,O)$ is close to $1$.
-  • Our approach to estimating the dimension of $C(\lambda,O)$ is to show that for small $\lambda$, there is a Cantor set of large thickness that is contained in $C(\lambda,O)$.
-This is the goal of Section 3, and to do so, we extend each map $\tilde{T}_{\lambda,a_{n}}$ to an Anosov diffeomorphisms on $\mathbb{T}^{2}$, and analyze the non-stationary dynamics of these extended maps.
-This Anosov family will be constructed to satisfy a common cone condition, and we derive a non-stationary version of the Stable Manifold Theorem for these maps in Section 3.4.
-  • Section 3.5 is dedicated to constructing a particular set of dynamical rectangles, which will help us estimate the dimension of $\mathcal{C}(\lambda,O)$.
-Roughly speaking, we will construct a family of curvilinear rectangles that have two opposite boundaries composed of pieces of the stable manifolds developed in Section 3.4, and these rectangles will have controlled size and contain the open set $O$.
-We use the common cone conditions to control the size of these rectangles.
-  • In Section 3.6, we use the properties of the dynamical rectangles constructed in Section 3.5 to construct a Cantor set of large thickness that is contained in $C(\lambda,O)$.
-Roughly speaking, this Cantor set is derived from considering the points of $\hat{L}_{\lambda}$ that do not enter a family of associated rectangles containing $O$, under our non-stationary iterations of extended Anosov maps.
-The dynamical nature of these rectangles allows us to estimate the thickness of the newly derived Cantor set.
-We use this result to prove Theorem 1.
-As a choice of notation, we will always use dist to denote the Hausdorff distance between sets and $d_{C^{0}}$ to denote the $C^{0}$ distance between mappings.
+  • In Section 4, we state and prove Theorem 4, which is a translation of Theorem 3 to a problem concerning the non-stationary dynamics of a sequence of toral maps.
+The proof of this statement is done through the entirety of Section 4, and hence, we give an outline of the argument immediately after the statement of Theorem 4.
+  • In Section 5, we use Theorem 4 to conclude the proofs of Theorem 3 and Theorem 1.
+As a choice of notation, we will always use dist to denote the minimal distance between sets and $\text{dist}_{H}$ to denote the Hausdorff distance between sets and $d_{C^{0}}$ to denote the $C^{0}$ distance between mappings.
 With a slight abuse of notation, we will denote $(x,y)$ to be the standard coordinates in both $\mathbb{T}^{2}$ and $\mathbb{R}^{2}$.
@@
 2.1 Torus Model for $\lambda=0$
-The surface $S_{0}$ is smooth everywhere except at $P_{1}=(1,1,1),\ P_{2}=(-1,-1,1),\ P_{3}=(1,-1,-1),$ and $P_{4}=(-1,1,-1)$.
-Let us denote $\mathbb{S}:=S_{0}\cap\{|x_{1}|\leq 1,|x_{2}|\leq 1,|x_{3}|\leq 1\}$.
-We know
+Notice that the surface $S_{0}$ is smooth everywhere except at each $P_{i}$, $i=1,\dots,4$. Let us denote $\mathbb{S}:=S_{0}\cap\{|x_{1}|\leq 1,|x_{2}|\leq 1,|x_{3}|\leq 1\}$. We know
 Lemma 1.
@@
 Proof.
-This is a direct calculation. ∎
+This is a direct calculation which can be found in Lemma 3.15 in [M].
 For each $a\in\mathbb{N}$, $A_{a}$ has two perpendicular eigenspaces with corresponding real eigenvalues for which one is less than $1$ in absolute value, and the other is greater than $1$.
@@
 The corresponding stable and unstable eigenvalues are given by
-$\mu^{s}(a):=\frac{a-\sqrt{1+a^{2}}}{2}\ \text{and}\ \mu^{u}(a):=\frac{a+\sqrt{1+a^{2}}}{2}.$
-For each $a>1$, the unstable eigenspace $E^{u}_{a}$ lies between the lines $E_{1}^{u}$ and the line $\{y=0\}$.
-Let $\beta\in(0.0.1]$, and consider the vector spaces $V_{1}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$.
+$\mu^{s}(a):=\frac{a-\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4+a^{2}}}}{2}\ \text{and}\ \mu^{u}(a):=\frac{a+\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4+a^{2}}}}{2}.$
+For each $a>1$, the unstable eigenspace $E^{u}_{a}$ lies between the lines $E_{1}^{u}$ and the line $\{y=0\}$ in the projective sense.
+Let $\beta\in(0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}0.1]$, and consider the vector spaces $V_{1}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$.
 We let $K^{u}_{\beta}$ be the collection of vectors that are between $V_{1}(\beta)$ and $V_{2}(\beta)$, which indeed contains each $E^{u}_{a}$.
@@
 $K^{u}_{\beta}:=\left\{c_{1}\begin{pmatrix}1\\ -\beta\end{pmatrix}+c_{2}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}:c_{1}c_{2}\geq 0\right\}$
-and we also set $V_{0}(\beta)$ to be the line mid-way between $V_{1}(\beta)$ and $V_{2}(\beta)$. Then, we have the following:
+and we also set $V_{0}(\beta)$ to be the line mid-way between $V_{1}(\beta)$ and $V_{2}(\beta)$. That is,
+$V_{0}(\beta):=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}+\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}.$
+Then, we have the following:
 Lemma 2.
 For each $\beta\in(0,0.1]$, for the cones given by $K^{u}_{\beta}$ and $K^{s}_{\beta}:=\left(K_{\beta}^{u}\right)^{\perp}$, there exists a $\overline{\mu}_{\beta}>1$ such that for each $a\in\mathbb{N}$, we have
-  • $A_{a}^{-1}\left(K^{s}_{\beta}\right)\subset\text{int}K^{s}_{\beta}$ and $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}K^{u}_{\beta}$
+  • $A_{a}^{-1}\left(K^{s}_{\beta}\right)\subset\text{int}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(K^{s}_{\beta}\right)}$ and $A_{a}\left(K^{u}_{\beta}\right)\subset\text{int}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(K^{u}_{\beta}\right)}$
   • $\overline{\mu}_{\beta}|v|\leq|A_{a}v|\leq\mu^{u}(a)|v|$ for all $v\in K_{\beta}^{u}$
-  • $\overline{\mu}_{\beta}|v|\leq\left|A_{a}^{-1}v\right|\leq\left(\mu^{s}(a)\right)^{-1}|v|$ for all $v\in K_{\beta}^{s}$
+  • $\overline{\mu}_{\beta}|v|\leq\left|A_{a}^{-1}v\right|\leq\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|\mu^{s}(a)\right|}\right)^{-1}|v|$ for all $v\in K_{\beta}^{s}$
 Proof.
@@
 $(1+2\beta)\left(a-\beta\right)>1\Rightarrow 1+2\beta>\frac{1}{a-\beta}$
-and hence $\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}$ belongs to a line strictly between $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$.
+and hence $\left(a-\beta\right)\begin{pmatrix}1\\ \frac{1}{a-\beta}\end{pmatrix}$ belongs to a line strictly between $V_{1}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ -\beta\end{pmatrix}\right\}$ and $V_{2}(\beta)=\text{span}\left\{\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}\right\}$ in the projective sense.
 A similar calculation shows that the same is true for $A_{a}\begin{pmatrix}1\\ 1+2\beta\end{pmatrix}$.
@@
 In other words, for any $a$ and $v\in K^{u}_{\beta}$,
-$\left|A_{a}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{v}{|v|}\right|$
-$\geq\left|A_{1}\frac{\left(1,-\beta\right)^{T}}{\left|\left(1,-\beta\right)^{T}\right|}\right|=\frac{\sqrt{\left(1-\beta\right)^{2}+1}}{\sqrt{\beta^{2}+1}}=:\overline{\mu}_{\beta}.$
+$\left|A_{a}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{v}{|v|}\right|\geq\left|A_{1}\frac{\left(1,-\beta\right)^{T}}{\left|\left(1,-\beta\right)^{T}\right|}\right|=\frac{\sqrt{\left(1-\beta\right)^{2}+1}}{\sqrt{\beta^{2}+1}}=:\overline{\mu}_{\beta}.$
 The vectors from $K^{u}_{\beta}$ that are expanded the most under $A_{a}$ are the ones from $E^{u}_{a}$, and these vectors are expanded by a factor of precisely $\mu^{u}(a)$. It follows that
@@
 The invariance of $K^{s}_{\beta}$ under $A_{a}^{-1}$ and expansion estimates are checked similarly. ∎
+Remark 1.
+The range of $\beta$ in this lemma can most certainly be increased, but for our arguments, we just need any interval containing 00.
 Let $\theta_{0}(\beta)$ denote the angle between boundary lines of $K^{u}_{\beta}$.
@@
 Proof.
-For each $n\in\mathbb{N}$, we consider $\hat{A}_{a_{n}}$ to be the action on $\mathbb{R}\mathbb{P}^{1}$ induced by $A_{a_{n}}$.
+For each $n\in\mathbb{N}$, we consider $\hat{A}_{a_{n}}$ to be the action on real projective space $\mathbb{R}\mathbb{P}^{1}$ induced by $A_{a_{n}}$.
 Denote $\hat{K}^{s}\subset\mathbb{R}\mathbb{P}^{1}$ to be the set obtained from projecting $K^{s}$ onto $\mathbb{R}\mathbb{P}^{1}$, which is compact.
 Then, from Lemma 2 we have that each map $\hat{A}_{a_{n}}^{-1}$ restricted to $\hat{K}^{s}$ maps $\hat{K}^{s}$ into $\hat{K}^{s}$ and is a contraction with contraction rate no greater than the contraction rate of $\hat{A}_{1}$, which is $\mu^{s}(1)^{2}$.
-By Proposition 1, there is a $L^{*}\in\hat{K}^{s}$ such that
+By Proposition A.1, there is a $L^{*}\in\hat{K}^{s}$ such that
 $\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}(L)=L^{*},$
@@
 In particular, we know that the projection of $\text{span}\left\{(0,1)^{T}\right\}$ onto $\mathbb{R}\mathbb{P}^{1}$, denoted by $\hat{(0,1)}$, lives in $\hat{K}^{s}$, and hence $\lim\limits_{n\rightarrow\infty}\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(0,1)}\right)=L^{*}.$ We however see that the line $\text{span}\left\{A_{a_{1}}^{-1}\circ\cdots\circ A_{a_{n}}^{-1}(1,0)^{T}\right\}$ projects onto $\hat{A}_{a_{1}}^{-1}\circ\cdots\circ\hat{A}_{a_{n}}^{-1}\left(\hat{(1,0)}\right)$.
-That $E^{s}_{\alpha}$ has slope $-\frac{1}{\alpha}$ follows from
 Claim 1.
@@
 and hence has slope $-\frac{1}{[a_{1},\dots,a_{n}]}$.
-Projecting these lines onto $\mathbb{T}^{2}$, we have that each line $\pi\left(E^{s}_{\alpha}\right)$ is dense in $\mathbb{T}^{2}$.
+This shows that $E^{s}_{\alpha}$ has slope $-\frac{1}{\alpha}$, proving the lemma.
+It is clear that each line $\pi\left(E^{s}_{\alpha}\right)$ is dense in $\mathbb{T}^{2}$.
 If $Q$ is a rational point in $\mathbb{T}^{2}$, and $L\subset\pi\left(E^{s}_{\alpha}\right)$ is a line segment with midpoint 00, the length of $L$ and continued fraction expansion of $\alpha$ determine the distance from $L$ and $Q$.
-From Proposition 2, we have the following:
+This exact distance can be estimated using Proposition B.1.
+In particular, we have:
 Lemma 4.
@@
 Without loss of generality, let us assume $Q=0$.
-Let $C^{*}>0$. By Proposition 2, there is a $n=n\left(C^{*},\mathcal{A}\right)$ so that for any $\alpha=[a_{1},a_{2},\dots]$ where $a_{i}\in\mathcal{A}$, we must have
+Let $C^{*}>0$. By Proposition B.1, there is a $n=n\left(C^{*},\mathcal{A}\right)$ so that for any $\alpha=[a_{1},a_{2},\dots]$ where $a_{i}\in\mathcal{A}$, we must have
 $\min\limits_{1\leq k\leq n}|k\alpha-y|<C^{*}$
@@
     $F_{\lambda}\circ\tilde{T}_{\lambda,a}=T_{a}\restriction_{\mathbb{S}_{\lambda,O_{\rho}}}\circ F_{\lambda},$
-    where $F_{\lambda}:=\pi^{-1}_{\lambda}\circ F$.
+    where $F_{\lambda}:=\pi^{-1}_{\lambda}\circ F$ and $F$ is defined in Lemma 1.
 Moreover, we have $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$, as $\lambda\rightarrow 0$, and $r(\rho)\rightarrow 0$ as $\rho\rightarrow 0$.
 Proof.
-One can choose $\rho^{\prime}\in(0,\rho)$ such that for $O_{\rho}=\bigcup_{i=1}^{4}B_{\rho}$ and $O_{\rho^{\prime}}=\bigcup_{i=1}^{4}B_{\rho^{\prime}},$ we must have
+One can choose $\rho^{\prime}\in(0,\rho)$ such that for $O_{\rho}=\bigcup_{i=1}^{4}B_{\rho}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(P_{i}\right)}$ and $O_{\rho^{\prime}}=\bigcup_{i=1}^{4}B_{\rho^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(P_{i}\right)},$ we must have
 $T_{a}^{-1}\left(O_{\rho^{\prime}}\right)\subset O_{\rho}.$
@@
 Also, there is an open set $O^{\prime}\subset\tilde{O}_{r(\rho)}$ such that $F^{-1}\circ\pi_{\lambda}\left(\mathbb{S}_{\lambda,O_{\rho^{\prime}}}\right)=\mathbb{T}^{2}\setminus O^{\prime}$.
-Let us denote $F_{\lambda}:=\pi_{\lambda}^{-1}\circ F$.
-Let $P\in\mathbb{T}^{2}$. Then, $F_{\lambda}(P)$ is contained in $\mathbb{S}_{\lambda,O_{\rho}}$ and $\left(T_{a}\circ F_{\lambda}\right)(P)\in\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, so that
+Let us denote $F_{\lambda}:=\pi_{\lambda}^{-1}\circ F$ and let $P\in\mathbb{T}^{2}$.
+Then, $F_{\lambda}(P)$ is contained in $\mathbb{S}_{\lambda,O_{\rho}}$ and $\left(T_{a}\circ F_{\lambda}\right)(P)\in\mathbb{S}_{\lambda,O_{\rho^{\prime}}}$, so that
 $\left(F^{-1}_{\lambda}\circ T_{a}\circ F_{\lambda}\right)(P)$
@@
 $F_{\lambda}\circ\tilde{T}_{\lambda,a}=T_{a}\restriction_{\mathbb{S}_{\lambda,O_{\rho}}}\circ F_{\lambda}.$
-Since $\left\|\pi_{\lambda}\circ T_{a}\circ\pi^{-1}_{\lambda}\restriction_{\mathbb{S}_{0,O_{\rho}}}-T_{a}\restriction_{\mathbb{S}_{0,O_{\rho}}}\right\|_{C^{2}}\rightarrow 0$ as $\lambda\rightarrow 0$, it follows that $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus O_{r(\rho)}$, as $\lambda\rightarrow 0$.
+Since $\left\|\pi_{\lambda}\circ T_{a}\circ\pi^{-1}_{\lambda}\restriction_{\mathbb{S}_{0,O_{\rho}}}-T_{a}\restriction_{\mathbb{S}_{0,O_{\rho}}}\right\|_{C^{2}}\rightarrow 0$ as $\lambda\rightarrow 0$, it follows that $\left\|\tilde{T}_{\lambda,a}-\hat{T}_{0,a}\right\|_{C^{2}}\rightarrow 0$, on $\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$, as $\lambda\rightarrow 0$.
 Lastly, since $\text{diam}\left(B_{\rho}(P_{i})\right)\rightarrow 0$ as $\rho\rightarrow 0$ for each $i=1,\dots,4$, we can choose $r(\rho)$ so that $r(\rho)\rightarrow 0$ as $\rho\rightarrow 0$.
-2.3 Main Problem Setup
+3 Dynamical Tools and Setup
 We will now setup various tools and notations that will be helpful in the main proof.
-2.3.1 Boxes and Local Cones
+Namely, in Section 4 we will invoke Lemma 5 and translate Theorem 3 into a problem concerning non-stationary dynamics of Anosov maps around small neighborhoods of each $Q_{i}$.
+Rather than working with arbitrary neighborhoods, we will define a notion of dynamical rectangles, that are analogous to Markov rectangles for hyperbolic maps, that will be useful in examining the dynamics near each $Q_{i}$.
+This section is dedicated to precisely defining these rectangles, and studying some basic geometric properties that they exhibit.
+3.1 Boxes and Local Cones
+Before defining the rectangles with dynamical properties, we will first restrict ourselves to small preliminary boxes around each $Q_{i}$.
+This subsection is dedicated to choosing these boxes, deriving some needed properties for later, and also introducing some notations.
 Recalling our definition of $V$ from (2.2), consider the projections $\pi_{V}:V\oplus V^{\perp}\rightarrow V$ and $\pi_{V^{\perp}}:V\oplus V^{\perp}\rightarrow V^{\perp}$, and set .$\|.\|_{V}:=\left|\pi_{V}(.)\right|$ and .$\|.\|_{V^{\perp}}:=\left|\pi_{V^{\perp}}(.)\right|$.
@@
 $\text{Box}_{r}(\boldsymbol{Q}):=\left\{\boldsymbol{P}:\|\boldsymbol{P}-\boldsymbol{Q}\|_{V}\leq r\right\}\times\left\{\boldsymbol{P}:\|\boldsymbol{P}-\boldsymbol{Q}\|_{V^{\perp}}\leq r\right\}.$
-Remark 2.
+For these choice of sets, it is important that each map $A_{a}$ overflows the square $\text{Box}_{1}(0)$. We use this fact later in Theorem 5 and Lemma 16. Formally, we have the following:
+Lemma 6.
+Let $L$ be a line segment through 00 that is contained in $K^{u}$ and connects the boundary lines of $\text{Box}_{1}(0)$ that are parallel to $V^{\perp}$.
+Then, for any $a\in\mathbb{N}$, we have that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$ (see Figure 3).
+Moreover, for each $\beta>\beta_{0}$ sufficiently close to $\beta_{0}$, there is a $\epsilon=\epsilon(a,\beta)>0$ such that if $f\in\text{Diff}^{1}\left(\mathbb{R}^{2}\right)$ and $\left\|A_{a}-f\right\|<\epsilon,$ then $f(L)$ is a curve that overflows $\text{Box}_{1}(0)$.
+That is, $f(L)$ is contained in $K^{u}_{\beta}$, and its endpoints belong to the compliment of $\text{Box}_{1}(0)$.
+Proof.
 Suppose that $a\in\mathbb{N}$ and consider the square $\text{Box}_{1}(0)$.
-Then, from our choice of $\beta_{0}$ (see 2.2), we know that the length of any segment contained in $K^{u}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length no greater than $\frac{2}{\sqrt{3}}$ and length no less than $1$.
-If we iterate this segment under $A_{a}$, the its length will be greater than $\overline{\mu}_{\beta_{0}}$.
+Then, from our choice of $\beta_{0}$ (see 2.2), we know that the length of any segment $L$ contained in $K^{u}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length strictly less than $\frac{2}{\sqrt{3}}$ and length no less than $1$.
+If we iterate this segment under $A_{a}$, then its length will be greater than $\overline{\mu}_{\beta_{0}}$.
 However, from (2.1), we have
 $\overline{\mu}_{\beta_{0}}\geq\sqrt{2}-0.01>\frac{2}{\sqrt{3}}.$
-This calculation gives that for any line segment $L$ through 00 that is contained in $K^{u}$ and connects the boundary lines of $\text{Box}_{1}(0)$ that are parallel to $V^{\perp}$, we must have that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$ (see Figure 1).
-If $\beta>\beta_{0}$ is close enough to $\beta_{0}$, then same property will hold for $K^{u}_{\beta}$ restricted to $\text{Box}_{1}(0)$.
-Notice that in this case, $V(\beta_{0})$ is not the center line of the cone $K^{u}_{\beta}$, but the angle between its center line and $V(\beta_{0})$ is small.
-The property also holds for a sufficiently small perturbation of each map $A_{a}$.
+It follows that $A_{a}(L)$ is a line through 00 that is contained in $K^{u}$ and extends outside of the square $\text{Box}_{1}(0)$.
+From the discussion before (2.1), for $\beta>\beta_{0}$ close enough to $\beta$, we can still guarantee that the length of any segment $L$ contained in $K^{u}_{\beta}$ that connects 00 and a boundary line of $\text{Box}_{1}(0)$ parallel to $V^{\perp}$, has length strictly less than $\frac{2}{\sqrt{3}}$.
+Then, by shrinking $\beta$ is needed, we have that
+$\overline{\mu}_{\beta}\geq(\sqrt{2}-0.01)-0.001$
+and hence, this expansion is still greater than $\frac{2}{\sqrt{3}}$.
+If $L$ is such a segment and $f$ is sufficiently $C^{1}$ close to $A_{a}$, then the endpoints of $f(L)$ will be close to those of $A_{a}(L)$, and we can guarantee that $Df\left(K^{u}_{\beta}\right)\subset K^{u}_{\beta}$.
+This would imply that the tangent line at each point of $f(L)$ is contained in $K^{u}_{\beta}$, and hence $f(L)\subset K^{u}_{\beta}$.
+∎
+We now define some sets and notations that will be used frequently throughout the paper.
 […]
-Figure 1: Illustration of Remark 2
+Figure 3: Illustration of Lemma 6
 […]
@@
 where $\pi\left(\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ is the rectangular neighborhood centered at $\pi(\boldsymbol{Q})$, with sides parallel to the lines $\pi(V)$ and $\pi(V^{\perp})$, respectively, of lengths given by $\frac{1}{5}$.
-For each $r\in\left(0,\frac{1}{10}\right)$, let us denote $\mathbb{B}_{r}(Q):=\pi\left(\text{Box}_{r}\left(\pi^{-1}(Q)\right)\right)$, which is the rectangular neighborhood centered at $Q$ of side lengths $2r$.
-We set $\partial^{s}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V^{\perp}\right)$ and $\partial^{u}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V\right)$ (see Figure 2).
-In particular, let us set $\mathbb{B}_{i}:=\mathbb{B}_{\frac{1}{10}}(Q_{i})$, $\partial^{s}\mathbb{B}_{i}:=\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q_{i})$ and $\partial^{u}\mathbb{B}_{i}:=\partial^{u}\mathbb{B}_{\frac{1}{10}}(Q_{i})$.
+Definition.
+For each $r\in\left(0,\frac{1}{10}\right)$, we define (see Figure 4)
+  • $\mathbb{B}_{r}(Q):=\pi\left(\text{Box}_{r}\left(\pi^{-1}(Q)\right)\right)$, which is the rectangular neighborhood centered at $Q$ of side lengths $2r$
+  • $\partial^{s}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V^{\perp}\right)$
+  • $\partial^{u}\mathbb{B}_{r}(Q)$ to be the two line segments of $\partial\mathbb{B}_{r}(Q)$ that are parallel to the line $\pi\left(V\right)$
+In particular, let we set $\mathbb{B}_{i}:=\mathbb{B}_{\frac{1}{10}}(Q_{i})$, $\partial^{s}\mathbb{B}_{i}:=\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q_{i})$ and $\partial^{u}\mathbb{B}_{i}:=\partial^{u}\mathbb{B}_{\frac{1}{10}}(Q_{i})$.
 […]
 Figure 4: Illustration of local boxes.
-For any cone $K$ containing $K^{u}$ and transversal to $K^{s}$, we define
+Definition.
+(see Figure 5) For any cone $K$ containing $K^{u}$ and transversal to $K^{s}$, we define
 $K_{-}:=K^{u}\cap\{x\leq 0\}\ \text{and}\ K_{+}:=K\cap\{x\geq 0\}.$ (3.2)
-For a cone $K\subset\mathbb{R}^{2}$, define the local projected cone $\tilde{K}(Q):=\pi\left(\left(\boldsymbol{Q}+K\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$.
-If $K$ is a cone containing $K^{u}$ and transversal to $K^{s}$, then set $\tilde{K}_{+}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{+}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ and $\tilde{K}_{-}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{-}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$ (see Figure 3).
+For a cone $K\subset\mathbb{R}^{2}$, define the local projected cone $\tilde{K}(Q)\subset\mathbb{B}_{\frac{1}{10}}(Q)$ via
+$\tilde{K}(Q):=\pi\left(\left(\boldsymbol{Q}+K\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$
+If $K$ is a cone containing $K^{u}$ and transversal to $K^{s}$, then set
+$\tilde{K}_{+}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{+}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)\ \text{and}\ \tilde{K}_{-}(Q):=\pi\left(\left(\boldsymbol{Q}+K_{-}\right)\cap\text{Box}_{\frac{1}{10}}(\boldsymbol{Q})\right)$
+With this in mind, we define
+  • +$\partial^{s}_{+}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{+}^{u}(Q)$
+  • $\partial^{s}_{-}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{-}^{u}(Q)$
+  • $K^{s,u}(x,y)$ to be the cone $K^{s,u}$ contained in the tangent space $T_{(x,y)}\mathbb{T}^{2}$
+Lastly, we set
+$\mathcal{K}^{s}:=\left\{K^{s}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}\ \text{and}\ \mathcal{K}^{u}:=\left\{K^{u}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}},$ (3.3)
+and refer to these collections as stable, respectively unstable, cone fields.
 […]
 Figure 5: Illustration of a local cone.
-With this in mind, we set +$\partial^{s}_{+}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{+}^{u}(Q)$ and $\partial^{s}_{-}\mathbb{B}_{r}(Q)$ to be the line segment from $\partial^{s}\mathbb{B}_{r}(Q)$ that intersects $\tilde{K}_{-}^{u}(Q)$.
-We will let $K^{s,u}(x,y)$ be the cone $K^{s,u}$ contained in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, and set
-$\mathcal{K}^{s}:=\left\{K^{s}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}\ \text{and}\ \mathcal{K}^{u}:=\left\{K^{u}(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}},$ (2.5)
-and refer to these collections as stable, respectively unstable, cone fields.
-Remark 3.
+The next lemma is another property concerning these boxes that will be frequently used later in Section 4.
+Lemma 7.
 There is a $C_{1}>0$ such that if $\gamma\subset\mathbb{T}^{2}$ is a curve tangent to $\mathcal{K}^{u}$, and is contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, for some $Q\in\mathbb{T}^{2}$, with endpoints $P_{1}$ and $P_{2}$, then
 $\text{length}(\gamma)\leq C_{1}\left|P_{1}-P_{2}\right|.$
-2.3.2 Dynamical Rectangles
+Proof.
+This follows from the fact that any curve tangent to $\mathcal{K}^{u}$ contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ is the graph of a Lipschitz map from $\pi(V)\cap\mathbb{B}_{\frac{1}{10}}(Q)$ to $\pi\left(V^{\perp}\right)\cap\mathbb{B}_{\frac{1}{10}}(Q)$ with Lipschitz constant determined by the cone $K^{u}$.
+∎
+3.2 Dynamical Rectangles
+We are now ready to define these so-called dynamical rectangles.
 Definition.
 A dynamical rectangle (or rectangle) is a curvilinear rectangle $R$, such that two opposite boundary curves of $R$ are $C^{1}$ curves tangent to $\mathcal{K}^{s}$, referred to as the stable boundaries, denoted by $\partial^{s}R$, and the other two boundary curves are transversal to $\mathcal{K}^{s}$, referred to as the non-stable boundaries, and denoted by $\partial^{ns}R$.
-We say that a $C^{1}$ curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ that is tangent to $\mathcal{K}^{s}$ connects opposite boundaries of $\tilde{K}^{u}(Q)$ if its endpoints belong to $\partial\tilde{K}_{+}^{u}(Q)$.
-For any $Q\in\mathbb{T}^{2}$ and points $P,P^{\prime}\in\mathbb{B}_{\frac{1}{10}}(Q)$, we say that a curve $\gamma$ connects $P$ and $P^{\prime}$ if these are the endpoints of $\gamma$, and $\gamma$ is strictly contained in $\overline{B_{\frac{1}{10}}(Q)}$.
-If $\gamma_{1}$ and $\gamma_{2}$ are disjoint curves contained in $B_{\frac{1}{10}}(Q)$, then $\gamma$ connects these curves if it connects points from these curves.
-For each $Q\in\mathbb{T}^{2}$ and cone $K\supset K^{u}$ transversal to $K^{s}$, we say that two $C^{1}$ curves $\gamma_{1}$ and $\gamma_{2}$ contained in $\mathbb{B}_{\frac{1}{10}}(Q)$, tangent to $\mathcal{K}^{s}$, are centered at $Q$ with respect to $\tilde{K}(Q)$, if there is a sub-curve of $\gamma_{1}$ that connects opposite boundaries of $\tilde{K}_{+}(Q)$ and does not intersect $\tilde{K}_{-}(Q)$, and if there is a sub-curve of $\gamma_{2}$ that connects opposite boundaries of $\tilde{K}_{-}(Q)$ and does not intersect $\tilde{K}_{+}(Q)$, or vice versa (see Figure 4).
+Definition.
+Let $\gamma$ be a $C^{1}$ curve contained in $\mathbb{B}_{\frac{1}{10}}(Q)$ that is tangent to $\mathcal{K}^{s}$. We say that $\gamma$
+  • connects opposite boundaries of $\tilde{K}^{u}(Q)$ if its endpoints belong to $\partial\tilde{K}_{+}^{u}(Q)$
+  • connects points $P,P^{\prime}\in\mathbb{B}_{\frac{1}{10}}(Q)$, if these are the endpoints of $\gamma$, and $\gamma$ is strictly contained in $\overline{B_{\frac{1}{10}}(Q)}$
+  • connects disjoint curves $\gamma_{1},\gamma_{2}\subset B_{\frac{1}{10}}(Q)$ if it connects points of these curves
+Definition.
+Let $Q\in\mathbb{T}^{2}$ and $K\supset K^{u}$ be a cone transversal to $K^{s}$. We say
+  • Two $C^{1}$ curves $\gamma_{1},\gamma_{2}\subset\mathbb{B}_{\frac{1}{10}}(Q)$ tangent to $\mathcal{K}^{s}$, are centered at $Q$ with respect to $\tilde{K}(Q)$, if
+    • There is a sub-curve of $\gamma_{1}$ that connects opposite boundaries of $\tilde{K}_{+}(Q)$ and does not intersect $\tilde{K}_{-}(Q)$
+    • There is a sub-curve of $\gamma_{2}$ that connects opposite boundaries of $\tilde{K}_{-}(Q)$ and does not intersect $\tilde{K}_{+}(Q)$
+    or vice versa (see Figure 6).
+  • A rectangle $R\subset B_{\frac{1}{10}}(Q)$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and its stable boundary curves are centered at $Q$ with respect to $\tilde{K}(Q)$.
+In this case we denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$.
+(see Figure 6)
+Remark 3.
 Notice that any pair of curves centered at $Q$ with respect to $\tilde{K}(Q)$, must be centered at $Q$ with respect to $\tilde{K}^{u}(Q)$.
 If not specified otherwise, we say that two curves are centered at $Q$, if they are centered at $Q$ with respect to $\tilde{K}^{u}(Q)$.
-A rectangle $R\subset B_{\frac{1}{10}}(Q)$ that contains $Q$ in its interior is centered at $Q$ with respect to $\tilde{K}(Q)$ if its stable boundary curves are centered at $Q$ with respect to $\tilde{K}(Q)$ (see Figure 4).
-We denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$.
+Also, we denote +$\partial^{s}_{+}R$ to be the curve from $\partial^{s}R$ that has a segment connecting opposite boundaries of $\tilde{K}^{u}_{+}(Q)$ and $\partial^{s}_{-}R$ to be the other curve which has a segment that connects opposite boundaries of $\tilde{K}^{u}_{-}(Q)$.
 It should also be noted that ±$\partial^{s}_{\pm}R\cap\tilde{K}^{u}_{\mp}(Q_{i})=\emptyset$.
+3.3 Geometric Properties of Dynamical Rectangles
+We now derive two propositions that describe useful geometric properties concerning these local cones and dynamical rectangles. The next proposition is illustrated by Figure 7.
+Proposition 1.
+There is a $C^{*}_{0}>0$ such that for any $i=1,\dots,4$, the following hold:
+  • (1)
+    If $\gamma$ is a $C^{1}$ curve tangent to $\mathcal{K}^{s}$ that is contained in $\overline{\mathbb{B}}_{i}$ satisfying
+    $\text{dist}(\gamma,Q_{i})<C^{*}_{0},$
+    then $\gamma$ is disjoint from $\partial^{s}\mathbb{B}_{i}$.
+  • (2)
+    If $\kappa$ is a curve tangent to $\mathcal{K}^{u}$ with endpoints $P_{1}$ and $P_{2}$ satisfying $P_{1}\in\tilde{K}_{+}(Q_{i})$, $P_{2}\in\tilde{K}_{-}(Q_{i})$, and
+    $\max\left\{\left|P_{1}-Q_{i}\right|,\left|P_{2}-Q_{i}\right|\right\}<C^{*}_{0},$
+    we must have $\kappa\subset\mathbb{B}_{i}$. That is $\kappa$ is disjoint from $\partial^{u}\mathbb{B}_{i}$.
+Proof.
+We prove part (1).
+Since $\text{dist}\left(\partial^{s}\mathbb{B}_{i},\tilde{K}^{s}(Q_{i})\right)>0$, by construction of $K^{s}$, if $L\subset\overline{\mathbb{B}}_{i}$ is a line segment parallel to a segment of $\partial\tilde{K}^{s}(Q_{i})$, and sufficiently close to $Q_{i}$, we must also have $\text{dist}\left(\partial^{s}\mathbb{B}_{i},L\right)>0.$ It follows that the same is true for any curve $\gamma\subset\overline{\mathbb{B}}_{i}$ that is tangent to $\mathcal{K}^{s}$.
+The proof of (2) is similar.
 […]
 $\partial^{s}_{-}R$$\partial^{ns}R$$R$+$\partial^{s}_{+}R$$Q$$\tilde{K}(Q)$$\boldsymbol{-}$$\boldsymbol{+}$
-Figure 4: The right figure depicts two curves $\gamma_{1}$ and $\gamma_{2}$ that are centered at $Q$ with respect to a cone $\tilde{K}(Q)$, and the left figure depicts a dynamical rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$.
-Remark 4.
-Notice there is a $C^{*}_{0}>0$ such that for each $i=1,\dots,4$, if $\gamma$ is a $C^{1}$ curve tangent to $\mathcal{K}^{s}$ that is contained in $\overline{\mathbb{B}}_{i}$ satisfying $\text{dist}(\gamma,Q_{i})<C^{*}_{0}$, then it is disjoint from $\partial^{s}\mathbb{B}_{i}$ (see Figure 5).
-Moreover, if $\kappa$ is a curve tangent to $\mathcal{K}^{u}$ with endpoints $P_{1}$ and $P_{2}$ satisfying $P_{1}\in\tilde{K}_{+}(Q_{i})$, $P_{2}\in\tilde{K}_{-}(Q_{i})$, and $\max\left\{\left|P_{1}-Q_{i}\right|,\left|P_{2}-Q_{i}\right|\right\}<C^{*}_{0}$, we must have $\kappa\subset\mathbb{B}_{i}$.
-That is $\kappa$ is disjoint from $\partial^{u}\mathbb{B}_{i}$.
-$Q$$\tilde{K}^{s}(Q)$$\partial^{s}\mathbb{B}_{\frac{1}{10}}(Q)$$\gamma$
-$\tilde{K}(Q)$$\tilde{K}^{u}(Q)$$Q$$\kappa$
-Figure 5: Illustration of Remark 4.
+Figure 6: The left figure depicts two curves $\gamma_{1}$ and $\gamma_{2}$ that are centered at $Q$ with respect to a cone $\tilde{K}(Q)$, and the right figure depicts a dynamical rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$.
+[…]
+[…]
+Figure 7: Illustration of Proposition 1.
 On the top figure, any curve $\gamma$ tangent to $\mathcal{K}^{s}$ that intersects the center “diamond” must stay between the dotted lines.
@@
 So, one can choose a neighborhood $O$ of $Q$ accordingly.
-Remark 5.
-If $K\supsetneq K^{u}$ is a cone disjoint from $K^{s}$, then for any rectangle $R$ centered at $Q$ with respect to $\tilde{K}(Q)$, there is a neighborhood $O$ of $Q$ such that for any curve $\gamma$ tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects $O$ must not intersect $\partial^{ns}R$ (see Figure 6).
-The size of the neighborhood $O$ depends on the cone $K$ and $\text{dist}(\partial R,Q)$.
-In particular, if $\{R_{n}\}$ is a family of rectangles that are centered at $Q$ with respect to $\tilde{K}(Q)$ and there is a $l^{*}>0$ such that $\inf_{n}\text{dist}\left(\partial R_{n},Q\right)\geq l^{*}$, then there is a $r=r(K,l^{*})>0$ such that for each $n$ and any curve $\gamma$ tangent to $\mathcal{K}^{u}$, we must have that any component of $\gamma\cap R_{n}$ that intersects $B_{r}(Q)$, must not intersect $\partial^{ns}R_{n}$.
-That is, for any curve $\gamma$ tangent to $\mathcal{K}^{u}$ connecting $\partial_{+}^{s}R_{n}$ and $\partial_{-}^{s}R_{n}$ with $\gamma\cap B_{r}(Q)\neq\emptyset$, we must have $\gamma\subset R_{n}$ and $\gamma\cap\partial^{ns}R=\emptyset$.
+The next proposition is illustrated by Figure 8.
+Proposition 2.
+Suppose $K\supsetneq K^{u}$ is a cone disjoint from $K^{s}$, and $R$ is a rectangle centered at $Q$ with respect to $\tilde{K}(Q)$.
+Then, there is a neighborhood $O$ of $Q$ such that for any curve $\gamma$ tangent to $\mathcal{K}^{u}$, we must have that any component of $\gamma\cap R$ that intersects $O$ must not intersect $\partial^{ns}R$.
+Proof.
+The proof is analogous to that of Proposition 1. ∎
+It should be noted that the size of the neighborhood $O$ in this proposition depends only on the cone $K$ and $\text{dist}(\partial R,Q)$.
 $R$$\tilde{K}(Q)$$\tilde{K}^{u}(Q)$$\gamma$
-Figure 6: Illustration of Remark 5.
+Figure 8: Illustration of Proposition 2.
 The rectangle $R$ is centered at $Q$ with respect to $\tilde{K}(Q)$ and has the property that for any curve $\gamma$ that is tangent to $\mathcal{K}^{u}$, any component of $\gamma\cap R$ that intersects the center quadrilateral must stay within the dotted lines and hence not intersect $\partial^{ns}R$.
 4 Main Estimates and Non-Stationary Trace Map Dynamics
-For the rest of the paper, let us consider an arbitrary finite set $\mathcal{A}\subset\mathbb{N}$ and an arbitrary sequence $\overline{a}:=(a_{n})$ such that $a_{n}\in\mathcal{A}$ for all $n\in\mathbb{N}$.
-Let $M>0$. We will show that there is a $\lambda_{0}>0$ such that for all $\lambda\in[0,\lambda_{0}]$, we have
-$\text{dim}_{H}\left(L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}.$ (3.1)
-Our approach is to invoke Lemma 5 and choose a neighborhood $O$, that is composed of small neighborhoods containing the points $Q_{i},$ so that the set
-$C(\lambda,O):=\{P\in\hat{L}_{\lambda}\setminus O:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\}$
-is diffeomorphic to a subset of $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$.
-We will choose the neighborhood $O$ so that the dimension of this set is bigger than the quantity $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ for small $\lambda$.
-More specifically, the neighborhood $O$ is chosen in a way so that for a sufficiently small range of $\lambda$, the set $C(\lambda,O)$ contains a Cantor set of thickness larger than $M$.
-In the next section, we introduce a series of constants that will be used throughout the proof, and these constants will determine the size of the neighborhoods that make up $O$.
+We will now translate Theorem 3 into a problem concerning a sequence of Anosov maps of $\mathbb{T}^{2}$.
+For the rest of this section, let us consider an arbitrary finite set $\mathcal{A}\subset\mathbb{N}$ and an arbitrary sequence $\overline{a}:=(a_{n})$ such that $a_{n}\in\mathcal{A}$ for all $n\in\mathbb{N}$.
+Recalling Lemma 5, given $\rho>0$, we set $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus\tilde{O}_{r(\rho)}$ to be the curve given by $F_{\lambda}^{-1}\left(L_{\lambda}\setminus O_{\rho}\right)$.
+Now, for appropriate range of $\lambda$, we define
+$C(\lambda,O):=\{P\in\hat{L}_{\lambda}\setminus O:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O\ \text{for all}\ n>0\}.$
+Theorem 3 will follow from
+Theorem 4.
+For each $M>0$, there is a $\rho>0$ and $\lambda_{0}\in\left(0,\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}\right)$, such that
+$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (4.1)
+for all $\lambda\in[0,\lambda_{0}]$.
+The proof of this theorem will be developed throughout Section 4 and will be concluded in Subsection 4.8. Let us fix $M>0$.
+4.1 Outline of Proof of Theorem 4
+We outline the proof of the argument to come
+  • In Subsection 4.2, we introduce a series of constants that will be used throughout the proof of this theorem, and these constants will determine $\rho$, and hence the size of the neighborhoods that make up $\tilde{O}_{r(\rho)}$.
+  • Subsection 4.3 is dedicated to choosing $\rho$ explicitly, and once chosen, in Subsection 4.4, we extend each map $\tilde{T}_{\lambda,a}$, for a small range of $\lambda$ and $a\in\mathcal{A}$, to a toral Anosov map, and these maps will satisfy a common cone condition.
+  • In Subsection 4.5, we then derive useful dynamical properties of these maps such as a bounded distortion estimate, and a non-stationary version of the Stable Manifold Theorem for these maps.
+  • In Subsection 4.6, we construct families of sequences of dynamical rectangles containing each component of $\tilde{O}_{r(\rho)}$, that will be useful in controlling the dynamics of our maps near this open set.
+The stable sides of these rectangles will be composed of the stable manifolds constructed in Subsection 4.5.
+  • Lastly, in Subsection 4.7, for each $\lambda$ sufficiently near 00, we construct a subset of $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$, that is roughly composed of the points whose orbits do not enter the interior of the constructed rectangles from Subsection 4.6.
+The proof is concluded in Subsection 4.8 by showing that this subset has large thickness (see Subsection 4.7.1 for a formal definition), and in particular, Hausdorff dimension greater than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
 4.2 Development of Constants
@@
 Recalling (2.2), we fix a cone $K:=K_{\beta_{1}}^{u}$, for some $\beta_{1}\in(\beta_{0},0.1)$, so that it strictly contains $K^{u}$ and is transversal to $K^{s}$.
-We also choose $\beta_{1}$ close enough to $\beta_{0}$ so that the property described in Remark 2 holds for the cone $K$.
+We also choose $\beta_{1}$ close enough to $\beta_{0}$ so that the property described in Lemma 6 holds for the cone $K$.
 If $K(x,y)$ denotes the cone $K$ in the tangent space $T_{(x,y)}\mathbb{T}^{2}$, then we denote $\mathcal{K}:=\left\{K(x,y)\right\}_{(x,y)\in\mathbb{T}^{2}}$.
@@
 $C^{*}:=\frac{C^{*}_{0}}{C_{1}},$ (4.2)
-where the constants $C_{1}$ and $C^{*}_{0}$ are taken from Remarks 3 and 4, respectively. Referring to Lemma 2, we set
+where the constants $C_{1}$ and $C^{*}_{0}$ are taken from Lemma 7 and Proposition 1, respectively. Referring to Lemma 2, we set
 $\mu_{1}:=\overline{\mu}_{\beta_{0}},\ \text{and}\ \mu_{2}:=\max_{n}\left\{\mu^{u}(a_{n})\right\}.$
@@
 for all $a\in\mathcal{A}$.
+Definition.
 We say that a $C^{1}$ diffeomorphism $T:\mathbb{T}^{2}\rightarrow\mathbb{T}^{2}$, satisfies Property (C) if for each $(x,y)\in\mathbb{T}^{2}$, we have
@@
 Certainly, all of our maps $\left\{\hat{T}_{0,a_{n}}\right\}$ satisfy Property (C) by Lemma 2. We set
-$C_{3}:=\sum_{n=-1}^{\infty}\frac{1}{(\mu_{1}-\delta)^{-(n+1)}},$ (3.3)
-Remark 6.
-There is a $\Theta>0$ such that if $(T_{n})$ is a sequence of maps satisfying Property (C) and $\sup_{n}\|T_{n}\|_{C^{2}}<C_{2}+1$, and if $\gamma$ is a smooth curve tangent to $\mathcal{K}^{u}$ so that $\sum_{i=0}^{n}\text{length}\left(T_{n}\circ\cdots\circ T_{1}(\gamma)\right)<C_{1}C_{3}\frac{\sqrt{2}}{10}$ for each $n$, then $\sum_{i=0}^{n}\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)<\Theta$ for each $n$, where $\text{Ang}(\kappa)$ denotes the supremum of the angles between tangent vectors of $\kappa$.
+$C_{3}:=\sum_{n=-1}^{\infty}\frac{1}{(\mu_{1}-\delta)^{-(n+1)}}.$ (4.3)
+Lemma 8.
+There is a $\Theta>0$ such that the following holds: If $(T_{n})$ is a sequence of maps satisfying Property (C) and $\sup_{n}\|T_{n}\|_{C^{2}}<C_{2}+1$, and if $\gamma$ is a smooth curve tangent to $\mathcal{K}^{u}$ so that
+$\sum_{i=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}\text{length}\left(T_{i}\circ\cdots\circ T_{1}(\gamma)\right)<C_{1}C_{3}\frac{\sqrt{2}}{10}$
+for some $m\geq 0$, then
+$\sum_{i=0}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}}\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)<\Theta,$
+where $\text{Ang}(\kappa)$ denotes the supremum of the angles between tangent vectors of $\kappa$.
+Proof.
+Let $\gamma$ be a curve tangent to $\mathcal{K}^{u}$.
+From Property (C) and since the $C^{2}$ norms of each $T_{n}$ are uniformly bounded, the curvature of each curve $T_{i}\circ\dots\circ T_{1}(\gamma)$, $1\leq i\leq m$, is uniformly bounded.
+Hence, there is a constant $\Theta^{\prime}>0$, that only depends on $C_{2}$ and the cones and constants in Property (C), such that
+$\text{Ang}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)\leq\Theta^{\prime}\text{length}\left(T_{i}\circ\dots\circ T_{1}(\gamma)\right)$
+for each $1\leq i\leq m$. The lemma follows.
 Let us define
 $C_{4}:=\exp{\left[(C_{2}+1)^{2}\left(\Theta+C_{1}C_{3}\frac{\sqrt{2}}{10}\right)\right]}.$ (4.4)
-By Lemma 4 we have
+Let us set $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$. By Lemma 4 and Proposition 1 we have
 Lemma 9.
-There is a $l^{*}>0$ such that if $L$ is a line through $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then for each $i=1,\dots,4$, there are segments $L^{+}_{i},\ L^{-}_{i}\subset L$ contained in $\mathbb{B}_{i}\setminus\partial^{s}\mathbb{B}_{i}$ such that
+There is a $l^{*}>0$ such that if $L$ is a line through $Q_{0}$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then for each $i=1,\dots,4$, there are segments $L^{+}_{i},\ L^{-}_{i}\subset L$ contained in $\mathbb{B}_{i}\setminus\partial^{s}\mathbb{B}_{i}$ such that
   • 1.)
@@
     For any $P\in\left(L^{+}_{i}\cup L^{-}_{i}\right)\cap\tilde{K}(Q_{i})$, we have $|P-Q_{i}|<\cos(\theta)\frac{C^{*}}{4}$
-where $\theta$ is the angle between boundary lines of $\tilde{K}(Q_{i})$ (or $K$).
-Let us choose a $\tilde{n}$ so that $\frac{\left(\mu_{1}-\delta\right)^{\tilde{n}}}{10}>l^{*}$ and set $l^{**}:=\left(\mu_{2}+\delta\right)^{\tilde{n}+1}$. Then, again by Lemma 4, we have
+where $\theta$ is the angle between boundary lines of $\tilde{K}(Q_{i})$ (explicitly, $\theta=\theta_{0}(\beta_{1})$, see discussion before (2.2)).
+Proof.
+Let $i\in\{1,2,3,4\}$.
+By Lemma 4, we can choose a $l^{*}>0$ such that if $L$ is a line through $Q_{0}$ of slope $\frac{-1}{[b_{1},b_{2},\dots]}$, $b_{i}\in\mathcal{A}$, and $\text{length}(L)\geq l^{*}$, then there is a segment $L^{\prime}_{i}\subset L$ such that
+  • $L^{\prime}_{i}\subset\tilde{K}_{+}(Q_{i})$ and connects the boundaries of $\tilde{K}_{+}(Q_{i})$
+  • For each $P\in L_{i}^{\prime}$, we have $|P-Q_{i}|\leq\min\left\{C_{0}^{*},\ \cos(\theta)\frac{C^{*}}{4}\right\}.$
+By increasing $l^{*}$ if needed, we can even assume that $L^{\prime}_{i}$ connects the boundary lines of $\tilde{K}_{+}(Q_{i})$.
+Let $L$ be such a line and let $L^{+}_{i}$ be the longest segment of $L\cap\overline{\mathbb{B}}_{i}$ that contains $L^{\prime}_{i}$. Then, by Proposition 1, since
+$\text{dist}\left(L_{i}^{+},Q_{i}\right)<C_{0}^{*},$
+we have that $L_{i}^{+}\cap\partial^{s}\mathbb{B}_{i}=\emptyset.$ By increasing $l^{*}$ again if needed, we can also ensure that $L_{i}^{+}$ connects the components of $\partial^{u}\mathbb{B}_{i}$ and Proposition 1 will still guarantee that this extended segment is disjoint from $\partial^{s}\mathbb{B}_{i}$.
+By repeating this argument for each $i$ and with $+$ replaced with --, the lemma is proved.
+Let us choose a $\tilde{n}$ so that $\frac{\left(\mu_{1}-\delta\right)^{\tilde{n}}}{10}>l^{*}$ and set $l^{**}:=\left(\mu_{2}+\delta\right)^{\tilde{n}+1}$. For $m\geq 1$, denote
+$\text{Orb}_{0}\left(Q_{0}\right):=\left\{\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(Q_{0}\right)\right\}_{n\geq 0}\cup\left\{Q_{0}\right\}.$
+It is a calculation to check that $\text{Orb}_{0}\left(Q_{0}\right):=\bigcup_{m\geq 0}\text{Orb}_{0}\left(Q_{0}\right)$ is a finite set and has distance from each $Q_{i}$ that is at least $\frac{1}{4}$.
+Again by Lemma 4, we have
 Lemma 10.
-There is a $C^{**}\in\left(0,\frac{1}{4}-\frac{1}{10}\right)$ such that if $L$ is a line through $Q_{0}$ of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\leq l^{**}$, then
+There is a $C^{**}\in\left(0,\frac{1}{4}-\frac{1}{10}\right)$ such that if $L$ is a line through $Q$, for some $Q\in\text{Orb}_{0}\left(Q_{0}\right)$, of slope $-\frac{1}{[b_{1},b_{2},\dots]}$ with $b_{k}\in\mathcal{A}$, and $\text{length}(L)\leq l^{**}$, then
 $\text{dist}(L,Q_{i})\geq C^{**},$ (4.5)
 for each $i=1,\dots,4$.
-Recalling our constant $C_{1}$ from Remark 3, we will choose $N^{*}\in\mathbb{N}$ such that
+Proof.
+The proof is analogous to that of Lemma 9 but through an application Lemma 4 part (2).
+The quantity $\frac{1}{4}-\frac{1}{10}$ comes from the distance between $\text{Orb}_{0}\left(Q_{0}\right)$ and each $Q_{i}$, and the size of each $\mathbb{B}_{i}$.
+Recalling our constant $C_{1}$ from Lemma 7, we will choose $N^{*}\in\mathbb{N}$ such that
 $C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-N^{*}}}-C_{4}^{-1}\\ \geq M$ (4.6)
@@
 $C_{1}(\mu_{1}-\delta)^{-N^{*}}C^{*}\leq\frac{C^{**}}{4}.$ (4.7)
-3.2 Translation to a Toral Map Problem
+4.3 Choice of Neighborhood
 We are now ready to choose the neighborhood $O$ that was described at the beginning of this section.
-From Remark 5, there is a $r_{0}=r_{0}\left(N^{*},C^{**},K\right)$ such that
+From Proposition 2, there is a $r_{0}=r_{0}\left(N^{*},C^{**},K\right)$ such that
 $r_{0}<(\mu_{2}+\delta)^{-N^{*}-1}\frac{C^{**}}{4},$ (4.8)
@@
 Referring to Lemma 5, let us choose $\rho>0$ so small that $r(\rho)<\frac{r_{0}}{2}$ and set $\lambda_{1}:=\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$.
-We see that for $\lambda\in[0,\lambda_{1}]$, the set
-$\left\{\boldsymbol{x}\in L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}:T_{a_{n}}\circ\cdots\circ T_{a_{1}}(\boldsymbol{x})\not\in O_{\rho}\ \text{for all}\ n>0\right\}$ (3.9)
-is a one-dimensional compact set that is contained in $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$.
-It is a straightforward calculation to check that the line $\hat{L}_{0}:=\left[0,\frac{1}{2}\right]\times\{0\}\subset\mathbb{T}^{2}$ is mapped into $L_{0}\cap\mathbb{S}$ under the map $F$.
+For $\lambda\in(0,\lambda_{1})$, it is a straightforward calculation to check that the line $\hat{L}_{0}:=\left[0,\frac{1}{2}\right]\times\{0\}\subset\mathbb{T}^{2}$ is mapped into $L_{0}\cap\mathbb{S}$ under the map $F$.
 Also, the projection map $\pi_{\lambda}$ sends $L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}$ onto a curve in $\mathbb{S}$ that is $C^{2}$ close to a segment of $L_{0}$, provided $\lambda$ is small.
-That is, for small $\lambda$, there is a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O_{r(\rho)}$, that is $C^{2}$ close to a segment of $\hat{L}_{0}$, that is mapped into $L_{\lambda}$ under the map $F_{\lambda}$, and this curve tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ as $\lambda\rightarrow 0$.
-As the map $F_{\lambda}$ preserves Hausdorff dimension, it suffices to estimate the Hausdorff dimension of the set
-$\mathcal{C}\left(\lambda,O_{r(\rho)}\right):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O_{r(\rho)}\ \text{for all}\ n>0\right\}.$ (3.10)
-That is, it suffices to show that there is a $\lambda_{0}\in(0,\lambda_{1})$ such that
-$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,O_{r(\rho)}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (3.11)
+That is, for small $\lambda$, the curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ is $C^{2}$ close to a segment of $\hat{L}_{0}$, and in particular, $\hat{L}_{0}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ as $\lambda\rightarrow 0$ in the $C^{2}$-sense.
+So, to complete the proof of Theorem 4, we will show that there is a $\lambda_{0}\in(0,\lambda_{1})$ such that
+$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (4.9)
 for all $\lambda\in[0,\lambda_{0}]$.
-3.3 Extended Torus Maps
+4.4 Extended Toral Maps
 Now, for each $a\in\mathcal{A}$ and $\lambda\in[0,\lambda_{1}]$, we will construct an Anosov diffeomorphism that coincides with $\tilde{T}_{\lambda,a}$ outside of $O_{r_{0}}$.
@@
 $\tilde{T}_{0,a}(x,y)=\left(\tilde{T}_{0,a}^{1}(x,y),\tilde{T}_{0,a}^{2}(x,y)\right)\ \text{and}\ \tilde{T}_{\lambda,a}(x,y)=\left(\tilde{T}_{\lambda,a}^{1}(x,y),\tilde{T}_{\lambda,a}^{2}(x,y)\right)$
-where $\tilde{T}_{0,a}^{1}(x,y)=ax+y$, $\tilde{T}_{0,a}^{2}(x,y)=x$, and $\tilde{T}_{\lambda,a}^{1}(x,y),\ \tilde{T}_{\lambda,a}^{2}(x,y):\mathbb{T}^{2}\setminus O_{r(\rho)}\rightarrow S^{1}$, we set
+where $\tilde{T}_{0,a}^{1}(x,y)=ax+y$, $\tilde{T}_{0,a}^{2}(x,y)=x$, and $\tilde{T}_{\lambda,a}^{1}(x,y),\ \tilde{T}_{\lambda,a}^{2}(x,y):\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\rightarrow S^{1}$, we set
 $\hat{T}_{\lambda,a}(x,y):=\left(\hat{T}_{\lambda,a}^{1}(x,y),\hat{T}_{\lambda,a}^{2}(x,y)\right),$
@@
 To ease notation, we will denote $f_{\lambda,n,m}:=\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}$, for $n,m\geq 0$, where $\hat{T}_{\lambda,0}:=\text{id}$.
-Remark 7.
+Lemma 12.
+For each $\lambda\in\left[0,\lambda_{1}\right]$, we have
+$\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)$ (4.10)
+for any $m\geq 0$ and $n\in\mathbb{N}$.
+Proof.
 Based on the compositions from (1.2) and construction of $\pi_{\lambda}$ in Lemma 5, one can check that for any $\lambda\in[0,\lambda_{1}]$, the set
@@
 $\left(\pi_{\lambda}\circ T_{a_{n}}\circ\cdots\circ T_{a_{1}}\right)\left(-\sqrt{1+\frac{\lambda^{2}}{4}},0,0\right)=T_{a_{n}}\circ\cdots\circ T_{a_{1}}(-1,0,0).$
-Since $F\left(\frac{1}{4},\frac{1}{4}\right)=(-1,0,0)$, and $\left(\frac{1}{4},\frac{1}{4}\right)\in\mathbb{T}^{2}\setminus O_{r_{0}}$ (see 3.8), this gives
+Since $F\left(\frac{1}{4},\frac{1}{4}\right)=(-1,0,0)$, and $\left(\frac{1}{4},\frac{1}{4}\right)\in\mathbb{T}^{2}\setminus O_{r_{0}}$ (see 4.8), this gives
 $\hat{T}_{\lambda,a_{n}}\circ\cdots\circ\hat{T}_{\lambda,a_{1}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n}}\circ\cdots\circ\hat{T}_{0,a_{1}}\left(\frac{1}{4},\frac{1}{4}\right)$
-for any $n\in\mathbb{N}$, and this non-stationary orbit is a finite set. It is similar to check that
-$\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)$
-for any $m\geq 0$ and $n\in\mathbb{N}$.
+for any $n\in\mathbb{N}$, and this non-stationary orbit is a finite set. This same argument gives (4.10) for any $m\geq 0$. ∎
+4.4.1 Further Choice of Constants.
 Now, one may choose a $\lambda_{2}\in(0,\lambda_{1})$ such that for any $\lambda\in[0,\lambda_{2}]$ and $n\in\mathbb{N}$, we have
@@
 The third part comes from the fact that $A_{a}$ has a positive unstable eigenvalue for each $a$, so if $\hat{T}_{\lambda,a}$ is sufficiently $C^{1}$-close to $\hat{T}_{0,a}$, then the differential of $\hat{T}_{\lambda,a}$ will preserve orientation of vectors from the cone $K$.
-The last part comes from the fact that the curve $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ in the $C^{2}$-sense.
-From these cone conditions, we have
+The last part comes from the fact that the curve $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ in the $C^{2}$-sense.
+4.5 Properties of the Extended Toral Maps
+This subsection is dedicated to deriving useful dynamical properties of our extended maps $\left\{\hat{T}_{\lambda,a}\right\}$.
+4.5.1 Lengths and Orientation of Curves
+From Property (C), we have
 Lemma 13.
@@
 Lemma 14.
-For each $\lambda\in[0,\lambda_{2}]$, if $\gamma$ is a sub-curve of $\hat{L}_{\lambda}$ that has endpoints $P$ and $Q$ and satisfies $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$ for some $i=1,\dots,4$, then for any point $Q^{\prime}\in\gamma\setminus\{P,Q\}$, we must have
+For each $\lambda\in[0,\lambda_{2}]$, if $\gamma$ is a sub-curve of $\hat{L}_{\lambda}$ that has endpoints $P$ and $Q$ and satisfies $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$ for some $i=1,\dots,4$ and $m\geq 0$, then for any point $Q^{\prime}\in\gamma\setminus\{P,Q\}$, we must have
 $C_{4}^{-1}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})}$ $\leq\frac{\text{dist}_{f_{\lambda,m,0}(\gamma)}\left(f_{\lambda,m,0}(P),f_{\lambda,m,0}(Q^{\prime})\right)}{\text{dist}_{f_{\lambda,m,0}(\gamma)}\left(f_{\lambda,m,0}(Q),f_{\lambda,m,0}(Q^{\prime})\right)}$
-$\leq C_{4}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})},$
-for any $m\geq 0$.
+$\leq C_{4}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})}.$
 Proof.
-Let $\lambda\in[0,\lambda_{2}]$ and $m\geq 0$.
-Since $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$, it must have length less than $C_{1}\frac{\sqrt{2}}{10}$.
+Let $\lambda\in[0,\lambda_{2}]$.
+If $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$, it must have length less than $C_{1}\frac{\sqrt{2}}{10}$.
 Thus, from Lemma 13 and recalling (4.3), it must be the case that for each $k$ with $0\leq k\leq m$, we have
@@
 $\sum_{k=0}^{m}\text{length}(f_{\lambda,k,0}(\gamma))\leq C_{1}\frac{\sqrt{2}}{10}\sum_{k=0}^{m}(\mu_{1}-\delta)^{-(m-k)}\leq C_{1}C_{3}\frac{\sqrt{2}}{10}.$
-In addition, since we have that $\hat{L}_{\lambda}$ is a smooth curve tangent to $\mathcal{K}^{u}$ and $\left\|\hat{T}_{\lambda,a_{n}}\right\|\leq C_{2}+1$ for all $n\in\mathbb{N}$, it follows from Remark 6 that
-$\sum_{k=0}^{m}\text{Ang}\left(f_{\lambda,k,0}\right)<\Theta.$
+In addition, since we have that $\hat{L}_{\lambda}$ is a smooth curve tangent to $\mathcal{K}^{u}$ and $\left\|\hat{T}_{\lambda,a_{n}}\right\|\leq C_{2}+1$ for all $n\in\mathbb{N}$, it follows from Lemma 8 that
+$\sum_{k=0}^{m}\text{Ang}\left(f_{\lambda,k,0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\gamma\right)}\right)<\Theta.$
 Thus, using Theorem 2 of [BM], the result holds for $C_{4}$ as defined in (4.4). ∎
-3.4 Non-Stationary Stable Manifolds
-In this section, we establish a non-stationary version of the Stable Manifold Theorem for our collection of maps.
+4.5.3 Non-Stationary Stable Manifolds
+In this subsection, we establish a non-stationary version of the Stable Manifold Theorem for our collection of maps.
 Consider the set
-$W^{s}_{\lambda}\left(Q,m\right):=\left\{P\in\mathbb{T}^{2}:\lim\limits_{n\rightarrow\infty}\text{dist}\left(f_{\lambda,n,m}(P),f_{\lambda,n,m}(Q)\right)=0\right\},$
+$W^{s}_{\lambda}\left(Q,m\right):=\left\{P\in\mathbb{T}^{2}:\lim\limits_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|f_{\lambda,n,m}(P)-f_{\lambda,n,m}(Q)\right|}=0\right\},$
 which we call the non-stationary stable manifold of $\left(T_{a_{n+m}}\right)_{n\geq 1}$ at $Q$. For $\eta>0$, we define
@@
     For each $\epsilon>0$, there is a $\lambda^{\prime}(\epsilon)>0$ such that if $\lambda<\lambda^{\prime}(\epsilon)$, then
-    $\text{dist}\left(W^{s}_{0,\text{loc}}(Q,k),W^{s}_{\lambda,\text{loc}}(Q,k)\right)<\epsilon$
+    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(W^{s}_{0,\text{loc}}(Q,k),W^{s}_{\lambda,\text{loc}}(Q,k)\right)<\epsilon$
     for all $k\in\mathbb{N}$ and any point $Q$ satisfying $f_{\lambda,n,k}(Q)=f_{0,n,k}(Q)$ for any $n,k\geq 0$ and each $\lambda$.
@@
 We prove the result first for $Q=0$.
-Recalling our notations from Section 2.3.1, we work locally in the neighborhood $\mathbb{B}_{1}=\pi\left(\text{Box}_{\frac{1}{10}}(0)\right)$ and in the coordinate system $\tilde{V}^{\perp}\oplus\tilde{V}$ where $\tilde{V}=\pi\left(V\cap\text{Box}_{\frac{1}{10}}(0)\right)$
+Recalling our notations from Section 3.1, we work locally in the neighborhood $\mathbb{B}_{1}=\pi\left(\text{Box}_{\frac{1}{10}}(0)\right)$ and in the coordinate system $\tilde{V}^{\perp}\oplus\tilde{V}$ where $\tilde{V}=\pi\left(V\cap\text{Box}_{\frac{1}{10}}(0)\right)$
 Notice that there is a $l<1$ such that $\varphi:\tilde{V}^{\perp}\rightarrow\tilde{V}$ is tangent to $\mathcal{K}^{s}$ if and only if $\text{Lip}(\varphi)<l$.
@@
 $d_{X}\left(\text{graph}(\varphi),\text{graph}(\varphi^{\prime})\right):=d_{C^{0}}(\varphi,\varphi^{\prime}).$
-From Remark 2 and the cone conditions, we may choose a $\lambda_{3}>0$ such that for each $\lambda\in[0,\lambda_{3}]$ and each $a\in\mathcal{A}$, if $\text{graph}(\varphi)\in X$, then its image under $\hat{T}_{\lambda,a}^{-1}$ must overflow $\mathbb{B}_{1}$ and be tangent to $\mathcal{K}^{s}$.
+From Lemma 6 and the cone conditions, we may choose a $\lambda_{3}>0$ such that for each $\lambda\in[0,\lambda_{3}]$ and each $a\in\mathcal{A}$, if $\text{graph}(\varphi)\in X$, then its image under $\hat{T}_{\lambda,a}^{-1}$ must overflow $\mathbb{B}_{1}$ and be tangent to $\mathcal{K}^{s}$.
 For each $\lambda\in[0,\lambda_{3}]$ and $n\in\mathbb{N}$, let us define $\Gamma_{\lambda,a_{n}}:X\rightarrow X$ so that $\Gamma_{\lambda,a_{n}}\left(\text{graph}(\varphi)\right)$ is the segment of $\hat{T}_{\lambda,a_{n}}^{-1}\left(\text{graph}(\varphi)\right)\cap\mathbb{B}_{1}$ that contains 00.
@@
 Through the common cone conditions (Property (C)), for each $\lambda\in[0,\lambda_{3}]$, we find that $\left(\Gamma_{\lambda,a_{n}}\right)$ is a sequence of contractions with uniform contraction rate over $n$.
-Thus, by Proposition 1, for each $k\in\mathbb{N}$, there is a $\varphi_{\lambda,k}^{*}:V^{\perp}\rightarrow V$ such that $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)\in X$ and
+Thus, by Proposition A.1, for each $k\in\mathbb{N}$, there is a $\varphi_{\lambda,k}^{*}:V^{\perp}\rightarrow V$ such that $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)\in X$ and
 $\lim\limits_{n\rightarrow\infty}\Gamma_{\lambda,a_{k}}\circ\cdots\circ\Gamma_{\lambda,a_{n+k}}\text{graph}\left(\varphi\right)=\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$
@@
 To establish (2), we notice that the distance between $\Gamma_{0,n}$ and $\Gamma_{\lambda,n}$ in $(X,d_{X})$, can be made arbitrarily small for a sufficiently small range of $\lambda$ due to the fact that $\hat{T}_{\lambda,a_{n}}$ is a $C^{2}$ perturbation of $\hat{T}_{0,a_{n}}$, and this smallness can be made uniform in $n$.
-Hence, by Proposition 1, we can make the corresponding fixed points $\text{graph}\left(\varphi_{0,k}^{*}\right)$ and $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ arbitrary close in the metric $d_{X}$ for sufficiently small $\lambda$, and this closeness can be made uniform in $k$.
+Hence, by Proposition A.1, we can make the corresponding fixed points $\text{graph}\left(\varphi_{0,k}^{*}\right)$ and $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ arbitrary close in the metric $d_{X}$ for sufficiently small $\lambda$, and this closeness can be made uniform in $k$.
 For the case $Q\neq 0$, we repeat the process but for the family of maps given by
@@
 For each $\epsilon>0$ and $n_{0}\in\mathbb{N}$, there is a $\lambda^{\prime}\in\left(0,\lambda_{3}\right)$ such that if $\lambda\in\left[0,\lambda^{\prime}\right]$, we have that
-$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq n_{0}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)<\epsilon,$
+$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq n_{0}\\ m\in\mathbb{N}\end{subarray}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)<\epsilon,$
 for all $Q\in\text{Orb}_{\lambda}(Q_{0})$.
@@
 4.6 Rectangle Construction
-Recall that by Remark 7, for $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$, we have that
+We will now construct a specific family of dynamical rectangles that contain each neighborhood of $O_{r_{0}}$. Recall that by Lemma 12, for $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$, we have that
 $f_{\lambda,n,m}(Q_{0})=f_{0,n,m}(Q_{0})$
-for each $n,m\geq 0$ and $\lambda\in[0,\lambda_{2}]$, and it is also a calculation to see that $\text{Orb}_{0}(Q_{0})$ is a finite set and bounded away from each $Q_{i}$ by at least $\frac{1}{4}$.
+for each $n,m\geq 0$ and $\lambda\in[0,\lambda_{2}]$, and as pointed out before Lemma 10, $\text{Orb}_{0}(Q_{0})$ is a finite set and bounded away from each $Q_{i}$ by at least $\frac{1}{4}$.
 Now, for each $i=1,\dots,4$, $m\in\mathbb{N}$, and $\lambda\in[0,\lambda_{3}]$, we will construct a family of rectangles containing $Q_{i}$ that have stable sides belonging to $W^{s}_{\lambda}(Q_{0},m)$ and also have controlled size.
@@
 This means that for future iterates, the points that intersected the stable boundaries of $R$ will move away from the points $Q_{i}$.
-This dynamical behavior will allow us to construct a Cantor set contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has large thickness.
+This dynamical behavior will allow us to construct a Cantor set contained in $\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ that has Hausdorff dimension greater than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
 4.6.1 Base Set of Rectangles
 The construction of our rectangles will be done in an inductive process, so we start by constructing a base set of rectangles.
-Recalling the constant $C^{*}$ from (3.2) and $C^{**}$ from Lemma 7, we have
+Recalling the constant $C^{*}$ from (4.2) and $C^{**}$ from Lemma 10, we have (see Figure 9)
 Lemma 15.
@@
   • 1.)
-    $\partial^{s}R_{m}(\lambda,Q_{i})\subset W^{s}_{\lambda}(Q_{0},m)$ and in particular, +$\partial^{s}_{+}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{+}(Q_{i})$ and $\partial^{s}_{-}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{-}(Q_{i})$
+    $\partial^{s}R_{m}(\lambda,Q_{i})\subset W^{s}_{\lambda}(Q_{0},m)$ and in particular, +$\partial^{s}_{+}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{+}(Q_{i})$ and $\partial^{s}_{-}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{-}(Q_{i})$.
   • 2.)
-    For any $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$
+    For any $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$.
   • 3.)
@@
     $\min\left\{\text{dist}\left(P,\bigcup_{i=1}^{4}\{Q_{i}\}\right),\text{dist}\left(f_{\lambda,n,m}(P),\bigcup_{i=1}^{4}\{Q_{i}\}\right)\right\}\geq\frac{C^{**}}{2}$
-    for any $n\geq 0$
+    for any $n\geq 0$.
 Proof.
@@
 Step 1: We first construct the stable sides of the rectangles for $\lambda=0.$
-We know that $W^{s}_{0}(Q_{0},m)$ is a line through $Q_{0}$ of slope $-\frac{1}{[a_{m},a_{m+1},\dots]}$ and we know from Lemma 9 that
+We know that $W^{s}_{0}(Q_{0},m)$ is a line through $Q_{0}$ of slope $-\frac{1}{[a_{m},a_{m+1},\dots]}$ (see Lemma 3) and we know from Lemma 13 that
 $\text{length}\left(f^{-1}_{0,\tilde{n},m}\left(W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),m+\tilde{n}+1)\right)\right)\geq l^{*},$
@@
 from Lemma 13.
-Notice that $f_{0,\tilde{n},m}(P)\in W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)$, and so if $n\geq\tilde{n}$, then $f_{\lambda,n,m}(P)\in B_{\frac{1}{10}}(Q_{0})$ so that its distance from $\bigcup_{i=1}^{4}\{Q_{i}\}$ is greater than $\frac{3}{20}\frac{1}{4}-\frac{1}{10}$.
+Notice that $f_{0,\tilde{n},m}(P)\in W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)$, and so if $n\geq\tilde{n}$, then $f_{\lambda,n,m}(P)\in B_{\frac{1}{10}}(Q_{0})$ so that its distance from $\bigcup_{i=1}^{4}\{Q_{i}\}$ is greater than $\frac{3}{20}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=}\frac{1}{4}-\frac{1}{10}$.
 If $n<\tilde{n}$, then $f_{0,n,m}(P)\in f^{-1}_{0,\tilde{n}-n,m+n+1}W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),\tilde{n}+m+1)$, where $f^{-1}_{0,\tilde{n}-n,m+n+1}W^{s}_{0,\text{loc}}(f_{0,\tilde{n},m}(Q_{0}),\tilde{n}+m+1)$ is a line segment contained in $W^{s}_{0}(f_{0,n,m}(Q_{0}),m+n+1)$ that has slope $-\frac{1}{[a_{m+n+1},a_{m+n+2},\dots]}$, contains $f_{0,n,m}(Q_{0})\in\text{Orb}_{0,m}(Q_{0})$, and has length no greater than $\frac{(\mu_{2}+\delta)^{(\tilde{n}-n)+1}}{10}<l^{**}$.
@@
 […]
-From Corollary 1, the quantity
-$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq\tilde{n}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)$ (3.17)
+From Corollary 1 and Lemma 12, the quantity
+$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq\tilde{n}\\ m\in\mathbb{N}\end{subarray}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)$ (4.16)
 can be made arbitrarily small for a sufficiently small range of $\lambda$ for all $Q\in\text{Orb}_{\lambda}(Q_{0})$.
@@
 We will now use this base set of rectangles to construct dynamical rectangles of small controlled size.
-We will need rectangles of smaller scale to ensure that the Cantor set of $\mathcal{C}(\lambda,O_{r(\rho)})$, that we construct later on, will have small gaps in comparison to their bridges.
+We will need rectangles of smaller scale to ensure that the sub-Cantor set of $\mathcal{C}(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}})$, that we construct later on, will be thick and hence have Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
 The construction will come from taking the stable boundaries of rectangles from the collection $\left\{R_{\lambda}(m,Q_{i})\right\}$, and iterating them under the maps $f^{-1}_{\lambda,n,m}$, which will give two new curves tangent to $\mathcal{K}^{s}$ that have segments which are close together and centered at some $Q_{i}$, and these curves will serve as the stable boundaries for a new rectangle.
@@
     If $\kappa$ is a curve contained in $\mathbb{B}_{i}$ and tangent to $\mathcal{K}^{u}$ that connects $\partial_{+}^{s}R^{\prime}_{m}(\lambda,Q_{i},N)$ and $\partial_{-}^{s}R^{\prime}_{m}(\lambda,Q_{i},N)$, then
-    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}C^{*}$
+    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}C^{*}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$
 […]
-    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}\frac{C^{*}}{2}$
+    $\text{length}(\kappa)\leq C_{1}(\mu_{1}-\delta)^{-N}\frac{C^{*}}{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$
   • 3)
@@
     The non-stable boundaries $\partial^{ns}R^{\prime}_{m}(\lambda,Q_{i},N)$ of $R^{\prime}_{m}(\lambda,Q_{i},N)$ are line segments with endpoints in $\partial\tilde{K}(Q_{i})$.
-That is, $R^{\prime}_{m}(\lambda,Q_{i},N)$ has vertices belonging to $\partial\tilde{K}(Q_{i})$, and does not intersect $\tilde{K}(Q_{i})$ otherwise.
+That is, $R^{\prime}_{m}(\lambda,Q_{i},N)$ has vertices belonging to $\partial\tilde{K}(Q_{i})$, and does not intersect ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial}\tilde{K}(Q_{i})$ otherwise.
 Proof.
@@
 Consider the rectangle $R^{\prime}_{m+1}\left(\lambda,Q_{j},N-1\right)$, where $Q_{j}=f_{\lambda,0,m}(Q_{i})$, which is centered at $Q_{j}$ with respect to $\tilde{K}(Q_{j})$.
-Notice from (3.13) and Remark 2, provided $\lambda$ is small, the image of the cone $\tilde{K}(Q_{i})$, under the map $f_{\lambda,0,m}$, overflows the box $\mathbb{B}_{j}$ (see Figure 8).
+Notice from (4.12) and Lemma 6, provided $\lambda$ is small, the image of the cone $\tilde{K}(Q_{i})$, under the map $f_{\lambda,0,m}$, overflows the box $\mathbb{B}_{j}$ (see Figure 10).
 Without loss of generality, let us assume that this holds for all $\lambda\in[0,\lambda_{4}].$
@@
 4.7 Gap Estimates and Thickness
-In order to establish (3.11), we will show that for each $\lambda\in[0,\lambda_{4}]$, there is a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has thickness larger than $M$.
+In order to establish (4.9), we will show that for each $\lambda\in[0,\lambda_{4}]$, there is a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ contained in $\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ that has Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
+4.7.1 Thickness of Cantor Sets
+A fundamental tool we will use to estimate the Hausdorff dimension of this set is the notion of thickness for Cantor sets.
+Given a Cantor set $\mathcal{C}\subset\mathbb{R}$ with convex hull $I$, we say that $U\subset I\setminus\mathcal{C}$ is a gap of $\mathcal{C}$ if it is a connected component of $I\setminus\mathcal{C}$.
+Any enumeration $\mathcal{U}=\{U_{n}\}$ of the gaps of $\mathcal{C}$ is called a presentation of $\mathcal{C}$.
+Given a boundary point $u\in\mathcal{C}$ of some gap $U_{k}$, we call the connected component $B$ contained in $I\setminus\left(U_{1}\cup\dots\cup U_{k}\right)$ that contains $u$ a bridge of $U_{k}$.
+We define
+$\tau(\mathcal{C},\mathcal{U},u):=\frac{|B|}{|U_{k}|},$
+and set the thickness of $\mathcal{C}$ to be
+$\tau(\mathcal{C}):=\sup\limits_{\mathcal{U}}\inf\limits_{u}\tau(\mathcal{C},\mathcal{U},u).$
+It is related to the Hausdorff dimension of $\mathcal{C}$ by the inequality
+$\text{dim}_{H}(\mathcal{C})\geq\frac{\log 2}{\log\left(2+\frac{1}{\tau(\mathcal{C})}\right)}.$ (4.18)
+(see Chapter 4 of [PT]).
+It is clear that these notions carry over to a Cantor set that belongs to a $C^{1}$ curve $\gamma\subset\mathbb{T}^{2}$ where the lengths of gaps and bridges are measured using length relative to the curve $\gamma$.
+4.7.2 Ordering of $\hat{L}_{\lambda}$
 For each curve $\hat{L}_{\lambda}$, as it is close to a segment of $\hat{L}_{0}=\left[0,\frac{1}{2}\right]\times\{0\}$ and tangent to $\mathcal{K}^{u}$, for our range of $\lambda$, we assign a natural ordering to its elements as follows: Let $P^{*}$ be the endpoint of $\hat{L}_{\lambda}$ that is closest to $(0,0)$ and $P^{**}$ its other endpoint.
@@
 Recall from (4.7) we know $R^{\prime}_{k}(\lambda,Q_{i},N^{*})\subset B_{\frac{C^{**}}{4}}(Q_{i})$ for all $i=1,\dots,4$ and $k\in\mathbb{N}$.
-We will construct a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ that is contained in $\mathcal{C}\left({\lambda},O_{r(\rho)}\right)$ and show that it has thickness bigger than $M$:
+We will construct a Cantor set $\mathcal{C}\left(\lambda,N^{*}\right)$ that is contained in $\mathcal{C}\left({\lambda},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ and show that it has thickness bigger than $M$:
 Consider the line $\hat{L}_{\lambda}$ and recall that it is tangent to $\mathcal{K}^{u}$ for any $\lambda\in[0,\lambda_{4}]$.
@@
 Then, $J$ is the image of some segment $I_{0}^{\prime}\subset I_{0}$ under the map $f_{\lambda,k_{0},0}$.
-We will say that $\text{int}I^{\prime}_{0}$ is a gap of order $k_{0}$, where the interior is taken with respect to the subspace topology on $\hat{L}_{\lambda}$.
+We will say that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{int}\left(I^{\prime}_{0}\right)}$ is a gap of order $k_{0}$, where the interior is taken with respect to the subspace topology on $\hat{L}_{\lambda}$.
 Since $R^{\prime}_{k_{0}+1}\left(\lambda,Q_{j},N^{*}\right)$ is centered at $Q_{j}$ with respect to $\tilde{K}(Q_{j})$ and
-$f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is a curve tangent to $\mathcal{K}^{u}$, by Remark 5 and our discussion after Lemma 12, as $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)\cap O_{r_{0}}(Q_{j})\neq\emptyset$, we must have that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is fully contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and only intersects $\partial R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ non-trivially in the event that either of its endpoints belong to $\partial^{s}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
+$f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is a curve tangent to $\mathcal{K}^{u}$, by Proposition 2 and our discussion after Lemma 16, as $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)\cap O_{r_{0}}(Q_{j})\neq\emptyset$, we must have that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ is fully contained in $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and only intersects $\partial R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ non-trivially in the event that either of its endpoints belong to $\partial^{s}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
 However, from (4.19), and since we chose $N^{*}$ so large that $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})\subset B_{\frac{C^{**}}{4}}$ (see (4.7)), it must be the case that $f_{\lambda,k_{0},0}\left(I_{0}^{\prime}\right)$ indeed connects the stable boundaries of $R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
 In particular, by Lemma 17, if $I_{0}^{\prime}$ has endpoints $P_{1}$ and $P_{2}$ with $P_{1}<P_{2}$, then we have $f_{\lambda,k_{0},0}(P_{1})\in\partial^{s}_{-}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$ and $f_{\lambda,k_{0},0}(P_{2})\in\partial^{s}_{+}R^{\prime}_{k_{0}+1}(\lambda,Q_{j},N^{*})$.
-Suppose that we have constructed all of the gaps up to order $k\geq k_{0}$, and that they are all disjoint segments of $\hat{L}_{\lambda}$ of the form $\text{int}I_{m}$ where $I_{m}\subset I_{0}$ for some $m\leq k$ such that:
+Suppose that we have constructed all of the gaps up to order $k\geq k_{0}$, and that they are all disjoint segments of $\hat{L}_{\lambda}$ of the form ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{int}\left(I_{m}\right)}$ where $I_{m}\subset I_{0}$ for some $m\leq k$ such that:
   • $f_{\lambda,k,0}\left(I_{m}\right)$ is a curve contained in $R^{\prime}_{m+1}(\lambda,Q_{i},N^{*})$, for some $i=1,\dots,4$
@@
 The gaps of order $k+1$ will come from taking the image of such components that intersects $\bigcup_{i=1}^{4}O_{r_{0}}(Q_{i})$ non-trivially, under the map $f^{-1}_{\lambda,k+1,0}$.
-A similar application of Remark 5 and Lemma 13 as in the base case shows that any gap of order $k+1$ will be of the form above with index $m$ replaced with $k+1$.
+A similar application of Proposition 2 and Lemma 17 as in the base case shows that any gap of order $k+1$ will be of the form above with index $m$ replaced with $k+1$.
 We have now constructed all orders of gaps of $\mathcal{C}\left(\lambda,N^{*}\right)$ and shown they have satisfied a certain form. From (4.8) and definition of $r_{0}$, we know
@@
 for all $k\geq 0$, $i=1,\dots,4$ and $\lambda\in[0,\lambda_{4}]$. Thus, it follows that
-$C\left(\lambda,N^{*}\right)\subset C\left(\lambda,O_{r(\rho)}\right).$ (3.20)
-3.6.2 Thickness Estimate
+$C\left(\lambda,N^{*}\right)\subset C\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right).$ (4.20)
+4.8 Concluding the Proof of Theorem 4
+With the construction of the set $\mathcal{C}\left(\lambda,N^{*}\right)$, we are now ready to complete the proof of Theorem 4.
+Proof of Theorem 4..
 Let $\lambda_{0}=\min\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ and let $\lambda\in[0,\lambda_{0}]$. We will show that $C\left(\lambda,N^{*}\right)$ has thickness larger than $M$.
@@
 and the segment $J\subset f_{\lambda,k,0}(B\cup U)$ connecting $f_{\lambda,0,k}(Q)$ and $f_{\lambda,k,0}(P)$ is contained in $\mathbb{B}_{i}$.
-Proof.
+Proof of Lemma 18..
 By the construction of our gaps, the endpoint $P^{\prime\prime}$ of $B$ satisfies $f_{\lambda,k^{\prime},0}\left(P^{\prime\prime}\right)\in\partial^{s}R^{\prime}_{k^{\prime}+1}(\lambda,Q_{j},N^{*})$ for some $j=1,\dots,4$ and $k^{\prime}\leq k$.
@@
 $\frac{\text{length}(B)}{\text{length}(U)}=\frac{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime\prime}\right)}{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime}\right)}\geq\frac{\text{dist}_{\hat{L}_{\lambda}}\left(P,Q\right)}{\text{dist}_{\hat{L}_{\lambda}}\left(P,P^{\prime}\right)}$
-$\geq C_{4}^{-1}\frac{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(Q\right)\right)}{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)}\geq C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-N}}-C_{4}^{-1}$
+$\geq C_{4}^{-1}\frac{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(Q\right)\right)}{\text{dist}_{f_{\lambda,k,0}\left(\hat{L}_{\lambda}\right)}\left(f_{\lambda,k,0}(P),f_{\lambda,k,0}\left(P^{\prime}\right)\right)}\geq C_{4}^{-1}\frac{\frac{C^{**}}{4}}{C_{1}C^{*}\left(\mu_{1}-\delta\right)^{-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}N^{*}}}}-C_{4}^{-1}$
 $\geq M,$
@@
 $\text{dim}_{H}\left(C(\lambda,N^{*})\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}.$
-From (3.20), this then implies (3.11) holds, and hence (3.1) holds, for all $\lambda\in[0,\lambda_{0}]$.
-Since $L_{\lambda}\cap\Lambda^{\text{bnd}}_{\lambda}$ is affine equivalent to $\sigma_{\lambda,\alpha}$ and $M$ was arbitrary, this proves Theorem 1.
+From (4.20), this then implies (4.9) holds, for all $\lambda\in[0,\lambda_{0}]$. ∎
+5 Concluding the Proofs
+Proof of Theorem 3..
+Fix a bounded sequence $\overline{a}=(a_{n})$, $a_{n}\in\mathbb{N}$, and denote $\mathcal{A}:=\{a_{n}\}$. Let $\epsilon>0$ and suppose that $M>0$ is so large that
+$\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}>1-\epsilon.$
+By Theorem 4, there is a $\rho>0$ and $\lambda_{0}\in\left(0,\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}\right)$ such that $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$ has Hausdorff dimension larger than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
+Set $\lambda\in(0,\lambda_{0})$.
+Then, since $\lambda_{0}<\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$, from Lemma 5, we know that $\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}$ is the image of $\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)$ under the map $F_{\lambda}$.
+Since $F$ is a local diffeomorphism, it follows that
+$\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)=\text{dim}_{H}\left(\mathcal{C}\left(\lambda,\tilde{O}_{r(\rho)}\right)\right)\geq 1-\epsilon.$
+Proof of Theorem 1..
+Let $\alpha=[a_{1},a_{2},\dots]\in(0,1)$ be an irrational of bounded type.
+Set $\overline{a}=(a_{n}).$ Let $\epsilon>0$.
+By Theorem 3, there is a $\rho>0$ and $\lambda_{0}>0$ such that for any $\lambda\in[0,\lambda_{0}]$, we have
+$\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)\geq 1-\epsilon.$
+Let $\lambda\in[0,\lambda_{0}]$. From Theorem 2, and since $\Lambda^{\text{surv}}_{\lambda}\left(\overline{a},O_{\rho}\right)\subset\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)$, we have
+$\text{dim}_{H}\left(\sigma_{\lambda,\alpha}\right)=\text{dim}_{H}\left(L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}\left(\overline{a}\right)\right)\geq\text{dim}_{H}\left(\Lambda_{\lambda}^{\text{surv}}\left(\overline{a}\right)\cap L_{\lambda}\right)\geq 1-\epsilon.$
+Remark 6.
+We expect that Theorem 1 should hold for all irrational $\alpha\in(0,1)$.
+We would like to point out exactly where the assumption that $\alpha$ is of bounded type is used in our proofs.
+It is used in Proposition B.1, which is the main ingredient of Lemma 4, and this is what later allows us to control the distance of each piece of stable manifold $W^{s}_{\lambda}(Q_{0},m)$ from each $Q_{i}$ in the construction of the rectangles in Subsection 4.6.
+The assumption also implies that the collection of maps $\left\{\hat{T}_{\lambda,a}:\lambda\in[0,\lambda_{0}],\ a\in\mathcal{A}\right\}$ from Section 4 have uniformly bounded $C^{2}$ norms, which is required for the bounded distortion estimate (see Lemma 14).
+It also allows us to control the $C^{2}$-closeness of $\hat{T}_{\lambda,a}$ and $\hat{T}_{0,a}$ in a uniform way, which is used in many arguments such as Theorem 5 and Corollary 1.
+Appendix A Contraction Mappings
+The following proposition is a non-stationary version of the contraction mapping principle.
+Proposition A.1.
+Let $(X,d)$ be a complete metric space and $K\subset X$ a compact set. Suppose that there is a $\mu\in(0,1)$ and, for each $n\in\mathbb{N}$, a map $f_{n}:K\rightarrow K$ such that
+$d\left(f_{n}(x),f_{n}(y)\right)\leq\mu d(x,y),$
+for all $x,y\in K$. Then, the following hold:
+  • (1)
+    There is a $x^{*}\in K$ such that
+    $\lim\limits_{n\rightarrow\infty}f_{1}\circ\cdots\circ f_{n}(x)=x^{*},$
+    for any $x\in K$.
+  • (2)
+    For each $\epsilon>0$, there is a $\epsilon^{\prime}>0$ such that if for each $n\in\mathbb{N}$, there is a mapping $h_{n}:K\rightarrow K$ with
+    $d\left(h_{n}(x),h_{n}(y)\right)\leq\mu^{\prime}d(x,y)\ \text{and}\ \sup_{n}d_{C^{0}(K)}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{n}},h_{n})<\epsilon^{\prime},$
+    for some $\mu^{\prime}\in(0,1)$ and for all $x,y\in K$, then
+    $d\left(x^{*},y^{*}\right)<\epsilon,$
+    where $y^{*}$ is the element such that
+    $\lim\limits_{n\rightarrow\infty}h_{1}\circ\cdots\circ h_{n}(x)=y^{*},$
+    for each $x\in X$.
+Proof.
+  • (1)
+    The proof here is analogous to that of the Banach Fixed Point Theorem. For a proof of this exact result, see Theorem 3 in [LY].
+  • (2)
+    Set $\epsilon^{\prime}=\frac{\epsilon(1-\mu)}{2}$. Then, an induction argument gives that
+    $d\left(f_{1}\circ\cdots\circ f_{n}(x),h_{1}\circ\cdots\circ h_{n}(x)\right)\leq\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1},$
+    for any $x\in K$. We can deduce that
+    $d\left(x^{*},y^{*}\right)\leq d\left(f_{1}\circ\cdots\circ f_{n}(x),x^{*}\right)+d\left(y^{*},h_{1}\circ\cdots\circ h_{n}(x)\right)+\sum_{k=1}^{n}\epsilon^{\prime}\mu^{k-1}$
+    where the last quantity approaches $\frac{\epsilon^{\prime}}{1-\mu}=\frac{\epsilon}{2}$ as $n\rightarrow\infty$.
+Appendix B Continued Fractions
+Given an irrational $\alpha=[a_{1},a_{2},\dots]\in(0,1)$, set $p_{n}(\alpha)\in\mathbb{Z}$ and $q_{n}(\alpha)\in\mathbb{N}$ to be the coprime integers such that
+$\frac{p_{n}(\alpha)}{q_{n}(\alpha)}=[a_{1},a_{2},\dots,a_{n}].$
+We know that these values satisfy the following recurrence equations:
+$p_{k+1}(\alpha)=a_{k+1}p_{k}(\alpha)+p_{k-1}(\alpha),\ p_{1}(\alpha)=p_{0}(\alpha)=1$ (B.1)
+$q_{k+1}(\alpha)=a_{k+1}q_{k}(\alpha)+q_{k-1}(\alpha),\ q_{1}(\alpha)=a_{1},\ q_{0}(\alpha)=1$ (B.2)
+and that
+$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|<\frac{1}{q_{k}(\alpha)}.$ (B.3)
+In regards to an irrational rotation by $\alpha=[a_{1},a_{2},\dots]$ on $S^{1}$, the Three Distance Theorem (see for example Section 2.1 of [I]) states that for each positive integer $n$, the values $\{m\alpha\}$, $1\leq m\leq n$, divide the unit circle into $n+1$ many intervals of at most three possible lengths, one of which is the sum of the other two.
+The precise lengths of the intervals come from how $n$ relates to the values $p_{k}(\alpha)$ and $q_{k}(\alpha)$.
+In particular, if we choose $n=(r+1)q_{k}(\alpha)+q_{k-1}(\alpha)-1$ for some $k\geq 1$ and $1\leq r<a_{k+1}$, then the set of points $\{m\alpha\}$, $1\leq m\leq n$, split $S^{1}$ into $n+1$ many intervals of only two possible lengths given by
+$\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|\ \text{or}\ \left|q_{k-1}(\alpha)\alpha-p_{k-1}(\alpha)\right|-r\left|q_{k}(\alpha)\alpha-p_{k}(\alpha)\right|,$
+and by (B.3), both of these quantities are no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$. As a corollary, we obtain the following result which will be used later on.
+Proposition B.1.
+Suppose that $\mathcal{A}\subset\mathbb{N}$ is finite.
+For each $\epsilon>0$, there is a $n=n(\epsilon,\mathcal{A})\in\mathbb{N}$ such that for any $\alpha=[a_{1},a_{2},\dots]$ with $a_{i}\in\mathcal{A}$, we must have
+$\min\limits_{1\leq m\leq n}\left|m\alpha-y\right|<\epsilon$
+for any $y\in S^{1}$.
+Proof.
+Let $\epsilon>0.$ Set $Y:=\{\alpha=[a_{1},a_{2},\dots]:a_{i}\in\mathcal{A}\}\subset[0,1]\setminus\mathbb{Q}$ and let $a^{\prime}$ be the smallest element of $\mathcal{A}$.
+Then using the recurrence relation (B.2) and an induction argument, we find that for each $k\geq 1$, there is an $M_{k}>0$ such that
+$q_{k}\left(\left[\overline{a^{\prime}}\right]\right)\leq\sup\limits_{\alpha\in Y}|q_{k}(\alpha)|\leq M_{k}.$
+Choose $k$ large enough so that
+$\frac{1}{q_{k-1}\left(\left[\overline{a^{\prime}}\right]\right)}+\frac{1}{q_{k}\left(\left[\overline{a^{\prime}}\right]\right)}<\epsilon.$
+Set $n:=2M_{k}+M_{k-1}-1$ and let $\alpha\in Y$ be arbitrary.
+By the above discussion, for $n_{k}:=2q_{k}(\alpha)+q_{k-1}(\alpha)-1\leq n$, we have that the points $\{m\alpha\}$, $1\leq m<n_{k}$, split $S^{1}$ into $n_{k}+1$ many intervals of lengths no greater than $\frac{1}{q_{k-1}(\alpha)}+\frac{1}{q_{k}(\alpha)}$ which is less than $\epsilon$.
+So, it follows that the set $\{m\alpha\}_{m=1}^{n}$ splits $S^{1}$ into finitely many intervals of length less than $\epsilon$.
 Acknowledgements
-I would like to thank Anton Gorodetski for his guidance and support. This project was supported by NSF grant DMS-2247966 (PI: A. Gorodetski).
+I would like to thank Anton Gorodetski for his guidance and support, and Jake Fillman for looking at a preliminary draft of this text.
+I would also like to thank both anonymous referees for their helpful suggestions and input.
+This project was supported by NSF grant DMS-2247966 (PI: A.
+Gorodetski).